### v5

# Decoupled Data Consistency with Diffusion Purification for Image Restoration

Abstract
Diffusion models have recently gained traction as a powerful class of deep generative priors, excelling in a wide range of image restoration tasks due to their exceptional ability to model data distributions. To solve image restoration problems, many existing techniques achieve data consistency by incorporating additional likelihood gradient steps into the reverse sampling process of diffusion models. However, the additional gradient steps pose a challenge for real-world practical applications as they incur a large computational overhead, thereby increasing inference time. They also present additional difficulties when using accelerated diffusion model samplers, as the number of data consistency steps is limited by the number of reverse sampling steps. In this work, we propose a novel diffusion-based image restoration solver that addresses these issues by decoupling the reverse process from the data consistency steps. Our method involves alternating between a reconstruction phase to maintain data consistency and a refinement phase that enforces the prior via diffusion purification. Our approach demonstrates versatility, making it highly adaptable for efficient problem-solving in latent space. Additionally, it reduces the necessity for numerous sampling steps through the integration of consistency models. The efficacy of our approach is validated through comprehensive experiments across various image restoration tasks, including image denoising, deblurring, inpainting, and super-resolution.
Keywords: diffusion model, inverse problem, image restoration, consistency model, latent diffusion, diffusion purification
1 Introduction
Image restoration problems naturally arise in many computer vision and biomedical imaging applications, examples of which include image denoising, deblurring, inpainting, and super-resolution [1, 2, 3]. The objective of solving an image restoration problem, also known as an inverse problem, is to estimate an unknown signal $\bm{x}_{\star}\in\mathbb{R}^{n}$ from its corrupted or subsampled measurements $\bm{y}\in\mathbb{R}^{m}$ of the following form:
$\bm{y}\;=\;\mathcal{A}(\bm{x}_{\star})+\bm{n},$ (1)
where $\mathcal{A}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a (nonlinear) forward operator, and $\bm{n}\in\mathbb{R}^{m}$ is additive noise.
In practice, the number of observed measurements is often significantly smaller than the input dimension (i.e., $m\ll n$), making the problem highly ill-posed and resulting in infinitely many solutions that satisfy Equation (1). To deal with this challenge, a common approach is to regularize the solution via enforcing image priors. In the field of image restoration, traditional priors such as sparsity have been thoroughly explored [2]. However, recent advancements have shifted focus towards employing deep generative models as learned priors for image restoration. This shift largely came to fruition by the work of Bora et al. [4], who demonstrated that deep generative models could achieve superior image reconstruction quality with fewer number of measurements compared to traditional sparsity-based methods [4, 5, 6]. Following this groundbreaking study, there has been a surge of interest in investigating and exploring various deep generative models for solving image restoration tasks, including deep image priors [7], compressed sensing with generative models (CSGM) [4], and their subsequent developments [1].
More recently, there has been a lot of attention in using diffusion models as image priors, as diffusion models have shown to achieve superior image sampling quality compared to existing generative methods [8, 9, 10, 11]. However, unlike with other types of generative models, adapting diffusion models as image priors to solve inverse problems is not straightforward. This is because the sampling (or generative) process for diffusion models is generally formulated for sampling from an unconditional distribution $p(\bm{x})$, where $p(\bm{x})$ denotes the data distribution. In contrast, solving inverse problems involves sampling from the conditional distribution $p(\bm{x}|\bm{y})$ to ensure data consistency. To sample from this conditional distribution, we need access to the likelihood gradients $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$, which are intractable. Many earlier existing works tackle this issue with projection-type methods that project the unconditional sample $\bm{x}_{t}$ at each iteration $t$ onto the measurement subspace during the sampling process [9, 12, 13, 14, 15]. However, since the projection step is vulnerable to measurement noise and can only be applied to linear inverse problems, more recent works deal with this issue by directly approximating the conditional score function $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ throughout the sampling process under certain assumptions [16, 17, 18].
However, most (if not all) of these existing techniques encounter inherent challenges due to the coupled nature of data consistency and the sampling process. That is, the data consistency step usually takes place within the reverse sampling process of diffusion models, resulting in several fundamental limitations. Firstly, notice that the quality of the reconstructed image is highly dependent on the total number of data consistency steps that we perform. Since data consistency occurs within the reverse sampling process, the total number of data consistency steps can only be at most the total number of sampling steps. This severely limits the usage of accelerated diffusion model samplers, such as denoising diffusion implicit models (DDIMs) [11] and consistency models [19], which perform one-step sampling. Secondly, this coupling also makes it difficult to generalize existing methods to latent (or stable) diffusion, as recent works [20, 21] have demonstrated that simply adding the likelihood gradient for data consistency in the presence of latent diffusion models is insufficient for maintaining data consistency. This restriction hinders usability in large-scale image restoration tasks, where fast and accurate inference is essential.
Figure 1: Illustration of the difference between our proposed method and existing posterior sampling techniques for image restoration. Our method decouples the data consistency step from the diffusion process, whereas existing methods incorporate data consistency into the reverse sampling process of diffusion models by taking an additional gradient step towards the measurement space.
Our Contributions.
In this study, we address these challenges by decoupling the data consistency and reverse sampling processes. Our algorithm involves a two-stage process that alternates between (i) generating initial estimations that are consistent with the measurements by minimizing the data fidelity objective and (ii) refining the initial estimations by incorporating the image prior via diffusion purification [22]. This effectively decouples the two processes, which consequently has two advantages over the state-of-the-art diffusion model methods for image restoration:
  • Framework Versatility. Compared to existing methods, our framework can easily incorporate any diffusion model in a plug-and-play fashion. We demonstrate this flexibility by presenting experimental results that effectively leverage both latent diffusion and consistency models for image restoration.
  • Improved Efficiency. We demonstrate that by decoupling these processes, we can significantly reduce inference time by more than 5×5$\times$5 ×, all while achieving state-of-the-art performance. This contribution is particularly notable because we can incorporate accelerated samplers without leading to insufficient data consistency.
We conduct thorough experiments to validate our method and its advantages across a wide range of image restoration problems, including super-resolution, inpainting, and image deblurring.
Throughout these tasks, we demonstrate that our method can outperform existing methods across several quantitative metrics and perform extensive ablation studies to highlight our approach.
Notation and Organization.
The remainder of this paper is organized as follows. In Section 2, we briefly discuss the essential preliminaries. Subsequently, we introduce our proposed framework in Section 3. Moreover, we further introduce acceleration techniques of our approach in Section 4. In Section 5, we validate the effectiveness of our algorithm with experiments on various image restoration problems. In Sections 6 and 7, we discuss the relationship of our work with prior art, limitation and future directions. All extra technical details are postponed to the appendices.
We denote scalar (function) as regular lower-case letters (e.g. $s_{t},f(t)$), vectors (function) with bold lower-case letters (e.g. $\bm{x},\bm{s}(\bm{x}_{t},t)$). We use $[N]$ to denote the set $\{1,2,\ldots,N\}$, $\mathbb{P}(\cdot)$ to denote the probability, $\mathbb{E}\left[\cdot\right]$ to denote expectation, $||\cdot||_{2}$ to denote L2 norm, $\mathcal{N}(\cdot)$ to denote gaussian distribution. We use $\bm{x}_{t}$ to denote a sample $\bm{x}$ at time step $t$. We use $\bm{x}_{0}$ to denote a clean sample, and use $\bm{s}_{\bm{\theta}}$ and $\bm{s}_{\bm{\phi}}$ for the estimated score functions in the pixel and latent spaces, respectively (discussed in detail later).
2 Preliminaries
In this section, we first briefly review the basic fundamentals of diffusion models [23, 8, 9, 10, 24, 25] and then discuss how to adapt them for solving image restoration problems. Subsequently, we present the main ideas behind diffusion purification, which serve as the backbone of our proposed method.
2.1 Unconditional Diffusion Models
Diffusion models [23, 8, 9, 10] generate an image with two processes, generally referred to as forward and reverse processes. The forward process is to diffuse the data distribution $p_{0}(\bm{x})$ to a standard Gaussian $p_{T}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$ by gradually injecting noise. In the variance-preserving (VP) setting [26], the forward process can be written as a stochastic differential equation (SDE) of the form
$d\bm{x}_{t}\;=\;-\frac{\beta_{t}}{2}\bm{x}_{t}dt+\sqrt{\beta_{t}}d\bm{w},$ (2)
where $\beta_{t}>0$ controls the speed of diffusion and $\bm{w}$ follows the standard Wiener process. To sample an image, we can employ the reverse SDE, which takes the form
$d\bm{x}_{t}\;=\;$ $\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\beta_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]dt+\sqrt{\beta_{t}}d\bar{\bm{w}},$ (3)
where $p_{t}(\bm{x}_{t})$ is the distribution of $\bm{x}_{t}$ at the time $t$, $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the (Stein) score function, and $\bar{\bm{w}}$ follows the reverse Wiener process [27]. In practice, we estimate the score function using a deep neural network $\bm{s}_{\bm{\theta}}:\mathbb{R}^{n}\times[0,1]\rightarrow\mathbb{R}^{n}$ parameterized by $\bm{\theta}$, which is trained using the denoising score matching objective [28]:
$\underset{\bm{\theta}}{\min}\,\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{t}|\bm{x}_{0}}\left[\left\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{0t}(\bm{x}_{t}|\bm{x}_{0})\right\|_{2}^{2}\right],$
where $t\sim\mathcal{U}(0,1)$, $\bm{x}_{0}\sim p_{0}(\bm{x})$, and $\bm{x}_{t}|\bm{x}_{0}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0})$. In practice, both the forward and reverse SDE can be discretized, with
$\bm{x}_{t}$ $=\sqrt{1-\beta_{t}}\bm{x}_{t-1}+\sqrt{\beta_{t}}\bm{\epsilon}_{t-1},$
$\bm{x}_{t-1}$ $=\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{x}_{t}+\beta_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right)+\sqrt{\beta_{t}}\bm{\epsilon}_{t},$
where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\bm{I})$.
2.2 Diffusion Models for Image Restoration Problems
To solve image restoration problems with diffusion models, one of the most commonly used techniques is to modify the unconditional reverse sampling process in Equation 3 by replacing the unconditional score with a conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ based on $\bm{y}$. As such, the conditional reverse SDE has the following form
$d\bm{x}_{t}\;=\;\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\beta_{t}\{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})\}\right]dt+\sqrt{\beta_{t}}d\bar{\bm{w}}.$
Then, by Bayes rule, we can decompose the conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ into
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})=\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})+\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}).$
While the unconditional prior score $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ can be obtained from a pre-trained diffusion model, the likelihood gradient $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})$ is usually intractable.
To tackle this issue, recent works [16, 17, 18] resort to approximating the likelihood term with tractable surrogates, built upon the idea of diffusion posterior sampling (DPS) [16]. DPS approximates the likelihood gradient with
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})\approx\eta\nabla_{\bm{x}_{t}}\|\mathcal{A}(\hat{\bm{x}}_{0}(\bm{x}_{t}))-\bm{y}\|_{2}^{2},$ (4)
where $\eta>0$ can be viewed as a step size for controlling the data consistency strength and $\hat{\bm{x}}_{0}(\bm{x}_{t}):=\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ is the posterior mean, which can be estimated through Tweedie’s formula [29, 30]:
$\hat{\bm{x}}_{0}(\bm{x}_{t})$ $=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t}))$
$\approx\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)),$
where $\bar{\alpha}_{T}=\prod_{t=1}^{T}(1-\beta_{t})$.
However, despite being one of the most recognized algorithms, DPS suffers from several limitations, as previously briefly illustrated. Firstly, the inference time of the reverse sampling process is significantly increased, as the likelihood estimation needs to backpropagate over the entire score function $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ for each iteration $t$ of the sampling process. Since the score function is generally parameterized as a large deep network, this process can be computationally challenging, especially for tasks involving high-dimensional images. Secondly, as the data consistency steps take place within the reverse sampling process, the number of gradient steps for data consistency is limited by the reverse sampling steps, which could lead to insufficient data consistency when using accelerated samplers such as consistency models [19]. Furthermore, the performance of DPS is sensitive to the step size $\eta$ due to the coupled data consistency and reverse sampling. When $\eta$ is small, the reconstructions will not be consistent with the measurements $\bm{y}$. On the other hand, a large $\eta$ may push $\bm{x}_{t}$ away from the noisy data manifold on which the diffusion model is trained, resulting in serious artifacts. Lastly, naively adapting DPS to latent diffusion models often fails due to the nonlinearity of the decoder. Since most diffusion-based image restoration solvers are built upon the idea of DPS, they all suffer from these fundamental issues, motivating the need to decouple the data consistency and reverse sampling processses.
Figure 2: An illustration of our overall framework. Our algorithm involves a total of $K$ iterations which alternate between (i) generating an initial reconstruction by solely optimizing the data fidelity objective and (ii) refining the image via diffusion purification.
2.3 Diffusion Purification
Recently, diffusion purification has shown to be an effective approach for removing adversarial perturbations [22, 31] and for image editing [32]. Specifically, for a given noisy or perturbed image $\bm{x}_{\textrm{adv}}$, this process involves first running the diffusion forward process as described in Equation (2) starting from $\bm{x}_{\textrm{adv}}$ to some intermediate timestep $T$. In the discrete case, $\bm{x}_{T}$ takes the form
$\bm{x}_{T}=\sqrt{\bar{\alpha}_{T}}\bm{x}_{\textrm{adv}}+\sqrt{(1-\bar{\alpha}_{T})}\bm{\epsilon}\>.$ (5)
Starting with $\bm{x}_{T}$, we implement the reverse diffusion process with Equation (3) to produce a “purified” image $\bm{x}_{\textrm{purified}}$, effectively eliminating the adversarial perturbations. The rationale behind this is that by carefully selecting $T$, the adversarial noise within $\bm{x}_{\textrm{adv}}$ becomes overwhelmed by the noise introduced during the forward phase, yet the core semantic information is preserved. Consequently, executing the reverse diffusion process leads to a “clean” image reconstruction. In our work, we use diffusion purification as the backbone of our approach, yet with a distinct objective diverging from existing methods that use diffusion purification to enhance the robustness of DNN-based classifiers [22] or remove artifacts in medical imaging [33, 31].
3 Proposed Methods
In this section, we first introduce our method for image restoration based upon variable splitting (Section 3.1). Then, in Section 3.2, we discuss enforcing data consistency, followed by our proposed method to use diffusion models as an image prior via diffusion purification (Section 3.3). The entire procedure of the proposed method is outlined in Algorithm 1.
3.1 Methods with Variable Splitting
Our proposed algorithm is inspired by the classical variable splitting technique [34, 35, 36, 37] in optimization. Specifically, consider the following regularized optimization formulation for the image restoration problem in Equation 1,
$\min_{\bm{x}}\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{x}),$
where $\mathcal{R}(\cdot)$ is the regularization for image prior, and $\lambda>0$ controls the penalty strength. The idea of variable splitting is to introduce an auxiliary variable $\bm{v}=\bm{x}$ to pose an equivalent problem
$\min_{\bm{x},\bm{v}}\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{v}),\;\mathrm{s.t.}\;\bm{x}=\bm{v}.$ (6)
This constrained optimization problem can be relaxed into an unconstrained problem by the half quadratic splitting technique (HQS) [38, 37], which we can reformulate as
$\min_{\bm{x},\bm{v}}\underbrace{\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\frac{\mu}{2}\|\bm{x}-\bm{v}\|_{2}^{2}+\lambda\mathcal{R}(\bm{v})\>,}_{=:\mathcal{L}(\bm{x},\bm{v})}$ (7)
for some properly chosen $\mu>0$. Then, to solve (7), we can use an iterative method such that for every iteration $k$, we can leverage the extra variable $\bm{v}$ to split the optimization into two sub-problems of the form
$\bm{x}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{x}}\;\mathcal{L}(\bm{x},\bm{v}_{k-1}),$ (8)
$\bm{v}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{v}}\;\mathcal{L}(\bm{x}_{k},\bm{v}),$ (9)
that decouples the data consistency term $\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}$ from the regularization term $\mathcal{R}(\bm{v})$. Then, instead of having an explicit regularizer $\mathcal{R(\cdot)}$ (e.g., $\ell_{1}$-norm to induce sparsity), we can use implicit regularizers. More specifically, we can constrain the output space as the range of the diffusion model via diffusion purification that acts as a form of regularization or prior. To enforce this implicit regularizer, our method involves alternating between the two following steps:
  1. 1.
    Data Consistency. Generate measurement-consistent reconstructions, which may have various artifacts, by solving the data consistency optimization problem Equation 8.
  2. 2.
    Regularization via Diffusion Purification. Refine the obtained reconstruction from the previous step with a pre-trained diffusion model via diffusion purification.
In the following subsections, we dive into the technical details of each step and present an illustration of our method in Figure 2. We also conduct an ablation study in Section 5.3, demonstrating that alternating between these two steps can decrease the distance between the reconstructed and the target image.
Algorithm 1 Image Restoration with Decoupled Data Consistency and Diffusion Model Prior
measurements $\bm{y}$, forward operator $\mathcal{A}(\cdot)$, pretrained model $\bm{s}_{\bm{\theta}}$, total number of iterations $K$, decaying time schedule $\{T_{k}\}_{k\in[K]}$ for diffusion purification, iteration number $\tau$, and $\bm{v}_{0}=\bm{0}\in\mathbb{R}^{n}$
for $k=1,\dots,K$ do
     $\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau)$triangleright solve via gradient based methods
     $\bm{v}_{k}=\texttt{DPUR}(\bm{x}_{k},\bm{s}_{\bm{\theta}},T_{k})$triangleright Diffusion Purification
end for
Return $\bm{v}_{k}$
3.2 Enforcing Data Consistency
Recall that the sub-problem in Equation (8) for maintaining data consistency can be written as
$\bm{x}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{x}}\;\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}+\mu\|\bm{x}-\bm{v}_{k-1}\|_{2}^{2}\>.$
The difficulty behind directly solving the above problem lies in selecting the proper hyperparameter $\mu$, as it may need to be frequently tuned. To tackle this issue, we instead seek an approximate solution $\bm{x}_{k}$ that minimizes solely the data fidelity term $\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}$ but starting the optimization with $\bm{v}_{k-1}$ as the initial point and run a limited number of $\tau$ gradient steps, resulting in a measurement-consistent reconstruction $\bm{x}_{k}$. The rationale behind this is two-fold: (i) by taking $\bm{v}_{k-1}$ as the initial point, we can obtain an output $\bm{x}_{k}$ that is sufficiently close to $\bm{v}_{k-1}$ without having to tune $\mu$, and (ii) taking a limited number of gradient steps $\tau$ reduces the computational burden. Overall, for simplicity, we denote this step as
$\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau).$ (10)
Lastly, we want to highlight that our technique for enforcing data consistency is significantly more efficient than existing methods, as our method does not require back propagation over the entire score function (i.e., the neural network) and hence is (very) computationally efficient. Furthermore, to enforce data consistency, we can use any momentum-based gradient descent optimizer (e.g., SGD, ADAM) for Equation 10. Throughout the experiment we fix the number of gradient steps $\tau$ for each round of the data fidelity optimization. When the measurements are corrupted by noise, one can further reduce the number of iteration to prevent overfitting. We discuss our choice of optimizers in Appendix A.
Figure 3: Demonstrating the regularization effects of diffusion purification. The estimates from the data fidelity step contain artifacts, making the overall reconstruction inaccurate, whereas regularization via diffusion purification can remove such artifacts.
3.3 Implicit Regularization by Diffusion Purification
Due to the underdetermined nature of the problem (1), the initial reconstruction $\bm{x}_{k}$ obtained from Equation 10 will suffer from task specific artifacts, as shown in Figure 3 (the third column). The next step is to refine $\bm{x}_{k}$ via enforcing the image prior learned by the diffusion model. To do so, we propose to utilize diffusion purification described in Section 2.3 as an implicit regularizer for effectively removing the artifacts in $\bm{x}_{k}$ as shown in Figure 3. More specifically, let $T_{k}$ denote the diffusion purification strength at iteration $k$. Following the diffusion purification step in Equation (5), we first define $\bm{x}_{k,0}:=\bm{x}_{k}$ and submerge this sample with noise up to time $T_{k}$:
$\bm{x}_{k,T_{k}}=\sqrt{\bar{\alpha}_{T_{k}}}\bm{x}_{k,0}+\sqrt{1-\bar{\alpha}_{T_{k}}}\bm{\epsilon}.$ (11)
Then, we run unconditional reverse sampling via the reverse SDE in Equation 3 starting from $\bm{x}_{k,T_{k}}$ to obtain the output $\bm{v}_{k}$. Concretely, the discretized version of the reverse SDE can be described as
$\bm{x}_{k,t-1}=$ $\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{x}_{k,t}+\beta_{t}\bm{s}_{\bm{\theta}}(\bm{x}_{k,t},t)\right)+\sqrt{\beta_{t}}\bm{\epsilon}_{t},$ (12)
for $t=T_{k},T_{k}-1,\ldots,1$. Here, recall that $\bm{s}_{\bm{\theta}}(\bm{x},t)$ denotes the score function which can be estimated by a pre-trained unconditional diffusion model. For simplicity, we denote this overall regularization step by
$\bm{v}_{k}=\texttt{DPUR}(\bm{x}_{k},\bm{s}_{\bm{\theta}},T_{k}).$ (13)
For the choice of $T_{k}$, we propose to use a linear decaying schedule for robust performance, which is capable of addressing the “hallucination” issue brought by the stochasticity of the diffusion purification process. In the early iterations of our method, when $T_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though may be inconsistent with the measurements, will be free from artifacts and can serve as good starting point for the subsequent iteration of data consistency optimization. As $T_{k}$ keeps decreasing, the algorithm will converge to an equilibrium point which is both consistent with the measurements and free from the artifacts. More technical details are deferred to Appendix A, and we study the convergence in Section 5.3.
4 Adaptability to Accelerated Samplers & Latent Diffusion
One of the main drawbacks of diffusion models is the increased inference time of the iterative sampling process compared to traditional methods such as variational autoencoders (VAEs) [39] and generative adversarial networks (GANs) [40]. In the literature, many advancements have been made to address this issue, which involve either decreasing the number of iterations via accelerated sampling techniques such as DDIMs [11] or carrying out the sampling process in a lower-dimensional, latent space [24, 25]. In the following sections, we briefly introduce these techniques and demonstrate how they can be used to decrease the inference time of diffusion purification, thereby accelerating our algorithm.
4.1 Faster Diffusion Purification
To improve sampling efficiency, we can adopt the use of accelerated samplers to decrease the sampling time of the reverse process of diffusion purification. For example, we can use the accelerated sampler DDIM, which constructs a non-Markovian chain of the reverse process. By doing so, we can perform each round of diffusion purification with as few as $20$ unconditional reverse sampling steps (i.e., $20$ number of function evaluations (NFEs)), with minimal performance degradation.
Here, we would like to remark that existing methods such as DPS do not have this advantage. DPS typically requires approximately $1000$ reverse sampling steps, as it requires more iterations to ensure “sufficient” data consistency steps.
In the following, we discuss the adaptability of our approach using one-step sampling approaches, including consistency models and approximations via Tweedie’s formula.
Consistency Models.
Recently, Song et al. [19] proposed a class of diffusion models termed consistency models (CMs), which can effectively generate high-quality samples in a single (or at least very few) time steps. CMs build on the fact that there exists a deterministic ODE for every corresponding SDE sampling trajectory of Equation 3 in the form of
$d\bm{x}_{t}\;=\;\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\frac{\beta_{t}}{2}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]dt,$ (14)
with which it shares the same probability densities $p_{t}(\bm{x}_{t})$ for all $t\in[0,1]$ [9]. This implies that for any image $\bm{x}_{0}$, there exists a unique ODE trajectory $\{\bm{x}_{t}\}_{t=0}^{1}$. By leveraging this property, CMs train a deep network $f_{\bm{\theta}}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ parameterized by $\bm{\theta}$ to map any point $\bm{x}_{t}$ to its corresponding $\bm{x}_{0}$ in a single step.
This approach can be seamlessly incorporated into our framework. Specifically, with a pre-trained consistency model $f_{\bm{\theta}}$, the diffusion purification step in Equation 13 in Algorithm 1 can be performed in a single step as
$\bm{v}_{k}=\texttt{DPUR}(\bm{f}_{\bm{\theta}},\bm{x}_{k},T_{k}):=f_{\bm{\theta}}(\bm{x}_{k},T_{k}).$ (15)
Approximation via Tweedie’s Formula.
Under a Gaussian assumption, Tweedie’s formula [29, 30] provides a way of computing the posterior mean by using
$\mathbb{E}\left[\bm{x}_{0}\mid\bm{x}_{t}\right]=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})),$
where $\bm{x}_{0}$ denotes a sample from the “clean” data distribution. Equipped with this formula, instead of running the reverse SDE in Equation 12 for diffusion purification, one can estimate the ground truth in a single step using this formula. Specifically, we can approximate $\bm{v}_{k}$ using
$\bm{v}_{k}$ $=\mathbb{E}\left[\bm{x}_{k,0}\mid\bm{x}_{k,T_{k}}\right]\approx\frac{1}{\sqrt{\bar{\alpha}_{T_{k}}}}(\bm{x}_{k,T_{k}}+(1-\bar{\alpha}_{T_{k}})\bm{s}_{\bm{\theta}}(\bm{x}_{k,T_{k}},T_{k}))\>,$
where the approximation arises from the score function estimation and $\bm{x}_{k,T_{k}}$ is a sample at iteration $k$ diffused upto switching time $T_{k}$ via the forward process.
Compared to reverse SDE sampling, employing Tweedie’s formula will lead to a tradeoff between computation time and reconstruction quality. As evidenced by our experimental results in Section 5, using Tweedie’s formula for approximating the reverse SDE may result in blurry reconstructions with lower perceptual quality, a phenomenon partly attributed to the ”regression to the mean” effects [41].
Algorithm 2 LDM Version of Our Method
measurements $\bm{y}$, forward operator $\mathcal{A}(\cdot)$, pretrained LDM $\bm{s}_{\bm{\phi}}$, a pair of pretrained encoder $\mathcal{E}$ and decoder $\mathcal{D}$, total number of iterations $K$, decaying time schedule $\{T_{k}\}_{k\in[K]}$ for diffusion purification, iteration number $\tau$, and $\widehat{\bm{v}}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$ if enforcing data consistency in the latent space or $\mathcal{D}(\widehat{\bm{v}}_{0})=\bm{0}$ if enforcing data consistency in the pixel space.
for $k=1,\dots,K$ do
     if Approach I: enforcing data consistency in the latent space  then
         $\bm{z}_{k}=\texttt{Data-Fidelity-LDM}(\widehat{\bm{v}}_{k-1},\bm{y},\mathcal{A},\mathcal{D},\tau)$triangleright Enforcing data consistency in the latent space
         $\bm{z}_{k}\leftarrow\mathcal{E}(\mathcal{D}(\bm{z}_{k}))$$\triangleright$ Re-Encode
     else if Approach II: enforcing data consistency in the pixel space  then
         $\bm{x}_{k}=\texttt{Data-Fidelity}(\mathcal{D}(\widehat{\bm{v}}_{k-1}),\bm{y},\mathcal{A},\tau)$triangleright Enforcing data consistency in the pixel space
         $\bm{z}_{k}=\mathcal{E}(\bm{x}_{k})$triangleright Mapping to the latent space
     end if
     $\widehat{\bm{v}}_{k}=\texttt{DPUR}(\bm{z}_{k},\bm{s}_{\bm{\phi}},T_{k})$triangleright Latent diffusion purification
     $\bm{x}_{k}=\mathcal{D}(\bm{v}_{k})$triangleright Mapping to the pixel space
end for
Return $\mathcal{D}(\bm{v}_{k})$
Lastly, notice that this approximation may seem similar to CMs, as they both necessitate only a single NFE. However, CMs avoid some of the tradeoffs discussed previously, as CMs were designed specifically for single-step sampling. Nevertheless, we illustrate the effectiveness of our framework on both techniques, demonstrating that one can use any model that is more readily available.
4.2 Adaptation to Latent Diffusion
Latent diffusion models (LDMs) [24, 25] operate in a lower-dimensional latent space by training a pair of encoder and decoder [42] to address the slow inference time of diffusion models. Specifically, given an encoder $\mathcal{E}:\mathbb{R}^{d}\to\mathbb{R}^{r}$ and a decoder $\mathcal{D}:\mathbb{R}^{r}\to\mathbb{R}^{d}$ where $r\ll d$, LDMs reduce the computational complexity by mapping each $\bm{x}\in\mathbb{R}^{d}$ to $\bm{z}=\mathcal{E}(\bm{x})\in\mathbb{R}^{r}$ and performing the entire diffusion process in the latent space. Then, upon running the reverse steps within the latent space, one can pass the sampled latent code $\bm{z}_{0}$ through the decoder to obtain a synthesized image $\bm{x}_{0}=\mathcal{D}(\bm{z}_{0})$.
Unlike previous methods [20, 21], the flexibility of our framework allows us to (almost) seamlessly incorporate LDMs. Similar to Section 3.1, we alternate between enforcing data consistency and diffusion purification but both in the latent space:
$\bm{z}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{z}}\;\mathcal{L}_{\mathrm{LDM}}(\bm{z},\widehat{\bm{v}}_{k-1}),$ (16)
$\widehat{\bm{v}}_{k}$ $=\texttt{DPUR}(\bm{z}_{k},\bm{s}_{\bm{\phi}},T_{k}),$ (17)
[…]
Approach I: Enforcing DC in the Latent Space.
One way of enforcing data consistency (DC) is strictly in the latent space. Specfically, we can consider the modified problem
$\bm{z}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{z}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}+\frac{\mu}{2}\|\bm{z}-\widehat{\bm{v}}_{k-1}\|_{2}^{2},$
where $\widehat{\bm{v}}_{k-1}$ is the outcome of the previous diffusion purification step in the latent space. We remark that this problem is similar to the CSGM problem considered in [4].
Then, similar to Section 3.2, we can approximately solve the above problem by taking a limited number of gradient based steps on $\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}$ w.r.t. $\bm{z}$ starting from $\widehat{\bm{v}}_{k-1}$:
$\bm{z}_{k}=\texttt{Data-Fidelity-LDM}(\widehat{\bm{v}}_{k-1},\bm{y},\mathcal{A},\mathcal{D},\tau).$ (18)
Before the diffusion purification step, we run an extra step on the solution obtained from Equation 18
$\bm{z}_{k}$ $\leftarrow\mathcal{E}(\mathcal{D}(\bm{z}_{k})),$ (19)
which we label “Re-Encode”. The rationale behind this seemingly unnecessary step is that due to the nonlinearity of the encoder $\mathcal{E}$, there exist infinitely many solutions $\bm{z}$ for the optimization problem in Equation 18. However, since the LDMs are trained only on $\bm{z}=\mathcal{E}(x)$, we need to first map $\bm{z}_{k}$ back to the image space with $\bm{x}_{k}=\mathcal{D}(\bm{z}_{k})$ and then re-encode $x_{k}$ as $z_{k}=\mathcal{E}(\bm{x}_{k})$.
Approach II: Enforcing DC in the Pixel Space.
Alternatively, we can also directly enforce the data consistency in the pixel space. Specifically, given $\widehat{\bm{v}}_{k-1}$ as the outcome of the previous diffusion purification step in the latent space, we can run
$\bm{x}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{x}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\bm{x})\|_{2}^{2}+\frac{\mu}{2}\|\bm{x}-\mathcal{D}(\widehat{\bm{v}}_{k-1})\|_{2}^{2}.$
Again, we can approximately solve the above problem by taking a limited number of gradient based steps on $\frac{1}{2}||\bm{y}-\mathcal{A}(\bm{x})||_{2}^{2}$ w.r.t. $\bm{x}$ starting from $\mathcal{D}(\widehat{\bm{v}}_{k-1})$:
[…]
which is the same as Equation 10 with $\bm{v}_{k-1}=\mathcal{D}(\widehat{\bm{v}}_{k-1})$. Followed by this, we run
$\bm{z}_{k}$ $=\mathcal{E}(\bm{x}_{k}),$
to use the encoder $\mathcal{E}$ to map $\bm{x}_{k}$ into the latent space $\bm{z}_{k}$ for the subsequent diffusion purification step in Equation 17.
Compared to Approach I, it is worth noting that Approach II is much more efficient. This is because here we do not require back propogation through the deep decoder $\mathcal{D}$ for computing the gradient. In practice, both methods are effective for a range of image restoration problems. However, we observe that Approach II tends to be more suitable for specific types of problems like motion deblurring and Gaussian deblurring. This is because it has a better capability to produce $\bm{x}_{k}$’s with superior quality compared to Approach I. See Figures 15 and 14 in Section A.1 for more detail.
5 Experiments
In this section, we demonstrate the effectiveness of our method under various settings. First, we introduce the basic experiment setup in Section 5.1. Second, we show our main experimental results in Section 5.2. Finally, we provide an ablation study in Section 5.3 in terms of noise robustness, computational efficiency, and convergence.
5.1 Experimental Setup
Types of Image Restoration Problems.
To demonstrate the effectiveness of our algorithm, we consider the following image restoration problems:
  • 1)
    Image inpainting with a $100\times 100$ size box mask.
  • 2)
    Image super-resolution with 4×4$\times$4 × down sampling.
  • 3)
    Gaussian image deblurring with Gaussian kernel of size $61\times 61$ and standard deviation $\sigma=3.0$.
  • 4)
    Motion image deblurring with motion deblurring kernel size $61\times 61$ and intensity value $0.5$.
These image restoration problems are commonly explored in numerous relevant works [20, 21, 16, 13] and also arise in many practical applications.
Evaluation Datasets and Metrics.
To demonstrate the effectiveness of our method, we conducted experiments on the FFHQ [43] and LSUN-Bedroom [44] datasets, which have an image size of $256\times 256\times 3$. For CMs, we test our method only on the LSUN-Bedroom dataset as a proof-of-concept. Generally, images of bedrooms have more complicated structures than those in “face” image datasets, making them a good indicator of the efficacy of our algorithm.
To quantitatively evaluate the performance of the algorithms, we use standard metrics: (i) Peak Signal-to-Noise-Ratio (PSNR), (ii) Structural Similarity Index Measure (SSIM) [45] and (iii) Learned Perceptual Image Patch Similarity (LPIPS) [46]. These metrics are commonly used in existing works that evaluate image restoration techniques [21], and are shown to be an accurate indicator of reconstruction performance.
Figure 4: Pixel space diffusion models for solving various image restoration problems. These results demonstrate the effectiveness of our algorithm. We can consistently recover images with the highest PSNR.
Method Super Resolution 4×4$\times$4 × Inpainting (Box)
[…]
DPS [16] $\underline{0.225}$$\pm 0.05$ $25.37\pm 3.17$ $0.717\pm 0.09$ $0.155\pm 0.04$ $\underline{23.58}\pm 3.49$ $0.795\pm 0.09$
MCG [13] $0.282\pm 0.11$ $18.94\pm 4.34$ $0.561\pm 0.15$ $0.159\pm 0.04$ $20.32\pm 3.54$ $0.789\pm 0.06$
ADMM-PnP [47] $0.349\pm 0.04$ $21.56\pm 2.16$ $0.676\pm 0.07$ $0.232\pm 0.02$ $18.41\pm 2.09$ $0.728\pm 0.03$
Ours Pixel (Version 1) $\bm{0.215}\pm 0.05$ $\underline{26.10}\pm 3.12$ $\underline{0.768}\pm 0.08$ $\underline{0.098}\pm 0.01$ $20.79\pm 2.66$ $\underline{0.880}\pm 0.02$
Ours Pixel (Version 2) $0.229\pm 0.05$ $\bm{26.92}\pm 3.33$ $\bm{0.800}\pm 0.07$ $\bm{0.094}\pm 0.02$ $\bm{23.66}\pm 3.42$ $\bm{0.906}\pm 0.02$
Latent-DPS $0.415\pm 0.06$ $22.96\pm 2.62$ $0.573\pm 0.11$ $0.312\pm 0.05$ $18.66\pm 2.53$ $0.656\pm 0.08$
PSLD [20] $0.353\pm 0.07$ $24.53\pm 2.66$ $0.694\pm 0.10$ $0.336\pm 0.07$ $22.01\pm 3.15$ $0.708\pm 0.11$
ReSample [21] $\bm{0.187}\pm 0.04$ $25.92\pm 3.20$ $0.764\pm 0.08$ $\bm{0.195}\pm 0.04$ $21.05\pm 3.11$ $0.762\pm 0.07$
Ours LDMs (Version 1) $\underline{0.244}\pm 0.05$ $\underline{26.52}\pm 3.21$ $\underline{0.778}\pm 0.08$ $\underline{0.215}\pm 0.04$ $\underline{23.75}\pm 3.32$ $\bm{0.831}\pm 0.05$
Ours LDMs (Version 2) $0.259\pm 0.05$ $\bm{26.88}\pm 3.28$ $\bm{0.789}\pm 0.08$ $0.239\pm 0.04$ $\bm{24.22}\pm 3.65$ $\underline{0.828}\pm 0.05$
Ours CMs $0.208\pm 0.05$ $26.17\pm 3.19$ $0.774\pm 0.08$ $0.108\pm 0.02$ $19.66\pm 2.48$ $0.874\pm 0.02$
Table 1: Quantitative results of super resolution and inpainting on the LSUN-Bedrooms dataset for both pixel-based and LDMs. The results colored in blue and in orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models. The best results are in bold and second best results are underlined for pixel space models and LDMs respectively.
Comparison Baselines.
Since we conduct experiments involving both pixel space and latent space, we consider the following baselines:
  • Diffusion Models in Pixel Space. For the pixel space diffusion models, we compare our algorithm to DPS [16], Manifold Constrained Gradients (MCG) [13] and Plug-and-Play ADMM [47] (PnP-ADMM).
  • Diffusion Models in Latent Space. For the experiments with LDMs, we compare to Latent-DPS [21], ReSample [21], and Posterior Sampling with Latent Diffusion Models (PSLD) [20].
To ensure a fair comparison, we use pre-trained pixel-space diffusion models from Chung et al.[16] and Dhariwal et al.[10], as well as pre-trained CMs from Song et al. [19], for both the baseline techniques and our methods in the pixel space. Similarly, for diffusion models in the latent space, we use models provided by Rombach et al. [24], which were previously used in existing methods [21].
Method Gaussian Deblurring Motion Deblurring
[…]
DPS [16] $0.213\pm 0.08$ $24.45\pm 3.72$ $0.691\pm 0.132$ $0.182\pm 0.03$ $24.45\pm 2.93$ $0.736\pm 0.06$
MCG [13] $0.311\pm 0.13$ $17.54\pm 5.06$ $0.551\pm 0.191$ $0.365\pm 0.11$ $20.17\pm 3.73$ $0.515\pm 0.36$
ADMM-PnP [47] $0.437\pm 0.04$ $20.76\pm 1.94$ $0.595\pm 0.09$ $0.524\pm 0.04$ $18.05\pm 2.05$ $0.493\pm 0.10$
Ours Pixel (Version 1) $\bm{0.196}\pm 0.05$ $\underline{27.13}\pm 3.26$ $\underline{0.804}\pm 0.07$ $\underline{0.065}\pm 0.01$ $\underline{33.56}\pm 3.92$ $\underline{0.947}\pm 0.02$
Ours Pixel (Version 2) $\underline{0.212}\pm 0.05$ $\bm{27.74}\pm 3.34$ $\bm{0.825}\pm 0.07$ $\bm{0.067}\pm 0.02$ $\bm{34.47}\pm 4.07$ $\bm{0.956}\pm 0.02$
Latent-DPS $0.337\pm 0.05$ $23.75\pm 2.53$ $0.622\pm 0.10$ $0.425\pm 0.06$ $21.90\pm 2.31$ $0.539\pm 0.10$
PSLD [20] $0.373\pm 0.07$ $24.26\pm 2.84$ $0.683\pm 0.11$ $0.469\pm 0.06$ $20.58\pm 2.32$ $0.562\pm 0.11$
ReSample [21] $\underline{0.240}\pm 0.05$ $25.76\pm 3.02$ $0.731\pm 0.09$ $0.188\pm 0.04$ $27.96\pm 3.07$ $0.806\pm 0.07$
Ours LDMs (Version 1) $0.246\pm 0.05$ $\underline{26.08}\pm 3.02$ $\underline{0.766}\pm 0.07$ $\underline{0.145}\pm 0.03$ $\underline{28.87}\pm 3.77$ $\underline{0.856}\pm 0.06$
Ours LDMs (Version 2) $\bm{0.217}\pm 0.05$ $\bm{27.11}\pm 3.26$ $\bm{0.807}\pm 0.07$ $\bm{0.121}\pm 0.03$ $\bm{29.40}\pm 3.81$ $\bm{0.875}\pm 0.05$
Ours CMs $0.202\pm 0.05$ $26.91\pm 3.11$ $0.805\pm 0.07$ $0.061\pm 0.02$ $34.10\pm 3.80$ $0.953\pm 0.02$
Table 2: Quantitative results of Gaussian and motion deblurring on the LSUN-Bedrooms dataset for both pixel-based and LDMs. The results colored in blue and orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models. The Best Results are in bold and second best results are underlined for pixel space models and LDMs respectively.
5.2 Main Experimental Results
Here, we present both quantitative and qualitative results for our proposed algorithm and the baselines. The algorithms are tested with clean measurements generated from $100$ randomly selected images from each of the FFHQ and LSUN-Bedrooms datasets. For conciseness, we only display results on the LSUN-Bedroom dataset and defer the results on FFHQ to Appendix B. We also defer the choices of hyperparameters in Section A.2.
5.2.1 Pixel Space Diffusion Models
In this section, we present the performance of our proposed method when using diffusion models trained in the pixel space. Recall that we have two versions of our method for the pixel space: vanilla reverse SDE sampling as well as an approximation method via Tweedie’s formula. We label the vanilla reverse sampling method as Version 1 (V1) and the Tweedie’s fomula approximation method as Version 2 (V2). We also demonstrate the use case of our algorithm using CMs as discussed in Section 4.1.
Figure 5: Demonstrating the differences between Versions 1 and 2 on super resolution (4×4$\times$4 ×). The two columns show the reconstructions with PSNRs and their corresponding “canny edges”, respectively.
The qualitative results are reported in Figure 4 and the quantitative comparison of pixel space diffusion model algorithms is shown in Table 1 and Table 2. In these tables, it is evident that our methods noted as V1 and V2 consistently outperform the existing baselines. However, even though the quantitative results imply that V2 is able to achieve higher performance on standard metrics such as PSNR and SSIM, the reconstructions are often blurrier and contain less perceptual detail compared to V1. As shown in Figure 5, the reconstruction of V1 has more high frequency information, visualized with “canny edges”, despite having lower PSNR. This phenomenon is commonly referred to as the “perception-distortion tradeoff” [48, 49].
Surprisingly, we observe that MCG seems to have a lot of variability in the results on the LSUN-Bedroom dataset, but performs well on FFHQ (shown in Appendix B). This might suggest that datasets such as faces are generally easier than datasets with more complicated structures such as bedrooms.
For the results with CMs, we present the quantitative results in Table 1 and  Table 2 highlighted in orange, with visual results in Figure 6. This highlights another use case of our algorithm, as it demonstrates that our algorithm can flexibly utilize CMs compared to existing methods.
Figure 6: Qualitative results for various image restoration problems by using CMs. Our proposed framework is capable of utilizing CMs for image restoration.
Figure 7: Latent space diffusion models for solving various image restoration problems. These results demonstrate the effectiveness of our algorithm. We can consistently recover images with the highest PSNR.
5.2.2 Latent Space Diffusion Models
For the results with LDMs, we display quantitative results in Table 1 and Table 2 highlighted in blue, with qualitative results in Figure 7. Recall that as outlined in Section 4.2, there are two approaches for ensuring data consistency. For the best performance, we use pixel space optimization (Approach 2) for motion deblurring and Gaussian deblurring and latent space optimization (Approach 1) for box inpainting and super-resolution. As observed in the tables, we observe that our algorithm outperforms the existing baselines on many of the inverse problems. Interestingly, we observe that PSLD seems to reconstruct images that are often noisy and blurry, where the learning rate plays a huge role on the reconstruction quality. While ReSample may have higher LPIPS scores on some inverse problem tasks, our algorithm is able to achieve better PSNR and SSIM scores.
5.3 Ablation Studies
Finally, we conduct ablation studies demonstrating computational efficiency, convergence, and robustness to measurement noise.
Computational Efficiency.
First, we compare the inference times (the time it takes to reconstruct a single image) of our method with the baselines and display the results in Table 3. Clearly, for pixel based diffusion models, our methods are much more efficient than DPS [16] and MCG [13], as we do not require backpropogation through the score function for data consistency. Furthermore, using the single step Tweedie’s formula approximation for diffusion purification is much faster than using reverse SDE which involves multiple time steps.
Second, for latent space methods, our methods are also faster than existing methods PSLD [20] and ReSample [21] as observed in Table 3. For our own methods, we observe that optimizing at the pixel level (Approach 2) for data consistency proves to be considerably more efficient than optimizing in the latent space (Approach 1). As previously mentioned, this is due to the need for backpropagating through the complex deep decoder $\mathcal{D}(\cdot)$. This is also one of the main reasons our methods are more computationally efficient than methods such as ReSample.
Algorithm Time (sec)
DPS [16] 387
MCG [13] 385
Our Pixel (Reverse SDE, Version 1) 28.7
Our Pixel (Tweedie’s Formula, Version 2) 2.14
PSLD [20] 188
ReSample [21] 118
Our Latent (Latent DC + Reverse SDE) 109
Our Latent (Pixel DC + Reverse SDE) 12.2
Our Latent (Latent DC + Tweedie’s Formula) 100
Our Latent (Pixel DC + Tweedie’s Formula) 2.59
Table 3: Comparison of inference times. The results colored in blue are for LDMs, whereas the others are for pixel space diffusion models.
Figure 8: Convergence of our methods in terms of LPIPS. The orange, green, red and blue curves correspond to the reconstruction process with data fidelity optimization only, our proposed method with a constant (large) purification strength, constant (small) purification strength and a linear decaying purificaiton strength, respectively. The x-axis denotes the total number of data fidelity gradient steps.
Convergence Study.
In this section, we investigate the convergence of our algorithm in terms of LPIPS. This study also serves to demonstrate the efficiency of a linear decaying time schedule for $T_{k}$ for diffusion purification across various image restoration problems. As shown in Figure 8, at the beginning of each diffusion purification and data fidelity optimization cycle, the LPIPS increases after performing diffusion purification due to stochasticity introduced by the forward-backward diffusion process. However, the subsequent iterations can successfully decrease LPIPS and result in reconstructions with higher qualities compared to pure data fidelity optimization, illustrating that the resulting purified image is able to serve as a good initialization point. In the same figure, also notice that the linear decaying schedule greatly enhances the performance of our algorithm compared to using a constant time schedule for $T_{k}$. This is expected, as a large constant purification strength will result in serious hallucination while a small constant purification strength is not able to fully alleviate the artifacts in the data fidelity based estimations. In Figure 9, we illustrate an example reconstruction process for super-resolution with noisy measurements, with more examples in the Appendix. It is evident that the quality of the reconstructed image gradually increases with more iterations.
Figure 9: Reconstruction process for super-resolution (4x) in the pixel space. To compensate the artifacts introduced by the measurement noise, the reconstruction process takes $K=10$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{10}=200$. The learning rate for data fidelity optimization is $lr=100$.
Super Resolution (4×4$\times$4 ×) Motion Deblurring
Metric DPS [16] Version 1 Version 2 DPS [16] Version 1 Version 2
[…]
Table 4: Case study of noise robustness of our methods compared to DPS on two image restoration problems in the pixel space. We compare our algorithm to DPS on super resolution (4×4$\times$4 ×) and motion deblurring to demonstrate that our method is effective even in the presence of measurement noise. Best results are bolded and second best results are underlined.
Noise Robustness.
Thus far, we considered experiments on image restoration tasks in which the measurements were noiseless. However, in many practical settings of interest, the measurements $\bm{y}$ can be corrupted with additive noise. In this section, we perform an ablation study on the performance of our algorithm in which the measurements have additive Gaussian noise $\mathcal{N}(0,\sigma^{2}_{\bm{y}})$, with varying values of $\sigma_{\bm{y}}$ for super resolution and motion deblurring. We compare our algorithm to DPS, as their theory states that the approximation of their gradient likelihood term actually becomes tighter as the measurement noise $\sigma_{\bm{y}}$ increases [16]. We conduct experiments on $100$ randomly selected images from the LSUN-Bedroom dataset for super-resolution, motion deblurring on standard deviations $0.05,0.075$, and $0.1$. For the hyperparameters, we tune the parameters according to $\sigma_{\bm{y}}=0.01$ and fix them throughout the experiments. The rest of the setup is the same as previously described. We present the results in Table 4. Here we observe that as the standard deviation increases, the performance of all algorithms slightly degrades, as expected. However, our algorithm consistently outperforms DPS in metrics such as PSNR and SSIM, whereas DPS often exhibits higher perceptual quality measured by LPIPS. This difference arises because DPS frequently suffers from hallucinations, introducing artifacts in the image that “sharpen” it. Our algorithm is less susceptible to such artifacts, resulting in “smoother” images with higher signal-to-noise ratios. Details of the hyperparameters and visual results can be found in Section A.3.
6 Relationship to Prior Arts
Here, we discuss the relationship of our work with existing methods for image restoration, including both methods derived in the pixel space and algorithms tailored for the latent space.
Pixel Space Diffusion Models.
Other than DPS [16], there are other diffusion model-based inverse problem solvers available for addressing image restoration tasks. Song et al. [17] introduced $\Pi$GDM, an algorithm that approximates the likelihood gradient using a pseudo-inverse of the linear sensing matrix used for generating the measurements. When the operator $\mathcal{A}(\cdot)$ is linear, it can be represented by a matrix $\bm{A}\in\mathbb{R}^{m\times n}$, for which one can calculate its pseudo-inverse. They leverage this property by first constructing the conditional distribution $p(\bm{y}|\bm{x}_{t})$ in terms of the sensing matrix $\bm{A}$ and derive the likelihood gradient using this conditional distribution. The work by Wang et al. [15] has a similar flavor; since the image restoration problem is ill-posed, the reconstructions without priors generally are a mixture of a sample in both the range and the nullspace of the matrix $\bm{A}$. They use the diffusion model as a prior to essentially “fill in” this nullspace, which is constructed by taking the pseudo-inverse of $\bm{A}$ and used during the reverse sampling process. The work presented in [50] introduced the Direct Diffusion Bridge method building upon the DPS approach. Here, the data fidelity is enforced during the reverse sampling procedure by taking a deeper gradient step which is designed for improved data consistency.
Another popular method for addressing image restoration problems is the Denoising Diffusion Restoration Models (DDRMs) proposed by Kawar et al. [51]. DDRMs construct a probabilistic graph conditioned on the measurements, allowing the use of a pre-trained unconditional diffusion model. Subsequently, they perform the reverse sampling process in the “spectral space”, which is constructed by taking the SVD of the sensing matrix $\bm{A}$. These works generally share two common themes: they require the inverse problem to be linear, and they incorporate the data consistency steps in the reverse diffusion process. Although we do not test on nonlinear inverse problems, our algorithm does not face these limitations and is flexible in this regard.
Lastly, the study presented by Luther et al. [52] shares close ties with Version 2 our algorithm in the pixel space. They introduced Diffusive Denoising of Gradient-based Minimization (DDGM), an image-space method for solving linear tomographic reconstruction problems, which iteratively applies data consistency steps and image denoising through a pre-trained diffusion model. However, as we have shown in Section 5.2, directly regularizing images with Tweedie’s formula will result in blurry images. In contrast, the proposed Version 1 of our algorithm alleviates this issue by performing diffusion purification with multiple diffusion reverse sampling steps. Furthermore, we propose a unified framework that can incorporate accelerated samplers, latent diffusion models, and consistency models, thereby significantly reducing the inference time. Hence, our work addresses a broader spectrum, not only limited to tomographic reconstruction, and can be readily adapted to applications requiring fast inference.
Latent Space Diffusion Models.
Throughout our experiments, we also presented results using LDMs, using PSLD [20] and ReSample [21] as baselines. ReSample [21] is an algorithm designed to solve inverse problems in the latent space that involves solving a data fidelity optimization problem during the reverse sampling process, a process they term “hard data consistency”. The data fidelity objective is similar to the approaches we take, where they alternate between minimizing the objective in either the latent space or the pixel space. Upon solving this objective, they devise a resampling scheme to project the data-consistent sample back onto the noisy manifold to continue the reverse sampling process. Nonetheless, by decoupling these two processes, our algorithm does not require a resampling stage. Additionally, ReSample cannot use accelerated samplers because of the coupling, as it would hinder the data fidelity stage by reducing the number of sampling steps, unlike our framework.
Posterior Sampling with Latent Diffusion (PSLD) [20] extends the idea of DPS to the latent space by addressing the challenge that the decoder and encoder constitute a many-to-one mapping. This implies that there may be multiple latent codes that fit the measurements, and they devise a gradient scheme to ensure the “correct” latent code is estimated during the reverse sampling process. However, their technique involves two gradient computations—one through the LDM decoder and one through the pre-trained LDM score function. Thus, PSLD suffers from increased inference time. Furthermore, it is observed in our experiment that the performance of PSLD is sensitive to the choice of hyperparameters.
Plug-and-Play (PnP) Methods.
The Plug-and-Play Priors (PnP) framework combines physical and learned models for effective computational imaging. PnP alternates between minimizing a data-fidelity term for enforcing data consistency and applying a learned regularizer, typically an image denoiser [53, 36, 37]. The PnP framework is commonly formulated using the Alternating direction method of multipliers (ADMM) [36] and fast iterative shrinkage/thresholding algorithms (FISTA) [54]. Our proposed method shares similarity with the PnP in the sense that both methods utilize image denoisers as the implicit prior. However, while PnP performs denoising directly on $\bm{x}_{k}$ by assuming that $\bm{x}_{k}$ contains certain pseudo noise, our proposed method uses diffusion models to explicitly add noise to $\bm{x}_{k}$ and then denoises it by either iterative sampling or Tweedie’s formula [52] estimation.
7 Discussion & Conclusion
Throughout this paper, we have discussed both the efficiency and flexibility of our algorithm, demonstrating that we can compete with many state-of-the-art techniques while offering numerous advantages. However, despite these strengths, our algorithm has a few steps that warrant further investigation. Recall that the input to the diffusion purification stage is an initial reconstruction estimated from minimizing the data fidelity objective. Notice that if this step does not return an estimate that is meaningful, then the diffusion purification stage cannot produce a high-quality reconstruction. Although this was not an issue for the image restoration problem that we considered, it may be the case for more complex problems, as simply minimizing the data fidelity objective in Equation (10) can lead to poor local minima. For these tasks, it would be of great interest to explore replacing Equation (10) with more advanced solvers.
Furthermore, while our algorithm generally outperforms others in metrics such as PSNR and SSIM, it often only shows marginal improvements in LPIPS, which measures perceptual quality. This outcome is expected, as coupled methods like DPS prioritize the prior step more meticulously, potentially leading to hallucinations. On the other hand, by decoupling data fidelity from the prior, our algorithm places greater emphasis on data fidelity. Therefore, exploring methods to carefully balance between the prior and data consistency with decoupling is an important avenue for future research.
8 Data Availability Statement
The code and instructions for reproducing the experiment results will be made available in the following link: https://github.com/Morefre/Decoupled-Data-Consistency-with-Diffusion-Purification-for-Image-Restoration.
Acknowledgement
QQ, XL, SMK acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, MICDE Catalyst Grant, and a gift grant from KLA. IA and SR acknowlege support from NSF CCF-2212065.
Appendix A Implementation Details
Notation Definition
$K\in\mathbb{N}$ Number of iterations for overall algorithm
$\tau\in\mathbb{N}$ Total number of iterations for data consistency optimization
$\alpha\in\mathbb{R}$ Learning rate for data consistency optimization
$T_{k}\in\mathbb{R}$ Diffusion purification strength at the $k^{\text{th}}$ iteration
$\bm{v}_{k}\in\mathbb{R}^{n}$ Prior consistent sample at $k^{\text{th}}$ iteration
$\bm{x}_{k}\in\mathbb{R}^{n}$ Data consistent sample at $k^{\text{th}}$ iteration
Table 5: Summary of the notation used throughout this work.
A.1 Discussion on Hyperparameters
Here, we discuss the hyperparameters associated with our algorithm and some typical choices we made throughout our experiments.
  • Total Number of Iterations. We run the algorithm for a total of $K$ iterations, which refers to reconstructing the image by alternately applying data fidelity optimization and diffusion purification for $K$ rounds. Empirically, we observe a $K$ ranging from 10 to 20 is sufficient for generating high quality reconstructions. We suggest users start with $K=10$ and then gradually decrease it if a faster computation speed is required.
  • Optimizer for Data Fidelity Optimization. Throughout our experiments, we use momentum-based gradient descent with the momentum hyperparameter set to 0.9. The optimal learning rate (LR) for data fidelity optimization varies from task to task. In general, the LR should neither be too large, which will not result in images with meaningful structures, nor too small, in which case the convergence speed of the overall algorithm will be too slow. Throughout the experiment we fix the number of gradient steps for each round of the data fidelity optimization. When the measurements are corrupted by noise, one can further perform early stopping to prevent overfitting to the noise.
  • Decaying Time Schedule for Diffusion Purification. A decaying time schedule plays the key role in the success of our proposed algorithm. As described in  Section 3.3, given a data fidelity based reconstruction $\bm{x}_{k}$, the goal of diffusion purification is to generate $\bm{v}_{k}$ that shares the overall structural and semantic information of $\bm{x}_{k}$, such that it is consistent with the measurements while at the same time being free from the artifacts visible in $\bm{x}_{k}$. Unfortunately, the $\bm{v}_{k}$ resulting from diffusion purification is sensitive to the choice of $T$, i.e., the amount of noise added to $\bm{x}_{k}$. When $T$ is small, the resulting $\bm{v}_{k}$ will exhibit similar structure to $\bm{x}_{k}$ but with artifacts, as there is not enough noise to submerge these artifacts through diffusion purification. On the other hand, when $T$ is too large, though $\bm{v}_{k}$ will be free from artifacts, its semantic information and overall structure will significantly deviate from $\bm{x}_{k}$. As a result, the resulting reconstructions will not be measurement-consistent.
    To this end, previous works that utilize diffusion purification [31, 14] depend on fine-tuning an optimal $T^{\star}$. However, we find that the fine-tuning strategy can be infeasible since the optimal $T^{\star}$ varies from task to task and even from image to image. To alleviate the need to identify the optimal $T^{\star}$, we implement a linear decaying schedule $\{T_{k}\}_{k\in[K]}$ for each iteration $k$. In the early iterations when $T_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though inconsistent with the measurements, will be free from artifacts and can serve as good starting point for the subsequent iteration of data fidelity optimization. As $T_{k}$ keeps decreasing, the algorithm will converge at an equilibrium point which is both consistent with the measurements and free from the artifacts. When the measurements are noiseless, we observe that for a diffusion model trained with 1000 total timesteps, a linear decaying time schedule starting from $T_{1}=400$ and ending at $T_{K}=0$ results in reconstructions with high quality for Super-Resolution, Gaussian Deblurring and Motion Deblurring. For Box Inpainting, we choose a higher initial timestep $T_{1}=700$ and $T_{1}=500$ for pixel space diffusion models and LDMs respectively to compensate the severe artifacts in the large box region. When the measurements are corrupted by noise, one may instead end at $T_{K}=T$, where $T>0$ to compensate the negative effect of the measurement noise. We demonstrate the reconstruction process for various tasks in Figures 10, 11, 12, 13 and 16. Notice that the diffusion purification removes the artifacts from the initial reconstructions obtained via data fidelity optimization especially in the early iterations and results in high quality reconstructions consistent with the measurements as timestep decreases.
    Figure 10: Reconstruction process for super-resolution with pixel space diffusion models. The reconstruction process takes $K=10$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{10}=0$. We demonstrate reconstructions at timesteps 1, 5, and 10.
    Figure 11: Reconstruction Process for Gaussian Deblurring with Pixel Space Diffusion Models. The reconstruction process takes $K=10$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{10}=0$. Due to the space limit, only reconstructions at timesteps 1, 5 and 10 are shown.
    Figure 12: Reconstruction Process for Inpainting with Pixel Space Diffusion Models. The reconstruction process takes $K=20$ iterations in total with a time schedule starting at $T_{1}=700$ and linearly decaying to $T_{20}=0$. Due to the space limit, only reconstructions at timesteps 1, 10 and 20 are shown.
    Figure 13: Reconstruction Process for Motion Deblurring with Pixel Space Diffusion Models. The reconstruction process takes $K=20$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{20}=0$. Due to the space limit, only reconstructions at timesteps 1, 10 and 20 are shown.
  • DDIM Steps for Diffusion Purification. With the help of accelerated diffusion samplers such as DDIMs [11], we can perform diffusion purification with as low as 20 unconditional sampling steps. As mentioned before, the total number of iterations $K$ ranging from 10-20 is sufficient for generating high quality reconstructions. This amounts to 200 to 400 unconditional reverse sampling steps in total, a significantly lower figure compared to DPS [16], which requires 1000 unconditional reverse sampling steps. Interestingly, we find that increasing the DDIM sampling steps does not necessarily improve the performance of diffusion purification. Therefore, we suggest users use 20 steps for diffusion purification.
  • Pixel Space Optimization and Latent Space Optimization. As discussed in Section 4.2, in the case of LDMs, one can use either Pixel Space Optimization Equation (20) or Latent Space Optimization (18) to enforce the data consistency. Depending on the specific task, one approach may generate $\bm{x}_{k}$’s with higher quality than the other. In practice, as shown in Figures 14 and 15, we observe that for motion deblurring and Gaussian deblurring, the initial reconstructions generated with Pixel Space Optimization have significantly higher quality compared to the ones generated with Latent Space Optimization. Furthermore, Pixel Space Optimization is much faster than the Latent Space Optimization since the gradient computation of the former approach does not require backpropagation through the deep decoder $\mathcal{D}(\cdot)$. To achieve the best reconstruction quality, we suggest users try both options and choose the one which gives the best performance. On the other hand, for applications that require fast reconstruction speed, we suggest using Pixel Space Optimization.
Figure 14: Comparison between Pixel Space Optimization and Latent Space Optimization for Motion Deblurring. Initial Reconstructions obtained via Pixel Space Pptimization have much better quality compared to the ones obtained from Latent Space Optimization which in turn result in final reconstructions with higher quality.
Figure 15: Comparison between Pixel Space Optimization and Latent Space Optimization for Gaussian Deblurring. Initial Reconstructions obtained via Pixel Space Optimization have much better quality compared to the ones obtained from Latent Space Optimization which in turn result in final reconstructions with higher quality.
A.2 Choice of Hyperparameters
In this section, we specify the hyperparameters used for generating the results in Section 5. For the baselines, we either followed the instructions given in each respective work or used the parameters provided in their code.
A.2.1 Experiments with Pixel Space Diffusion Models
For the hyperparameters used for the pixel space experiments, we tabulate all of our values for each dataset in Table 6. Recall that for $T_{k}$, we choose a linearly decaying time schedule. We start at a value $T_{0}$, which is specified in the table. For box inpainting, we used a larger value of $T_{0}$ to compensate for the large artifacts that exist within the box region. The diffusion purification is performed with 20 deterministic DDIM steps.
We optimize the data fidelity objective with momentum-based gradient descent with the momentum hyperparameter set to $0.9$ for both FFHQ and LSUN-Bedrooms datasets. To achieve the best performance, we use Pixel Space Optimization for Gaussian deblurring and motion deblurring while using Latent Space Optimization for box inpainting and super-resolution. Again, we fix the total number of gradient steps as 1000 throughout the $K$ iterations, which means we take $\frac{1000}{K}$ gradient steps in each round of data fidelity optimization.
Dataset Super Resolution (4×4$\times$4 ×) Gaussian Deblurring Motion Deblurring Inpainting (Box)
$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
LSUN-Bedrooms $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $20$ $400$ $10^{3}$ $20$ $700$
FFHQ $10^{3}$ 10 $400$ $10^{5}$ $10$ $400$ $10^{5}$ $20$ $400$ $10^{3}$ $20$ $700$
Table 6: Table of hyperparameters used for experiments in the pixel space. $\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule.
Dataset Super Resolution (4×4$\times$4 ×) Gaussian Deblurring Motion Deblurring Inpainting (Box)
$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
LSUN-Bedrooms $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{3}$ $20$ $500$
FFHQ $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $4$ $400$ $10^{3}$ $20$ $500$
Table 7: Table of hyperparameters used for experiments in the latent space. $\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule.
Dataset Super Resolution (4×4$\times$4 ×) Gaussian Deblurring Motion Deblurring Inpainting (Box)
$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
LSUN-Bedrooms $10^{3}$ $20$ $1$ $10^{5}$ $20$ $1$ $10^{5}$ $20$ $1$ $10^{3}$ $20$ $5$
Table 8: Table of hyperparameters used for experiments with CMs. $\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule. Different from the typical diffusion models used in previous experiments which use a noise scale in the range of [0,1000], CMs are trained in the range of [0.002,80].
Method Super Resolution 4×4$\times$4 × Inpainting (Box)
[…]
DPS [16] $\underline{0.176}$$\pm 0.04$ $28.45\pm 1.99$ $0.792\pm 0.05$ $0.115\pm 0.02$ $23.14\pm 3.14$ $0.829\pm 0.03$
MCG [13] $0.182\pm 0.03$ $24.45\pm 2.93$ $0.736\pm 0.06$ $0.119\pm 0.02$ $22.77\pm 3.06$ $0.827\pm 0.03$
ADMM-PnP [47] $0.307\pm 0.03$ $21.54\pm 3.12$ $0.766\pm 0.03$ $0.251\pm 0.02$ $18.26\pm 1.86$ $0.724\pm 0.02$
Ours Pixel (Version 1) $\bm{0.154}\pm 0.03$ $\underline{29.43}\pm 2.01$ $\underline{0.844}\pm 0.03$ $\underline{0.052}\pm 0.01$ $\underline{25.54}\pm 3.01$ $\underline{0.917}\pm 0.01$
Ours Pixel (Version 2) $0.185\pm 0.03$ $\bm{30.86}\pm 2.11$ $\bm{0.878}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\bm{28.02}\pm 2.74$ $\bm{0.935}\pm 0.01$
PSLD [20] $0.309\pm 0.06$ $26.22\pm 2.07$ $0.758\pm 0.06$ $0.248\pm 0.06$ $25.42\pm 2.15$ $0.788\pm 0.05$
Ours LDMs (Version 1) $\bm{0.183}\pm 0.04$ $\underline{29.16}\pm 1.65$ $\underline{0.826}\pm 0.03$ $\bm{0.166}\pm 0.03$ $\underline{26.50}\pm 2.20$ $\underline{0.860}\pm 0.02$
Ours LDMs (Version 2) $\underline{0.215}\pm 0.04$ $\bm{30.61}\pm 1.97$ $\bm{0.866}\pm 0.03$ $\underline{0.204}\pm 0.04$ $\bm{27.47}\pm 2.05$ $\bm{0.868}\pm 0.03$
Table 9: Quantitative results of super resolution and inpainting on the FFHQ dataset for both pixel-based and LDMs. The results colored in blue are for LDMs, while the others are for pixel space diffusion models. The Best Results are in bold and second best results are underlined for pixel space models and LDMs respectively.
Method Gaussian Deblurring Motion Deblurring
[…]
DPS [16] $\underline{0.162}\pm 0.03$ $27.97\pm 2.01$ $0.782\pm 0.05$ $0.219\pm 0.04$ $25.34\pm 2.21$ $0.690\pm 0.07$
MCG [13] $0.164\pm 0.03$ $25.03\pm 3.62$ $0.790\pm 0.06$ $0.343\pm 0.04$ $24.88\pm 2.01$ $0.708\pm 0.05$
ADMM-PnP [47] $0.421\pm 0.03$ $20.86\pm 2.16$ $0.664\pm 0.05$ $0.522\pm 0.02$ $17.91\pm 1.99$ $0.547\pm 0.06$
Ours Pixel (Version 1) $\bm{0.140}\pm 0.02$ $\underline{30.41}\pm 1.93$ $\underline{0.864}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\underline{37.56}\pm 1.77$ $\underline{0.963}\pm 0.01$
Ours Pixel (Version 2) $0.173\pm 0.03$ $\bm{31.67}\pm 2.12$ $\bm{0.891}\pm 0.03$ $\underline{0.060}\pm 0.01$ $\bm{38.37}\pm 1.88$ $\bm{0.970}\pm 0.01$
PSLD [20] $0.270\pm 0.05$ $27.46\pm 1.84$ $0.790\pm 0.05$ $0.379\pm 0.05$ $22.35\pm 1.85$ $0.641\pm 0.07$
Ours LDMs (Version 1) $\underline{0.198}\pm 0.03$ $\underline{28.91}\pm 1.98$ $\underline{0.825}\pm 0.04$ $\underline{0.152}\pm 0.03$ $\underline{30.11}\pm 1.93$ $\underline{0.851}\pm 0.03$
Ours LDMs (Version 2) $\bm{0.171}\pm 0.02$ $\bm{30.45}\pm 1.86$ $\bm{0.877}\pm 0.03$ $\bm{0.108}\pm 0.02$ $\bm{31.48}\pm 1.94$ $\bm{0.894}\pm 0.02$
Table 10: Quantitative results of Gaussian and motion deblurring on the FFHQ dataset for both pixel-based and LDMs. The results colored in blue are for LDMs, while the others are for pixel space diffusion models. The Best Results are in bold and second best results are underlined for pixel space models and LDMs respectively.
A.2.2 Experiments with LDMs
Similarly, we specify the hyperparameters used for experiments using LDMs in Table 7. For box inpainting, similar to the pixel space experiments, we used a larger value of $T_{0}$ to compensate for the artifacts seen in the box region. The diffusion purification is performed with 20 deterministic DDIM steps. For the data fidelity objectives, the same procedure done in the pixel space was applied for the LDM case.
A.2.3 Experiments with CMs
Lastly, we specify the hyperparameters used for experiments using CMs in Table 8. Notice that different from the typical diffusion model used previously which use a noise scale in the range of [0,1000], CMs are traind in the range of [0.002,80]. Again, the diffusion purification is performed with 20 deterministic DDIM steps.
A.3 Dealing with Measurement Noise
When the measurements are noiseless, a linear decaying schedule starts from $T_{0}$ and ends at $T_{K}=0$ is capable of resulting in high quality reconstructions and we can use a relatively high learning rate for the data fidelity optimization. On the contrary, in the presence of measurement noise, the data fidelity optimization in (10) will inevitably introduce additional artifacts even if $\bm{v}_{k-1}$ is the ground truth image. To achieve robust recovery, one may instead: (i) use a relatively small learning rate for data fidelity optimization to prevent overfitting to the noise, and (ii) end the time schedule at $t_{K}=T$, where $T>0$, to compensate for the artifacts resulting from the measurement noise. As shown in  Figure 16, compared to the clean measurements case, due to the additive noise, the data fidelity optimization always introduce artifacts in $\bm{x}_{k}$. This issue can be alleviated by utilizing a non-zero diffusion purification strength at the final iteration. For the experiments in Section 5.3, all the hyperparameters are kept the same as described in Section A.2.1, but with the following changes:
  • The learning rate for motion deblurring changes from $1e^{5}$ to $1e^{4}$.
  • The time schedules for super-resolution and motion deblurring start from $t_{0}=400$ and decay linearly to $250$ and $300$ respectively.
With the changes made above, our algorithm achieves reconstruction performance comparable to that of DPS with measurement noise. Notice that the hyperparameters are chosen such that they work under the most severe noise level, i.e., $\sigma=0.1$. Better performance can be achieved by decreasing $t_{K}$ in the case of $\sigma=0.05$ or $\sigma=0.075$.
Appendix B Additional Quantitative Results
In this section, we provide additional quantitative results for solving inverse problems on the FFHQ dataset. The results are presented in Tables 9 and 10. Throughout all of our experiments, we observe that our proposed algorithms outperform the baselines for various image restoration problems.
(a) Reference and Measurements for Super-Resolution (4x). The noisy measurements are corrupted by random Gaussian noise with $\sigma=0.1.$
(b) Reconstruction Process with Clean Measurements. The reconstruction process takes $K=10$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{10}=0$. The learning rate for data fidelity optimization is $lr=1000$. Due to the space limit, only reconstructions at timesteps 1, 3, 5, 7 and 10 are shown.
(c) Reconstruction Process with Noisy Measurements. To compensate the artifacts introduced by the measurement noise, the reconstruction process takes $K=10$ iterations in total with a time schedule starting at $T_{1}=400$ and linearly decaying to $T_{10}=200$. The learning rate for data fidelity optimization is $lr=100$.
Figure 16: Reconstruction Process for Super-Resolution (4x) with Pixel Space Diffusion Models.

### v6

@@
-Decoupled Data Consistency with Diffusion Purification for Image Restoration
+DCDP: Decoupled Data Consistency via Diffusion Purification for Solving General Inverse Problems
 Abstract
-Diffusion models have recently gained traction as a powerful class of deep generative priors, excelling in a wide range of image restoration tasks due to their exceptional ability to model data distributions.
-To solve image restoration problems, many existing techniques achieve data consistency by incorporating additional likelihood gradient steps into the reverse sampling process of diffusion models.
-However, the additional gradient steps pose a challenge for real-world practical applications as they incur a large computational overhead, thereby increasing inference time.
-They also present additional difficulties when using accelerated diffusion model samplers, as the number of data consistency steps is limited by the number of reverse sampling steps.
-In this work, we propose a novel diffusion-based image restoration solver that addresses these issues by decoupling the reverse process from the data consistency steps.
-Our method involves alternating between a reconstruction phase to maintain data consistency and a refinement phase that enforces the prior via diffusion purification.
-Our approach demonstrates versatility, making it highly adaptable for efficient problem-solving in latent space.
-Additionally, it reduces the necessity for numerous sampling steps through the integration of consistency models.
-The efficacy of our approach is validated through comprehensive experiments across various image restoration tasks, including image denoising, deblurring, inpainting, and super-resolution.
-Keywords: diffusion model, inverse problem, image restoration, consistency model, latent diffusion, diffusion purification
+Diffusion models have emerged as powerful generative priors, excelling in solving inverse problems owing to their ability to model complex data distributions.
+Existing methods typically enforce data consistency by guiding the reverse sampling process with additional likelihood gradients.
+However, computing these gradients introduces significant computational overhead, prolongs inference time, and most importantly constrains compatibility with accelerated samplers.
+To overcome these limitations, we propose DCDP (Decoupled Data Consistency via Diffusion Purification), a two-stage diffusion-based framework for solving inverse problems that disentangles data consistency from reverse sampling steps.
+Specifically, DCDP iteratively alternates between the following two phases: (i) a reconstruction phase that enforces data consistency, followed by (ii) a refinement phase that applies diffusion purification to impose the learned generative prior.
+DCDP is broadly applicable to pixel space and latent space diffusion, and consistency models, achieving significant speedup in inference time compared to most state-of-the-art methods.
+We demonstrate its effectiveness and generality across a range of inverse problems, including image denoising, (nonlinear) deblurring, inpainting, and super-resolution.
+Keywords.
+Diffusion model, inverse problems, latent diffusion, diffusion purification, plug-and-play.
+Originally released on arXiv in March 2024; revised in June 2025.
 1 Introduction
-Image restoration problems naturally arise in many computer vision and biomedical imaging applications, examples of which include image denoising, deblurring, inpainting, and super-resolution [1, 2, 3].
-The objective of solving an image restoration problem, also known as an inverse problem, is to estimate an unknown signal $\bm{x}_{\star}\in\mathbb{R}^{n}$ from its corrupted or subsampled measurements $\bm{y}\in\mathbb{R}^{m}$ of the following form:
-$\bm{y}\;=\;\mathcal{A}(\bm{x}_{\star})+\bm{n},$ (1)
-where $\mathcal{A}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a (nonlinear) forward operator, and $\bm{n}\in\mathbb{R}^{m}$ is additive noise.
-In practice, the number of observed measurements is often significantly smaller than the input dimension (i.e., $m\ll n$), making the problem highly ill-posed and resulting in infinitely many solutions that satisfy Equation (1).
-To deal with this challenge, a common approach is to regularize the solution via enforcing image priors.
-In the field of image restoration, traditional priors such as sparsity have been thoroughly explored [2].
-However, recent advancements have shifted focus towards employing deep generative models as learned priors for image restoration.
-This shift largely came to fruition by the work of Bora et al.
-[4], who demonstrated that deep generative models could achieve superior image reconstruction quality with fewer number of measurements compared to traditional sparsity-based methods [4, 5, 6].
-Following this groundbreaking study, there has been a surge of interest in investigating and exploring various deep generative models for solving image restoration tasks, including deep image priors [7], compressed sensing with generative models (CSGM) [4], and their subsequent developments [1].
-More recently, there has been a lot of attention in using diffusion models as image priors, as diffusion models have shown to achieve superior image sampling quality compared to existing generative methods [8, 9, 10, 11].
-However, unlike with other types of generative models, adapting diffusion models as image priors to solve inverse problems is not straightforward.
-This is because the sampling (or generative) process for diffusion models is generally formulated for sampling from an unconditional distribution $p(\bm{x})$, where $p(\bm{x})$ denotes the data distribution.
-In contrast, solving inverse problems involves sampling from the conditional distribution $p(\bm{x}|\bm{y})$ to ensure data consistency.
-To sample from this conditional distribution, we need access to the likelihood gradients $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$, which are intractable.
-Many earlier existing works tackle this issue with projection-type methods that project the unconditional sample $\bm{x}_{t}$ at each iteration $t$ onto the measurement subspace during the sampling process [9, 12, 13, 14, 15].
-However, since the projection step is vulnerable to measurement noise and can only be applied to linear inverse problems, more recent works deal with this issue by directly approximating the conditional score function $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ throughout the sampling process under certain assumptions [16, 17, 18].
-However, most (if not all) of these existing techniques encounter inherent challenges due to the coupled nature of data consistency and the sampling process.
-That is, the data consistency step usually takes place within the reverse sampling process of diffusion models, resulting in several fundamental limitations.
-Firstly, notice that the quality of the reconstructed image is highly dependent on the total number of data consistency steps that we perform.
-Since data consistency occurs within the reverse sampling process, the total number of data consistency steps can only be at most the total number of sampling steps.
-This severely limits the usage of accelerated diffusion model samplers, such as denoising diffusion implicit models (DDIMs) [11] and consistency models [19], which perform one-step sampling.
-Secondly, this coupling also makes it difficult to generalize existing methods to latent (or stable) diffusion, as recent works [20, 21] have demonstrated that simply adding the likelihood gradient for data consistency in the presence of latent diffusion models is insufficient for maintaining data consistency.
-This restriction hinders usability in large-scale image restoration tasks, where fast and accurate inference is essential.
-Figure 1: Illustration of the difference between our proposed method and existing posterior sampling techniques for image restoration.
-Our method decouples the data consistency step from the diffusion process, whereas existing methods incorporate data consistency into the reverse sampling process of diffusion models by taking an additional gradient step towards the measurement space.
+Inverse problems naturally arise in many image processing and scientific applications, such as image denoising, deblurring, inpainting, super-resolution, and phase retrieval [1, 2, 3].
+Mathematically, the goal of solving an inverse problem is to estimate an unknown signal $\bm{x}^{\star}\in\mathbb{R}^{n}$ from its subsampled (and possibly corrupted) measurements $\bm{y}\in\mathbb{R}^{m}$:
+$\bm{y}\;=\;\mathcal{A}(\bm{x}^{\star})+\bm{n},$ (1)
+where $\mathcal{A}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is a linear or nonlinear forward operator and $\bm{n}\in\mathbb{R}^{m}$ is the additive noise.
+Typically, these inverse problems are highly ill-posed with $m\ll n$, resulting in infinitely many solutions that satisfy (1).
+A common approach to address this challenge is to regularize the solution by enforcing certain data priors.
+In the past, many classical handcrafted priors such as sparsity or total variation have been investigated in the literature [2].
+However, handcrafted priors are often overly simplistic, limiting their ability to capture the complex structures inherent in natural images.
+To deal with the challenge, deep generative models have emerged as better priors, significantly outperforming handcrafted priors in solving inverse problems.
+This paradigm shift was pioneered by recent seminal work such as [4], which showed that deep generative priors enable superior image reconstruction with fewer measurements than traditional sparsity-based methods; see also subsequent developments [1].
+More recently, diffusion models have gained increasing attention as image priors [5, 6, 7, 8].
+These models represent a novel class of deep generative models that are capable of producing high-quality samples thanks to their powerful ability to model complex data distributions.
+Specifically, diffusion models operate by learning to reverse a forward process that progressively adds Gaussian noise to the data till it becomes indistinguishable from pure noise.
+The key to performing the reverse process is learning the score functions $\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})$, i.e., the gradients of the log-density of the noise-mollified data $\bm{x}_{t}$, which can be approximated with deep networks trained with the denoising score matching objective [9].
+When applied to solving inverse problems, the reverse process needs to be guided by additional likelihood gradients $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ to ensure consistency with the measurements $\bm{y}$.
+However, since the likelihood gradient $\nabla_{\bm{x}_{t}}\log p(\bm{y}|\bm{x}_{t})$ is generally intractable, existing methods either iteratively project $\bm{x}_{t}$ onto the measurement subspace [6, 10, 11, 12, 13], or approximate the gradient under simplified assumptions [14, 15, 16].
+These approaches face fundamental challenges due to the coupling of data consistency and the reverse sampling process.
+Specifically, data consistency is typically enforced at each diffusion step, so that the total number of consistency updates is constrained by the number of reverse steps.
+As a result, achieving high reconstruction quality often requires a large number of diffusion steps, limiting compatibility with accelerated samplers such as Denoising Diffusion Implicit Models (DDIMs) [8] or consistency models (CMs) [17], which operate with fewer steps by improved designs.
+Moreover, this coupling complicates the extension to latent diffusion models: recent works [18, 19] demonstrate that naively incorporating the likelihood gradient fails to ensure data consistency in the latent space, posing a significant challenge for fast and accurate inference in high-dimensional inverse problems [20].
+Figure 1: Illustration of the difference between our proposed method and existing posterior sampling techniques for solving inverse problems.
+Our method decouples data consistency from the diffusion process, unlike existing approaches that enforce it via gradient steps during reverse sampling.
 Our Contributions.
-In this study, we address these challenges by decoupling the data consistency and reverse sampling processes.
-Our algorithm involves a two-stage process that alternates between (i) generating initial estimations that are consistent with the measurements by minimizing the data fidelity objective and (ii) refining the initial estimations by incorporating the image prior via diffusion purification [22].
-This effectively decouples the two processes, which consequently has two advantages over the state-of-the-art diffusion model methods for image restoration:
-  • Framework Versatility.
-Compared to existing methods, our framework can easily incorporate any diffusion model in a plug-and-play fashion.
-We demonstrate this flexibility by presenting experimental results that effectively leverage both latent diffusion and consistency models for image restoration.
-  • Improved Efficiency.
-We demonstrate that by decoupling these processes, we can significantly reduce inference time by more than 5×5$\times$5 ×, all while achieving state-of-the-art performance.
-This contribution is particularly notable because we can incorporate accelerated samplers without leading to insufficient data consistency.
-We conduct thorough experiments to validate our method and its advantages across a wide range of image restoration problems, including super-resolution, inpainting, and image deblurring.
-Throughout these tasks, we demonstrate that our method can outperform existing methods across several quantitative metrics and perform extensive ablation studies to highlight our approach.
-Notation and Organization.
-The remainder of this paper is organized as follows.
-In Section 2, we briefly discuss the essential preliminaries.
-Subsequently, we introduce our proposed framework in Section 3.
-Moreover, we further introduce acceleration techniques of our approach in Section 4.
-In Section 5, we validate the effectiveness of our algorithm with experiments on various image restoration problems.
-In Sections 6 and 7, we discuss the relationship of our work with prior art, limitation and future directions.
-All extra technical details are postponed to the appendices.
-We denote scalar (function) as regular lower-case letters (e.g.
-$s_{t},f(t)$), vectors (function) with bold lower-case letters (e.g.
-$\bm{x},\bm{s}(\bm{x}_{t},t)$).
-We use $[N]$ to denote the set $\{1,2,\ldots,N\}$, $\mathbb{P}(\cdot)$ to denote the probability, $\mathbb{E}\left[\cdot\right]$ to denote expectation, $||\cdot||_{2}$ to denote L2 norm, $\mathcal{N}(\cdot)$ to denote gaussian distribution.
+In this study, we tackle the challenges of solving general inverse problems with diffusion models by introducing a new method termed Decoupled Data Consistency via Diffusion Purification (DCDP), which decouples data consistency steps from the reverse sampling process.
+As illustrated in Figure 1, our algorithm iteratively alternates between: (i) a reconstruction phase that enforces measurement consistency by minimizing a data fidelity objective, and (ii) a refinement phase that refines the estimate using a diffusion-based image prior via diffusion purification [21].
+Importantly, the data consistency update is executed independently of the reverse diffusion process, offering two key advantages over existing diffusion-based inverse problem solvers:
+  • Versatility.
+Our framework can seamlessly integrate a broad range of diffusion models in a plug-and-play fashion, including latent diffusion for improved memory efficiency and consistency models for accelerated sampling.
+  • Efficiency. By decoupling these processes, we can significantly reduce the inference time (see Figure 7), while achieving comparable performance compared to state-of-the-art algorithms.
+Empirically, we conducted comprehensive experiments validating our method on both linear (super-resolution, inpainting, and image deblurring) and nonlinear inverse problems (nonlinear deblurring) across various datasets including FFHQ [22], LSUN-Bedroom [23], and ImageNet [24], achieving competitive results over existing methods with a faster inference speed.
+Notations.
+Throughout this paper, we denote scalars (functions) with regular lower-case letters (e.g., $s_{t},f(t)$) and vectors (vector functions) with bold lowercase letters (e.g., $\bm{x},\bm{s}(\bm{x}_{t},t)$).
+We use $[N]$ to denote the set $\{1,2,\ldots,N\}$, $p(\cdot)$ to denote the probability, $\mathbb{E}\left[\cdot\right]$ to denote the expectation, $\|\cdot\|_{2}$ to denote the $\ell_{2}$ norm, and we use $\mathcal{N}(\cdot)$ to denote the Gaussian distribution.
 We use $\bm{x}_{t}$ to denote a sample $\bm{x}$ at time step $t$.
-We use $\bm{x}_{0}$ to denote a clean sample, and use $\bm{s}_{\bm{\theta}}$ and $\bm{s}_{\bm{\phi}}$ for the estimated score functions in the pixel and latent spaces, respectively (discussed in detail later).
-2 Preliminaries
-In this section, we first briefly review the basic fundamentals of diffusion models [23, 8, 9, 10, 24, 25] and then discuss how to adapt them for solving image restoration problems.
-Subsequently, we present the main ideas behind diffusion purification, which serve as the backbone of our proposed method.
+We use $\bm{x}_{0}$ to denote a clean sample, and $\bm{s}_{\bm{\theta}}$ and $\bm{s}_{\bm{\phi}}$ for the estimated score functions in the pixel and latent spaces, respectively.
+2 Background
 2.1 Unconditional Diffusion Models
-Diffusion models [23, 8, 9, 10] generate an image with two processes, generally referred to as forward and reverse processes.
-The forward process is to diffuse the data distribution $p_{0}(\bm{x})$ to a standard Gaussian $p_{T}(\bm{x})=\mathcal{N}(\bm{0},\bm{I})$ by gradually injecting noise.
+Diffusion models [25, 5, 6, 7] generate samples by reversing a pre-defined forward process.
+The forward process iteratively diffuses the data distribution $p_{0}(\bm{x})$ with Gaussian noise until a sufficiently large timestep $T$, so that the noise-mollified data distribution $p_{T}(\bm{x})$ becomes indistinguishable from an isotropic Gaussian distribution with covariance $\sigma^{2}_{T}\bm{I}$.
 In the variance-preserving (VP) setting [6], the forward process can be written as a stochastic differential equation (SDE) of the form
 $d\bm{x}_{t}\;=\;-\frac{\beta_{t}}{2}\bm{x}_{t}dt+\sqrt{\beta_{t}}d\bm{w},$ (2)
-where $\beta_{t}>0$ controls the speed of diffusion and $\bm{w}$ follows the standard Wiener process. To sample an image, we can employ the reverse SDE, which takes the form
+where $\beta_{t}>0$ controls the speed of diffusion and $\bm{w}$ follows the standard Wiener process.
+To sample an image, we can start from $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\sigma^{2}_{T}\bm{I})$ and reverse the forward process (2).
+As shown in [6], the forward process (2) can be reversed with either a reverse SDE:
 $d\bm{x}_{t}\;=\;$ $\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\beta_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]dt+\sqrt{\beta_{t}}d\bar{\bm{w}},$ (3)
-where $p_{t}(\bm{x}_{t})$ is the distribution of $\bm{x}_{t}$ at the time $t$, $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the (Stein) score function, and $\bar{\bm{w}}$ follows the reverse Wiener process [27].
-In practice, we estimate the score function using a deep neural network $\bm{s}_{\bm{\theta}}:\mathbb{R}^{n}\times[0,1]\rightarrow\mathbb{R}^{n}$ parameterized by $\bm{\theta}$, which is trained using the denoising score matching objective [28]:
-$\underset{\bm{\theta}}{\min}\,\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{t}|\bm{x}_{0}}\left[\left\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{0t}(\bm{x}_{t}|\bm{x}_{0})\right\|_{2}^{2}\right],$
-where $t\sim\mathcal{U}(0,1)$, $\bm{x}_{0}\sim p_{0}(\bm{x})$, and $\bm{x}_{t}|\bm{x}_{0}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0})$. In practice, both the forward and reverse SDE can be discretized, with
-$\bm{x}_{t}$ $=\sqrt{1-\beta_{t}}\bm{x}_{t-1}+\sqrt{\beta_{t}}\bm{\epsilon}_{t-1},$
-$\bm{x}_{t-1}$ $=\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{x}_{t}+\beta_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right)+\sqrt{\beta_{t}}\bm{\epsilon}_{t},$
-where $\bm{\epsilon}_{t}\sim\mathcal{N}(0,\bm{I})$.
-2.2 Diffusion Models for Image Restoration Problems
-To solve image restoration problems with diffusion models, one of the most commonly used techniques is to modify the unconditional reverse sampling process in Equation 3 by replacing the unconditional score with a conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ based on $\bm{y}$.
+or a reverse ordinary differential equation (ODE):
+$d\bm{x}_{t}\;=\;$ $\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\frac{\beta_{t}}{2}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]dt,$ (4)
+where $p_{t}(\bm{x}_{t})$ is the intermediate noise-mollified distribution at time $t$, $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the score function, and $\bar{\bm{w}}$ is the standard Wiener process [26].
+In practice, the score function is estimated using a deep neural network $\bm{s}_{\bm{\theta}}$ with parameter $\bm{\theta}$ trained using the weighted denoising score matching objective [9]:
+$\underset{\bm{\theta}}{\min}\;\mathbb{E}_{t,\bm{x}_{0},\bm{x}_{t}}\bigl{[}\lambda(t)\|\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)-\nabla_{\bm{x}_{t}}\log p_{0t}(\bm{x}_{t}\mid\bm{x}_{0})\|_{2}^{2}\bigr{]},$
+where $\lambda(t)$ is a positive weighting function, $t\sim\mathcal{U}(0,1)$, $\bm{x}_{0}\sim p_{0}(\bm{x})$, and $\bm{x}_{t}\sim p_{0t}(\bm{x}_{t}|\bm{x}_{0})$, where $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition kernel, i.e., the distribution of $\bm{x}_{t}$ induced by (2) given $\bm{x}_{0}$.
+In the VP setting, the transition kernel takes the following form:
+$p_{0t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\sqrt{\alpha_{t}}\bm{x}_{0},(1-\alpha_{t})\bm{I}),$ (6)
+[…]
+2.2 Diffusion Posterior Sampling for Solving Inverse Problems
+To solve inverse problems using diffusion models, one of the most commonly used techniques is to modify the unconditional reverse sampling process in (3) by replacing the unconditional score with a conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ conditioned on measurement $\bm{y}$.
 As such, the conditional reverse SDE has the following form
-$d\bm{x}_{t}\;=\;\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\beta_{t}\{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})\}\right]dt+\sqrt{\beta_{t}}d\bar{\bm{w}}.$
-Then, by Bayes rule, we can decompose the conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ into
-$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})=\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})+\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}).$
-While the unconditional prior score $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ can be obtained from a pre-trained diffusion model, the likelihood gradient $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})$ is usually intractable.
-To tackle this issue, recent works [16, 17, 18] resort to approximating the likelihood term with tractable surrogates, built upon the idea of diffusion posterior sampling (DPS) [16].
-DPS approximates the likelihood gradient with
-$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})\approx\eta\nabla_{\bm{x}_{t}}\|\mathcal{A}(\hat{\bm{x}}_{0}(\bm{x}_{t}))-\bm{y}\|_{2}^{2},$ (4)
-where $\eta>0$ can be viewed as a step size for controlling the data consistency strength and $\hat{\bm{x}}_{0}(\bm{x}_{t}):=\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ is the posterior mean, which can be estimated through Tweedie’s formula [29, 30]:
-$\hat{\bm{x}}_{0}(\bm{x}_{t})$ $=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t}))$
-$\approx\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)),$
-where $\bar{\alpha}_{T}=\prod_{t=1}^{T}(1-\beta_{t})$.
-However, despite being one of the most recognized algorithms, DPS suffers from several limitations, as previously briefly illustrated.
-Firstly, the inference time of the reverse sampling process is significantly increased, as the likelihood estimation needs to backpropagate over the entire score function $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ for each iteration $t$ of the sampling process.
-Since the score function is generally parameterized as a large deep network, this process can be computationally challenging, especially for tasks involving high-dimensional images.
-Secondly, as the data consistency steps take place within the reverse sampling process, the number of gradient steps for data consistency is limited by the reverse sampling steps, which could lead to insufficient data consistency when using accelerated samplers such as consistency models [19].
-Furthermore, the performance of DPS is sensitive to the step size $\eta$ due to the coupled data consistency and reverse sampling.
-When $\eta$ is small, the reconstructions will not be consistent with the measurements $\bm{y}$.
-On the other hand, a large $\eta$ may push $\bm{x}_{t}$ away from the noisy data manifold on which the diffusion model is trained, resulting in serious artifacts.
-Lastly, naively adapting DPS to latent diffusion models often fails due to the nonlinearity of the decoder.
-Since most diffusion-based image restoration solvers are built upon the idea of DPS, they all suffer from these fundamental issues, motivating the need to decouple the data consistency and reverse sampling processses.
-Figure 2: An illustration of our overall framework.
-Our algorithm involves a total of $K$ iterations which alternate between (i) generating an initial reconstruction by solely optimizing the data fidelity objective and (ii) refining the image via diffusion purification.
+$d\bm{x}_{t}\;=\;\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\beta_{t}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})\right]dt+\sqrt{\beta_{t}}d\bar{\bm{w}}.$ (7)
+Using Bayes rule, the conditional score function $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})$ can be decomposed into the unconditional score and the likelihood gradient:
+$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})=\underbrace{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})}_{\text{unconditional score}}+\underbrace{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})}_{\text{likelihood gradient}}.$ (8)
+While the unconditional score $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ can be approximated using a pre-trained diffusion model $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ (see Section 2.1), the likelihood gradient $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})$ is computationally intractable in general.
+In the past, many existing methods focused on estimating the likelihood gradient under simplified assumptions [27, 14, 15, 16].
+Among them, one of the most popular approaches is diffusion posterior sampling (DPS) [14], which approximates the likelihood gradient as
+$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})\approx\nabla_{\bm{x}_{t}}\|\mathcal{A}(\hat{\bm{x}}_{0}(\bm{x}_{t}))-\bm{y}\|_{2}^{2},$ (9)
+where $\hat{\bm{x}}_{0}(\bm{x}_{t}):=\mathbb{E}[\bm{x}_{0}|\bm{x}_{t}]$ is the posterior mean, which can be computed with the unconditional score function through Tweedie’s formula:
+$\hat{\bm{x}}_{0}(\bm{x}_{t})$ $=\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}+(1-\alpha_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t}))$ (10)
+$\approx\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}+(1-\alpha_{t})\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)).$ (11)
+Note the $\ell_{2}$ norm in (9) is a direct consequence of assuming Gaussian measurement noise; if the noise follows a different distribution, the norm can be replaced accordingly.
+Although effective, the DPS-based method still suffers from several limitations: (i) the number of data consistency steps is limited by the number of reverse sampling steps, potentially resulting in insufficient measurement consistency when employing accelerated samplers, and (ii) computing the approximated likelihood gradient (9) requires backpropagating through the entire score function $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ at each timestep $t$, which is computationally expensive since $\bm{s}_{\bm{\theta}}(\bm{x}_{t},t)$ is typically parameterized by a large deep network.
+Moreover, adapting DPS to latent diffusion models (LDMs) [28] is highly nontrivial due to the nonlinear nature of their encoders and decoders [19, 18].
 2.3 Diffusion Purification
-Recently, diffusion purification has shown to be an effective approach for removing adversarial perturbations [22, 31] and for image editing [32].
-Specifically, for a given noisy or perturbed image $\bm{x}_{\textrm{adv}}$, this process involves first running the diffusion forward process as described in Equation (2) starting from $\bm{x}_{\textrm{adv}}$ to some intermediate timestep $T$.
-In the discrete case, $\bm{x}_{T}$ takes the form
-$\bm{x}_{T}=\sqrt{\bar{\alpha}_{T}}\bm{x}_{\textrm{adv}}+\sqrt{(1-\bar{\alpha}_{T})}\bm{\epsilon}\>.$ (5)
-Starting with $\bm{x}_{T}$, we implement the reverse diffusion process with Equation (3) to produce a “purified” image $\bm{x}_{\textrm{purified}}$, effectively eliminating the adversarial perturbations.
-The rationale behind this is that by carefully selecting $T$, the adversarial noise within $\bm{x}_{\textrm{adv}}$ becomes overwhelmed by the noise introduced during the forward phase, yet the core semantic information is preserved.
+Recently, diffusion purification has been shown to be an effective approach for removing adversarial perturbations [21, 29].
+Specifically, for a given adversarially perturbed image $\bm{x}_{\textrm{adv}}$, diffusion purification first runs the forward process (2) starting from $\bm{x}_{\textrm{adv}}$ to some intermediate timestep $t^{\prime}$ such that $\bm{x}_{t^{\prime}}\sim p_{0t^{\prime}}(\bm{x}_{t^{\prime}}|\bm{x}_{\textrm{adv}})$:
+$\bm{x}_{t^{\prime}}=\sqrt{\alpha_{t^{\prime}}}\bm{x}_{\textrm{adv}}+\sqrt{(1-\alpha_{t^{\prime}})}\bm{\epsilon}\>,$ (12)
+where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$.
+Starting from $\bm{x}_{t^{\prime}}$, diffusion purification performs the reverse diffusion process (3) or (4) to produce a “purified” image $\bm{x}_{\textrm{purified}}$, effectively eliminating the adversarial perturbations.
+The rationale behind this approach is that by carefully selecting $t^{\prime}$, the adversarial noise within $\bm{x}_{\textrm{adv}}$ becomes overwhelmed by the noise introduced during the forward process, yet the core semantic and structural information is preserved.
 Consequently, executing the reverse diffusion process leads to a “clean” image reconstruction.
-In our work, we use diffusion purification as the backbone of our approach, yet with a distinct objective diverging from existing methods that use diffusion purification to enhance the robustness of DNN-based classifiers [22] or remove artifacts in medical imaging [33, 31].
-3 Proposed Methods
-In this section, we first introduce our method for image restoration based upon variable splitting (Section 3.1).
-Then, in Section 3.2, we discuss enforcing data consistency, followed by our proposed method to use diffusion models as an image prior via diffusion purification (Section 3.3).
-The entire procedure of the proposed method is outlined in Algorithm 1.
-3.1 Methods with Variable Splitting
-Our proposed algorithm is inspired by the classical variable splitting technique [34, 35, 36, 37] in optimization.
-Specifically, consider the following regularized optimization formulation for the image restoration problem in Equation 1,
-$\min_{\bm{x}}\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{x}),$
-where $\mathcal{R}(\cdot)$ is the regularization for image prior, and $\lambda>0$ controls the penalty strength.
-The idea of variable splitting is to introduce an auxiliary variable $\bm{v}=\bm{x}$ to pose an equivalent problem
-$\min_{\bm{x},\bm{v}}\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\lambda\mathcal{R}(\bm{v}),\;\mathrm{s.t.}\;\bm{x}=\bm{v}.$ (6)
-This constrained optimization problem can be relaxed into an unconstrained problem by the half quadratic splitting technique (HQS) [38, 37], which we can reformulate as
-$\min_{\bm{x},\bm{v}}\underbrace{\frac{1}{2}\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}+\frac{\mu}{2}\|\bm{x}-\bm{v}\|_{2}^{2}+\lambda\mathcal{R}(\bm{v})\>,}_{=:\mathcal{L}(\bm{x},\bm{v})}$ (7)
+Distinct from existing methods that use diffusion purification to improve the robustness of DNN-based classifiers [21], in our work, we use diffusion purification as an operator that enforces the image prior embedded in the diffusion models.
+3 Proposed Method
+This section first motivates DCDP from the perspective of classical variable splitting techniques (section 3.1).
+Then we present our overall algorithm (section 3.2) and its extension to latent diffusion models and consistency models (sections 3.3 and 3.4).
+3.1 Motivation: Methods with Variable Splitting
+Our method is inspired by the classical variable splitting techniques [30, 31, 32, 33] in optimization. Consider the following regularized optimization formulation for solving (1),
+$\min_{\bm{x}}f(\bm{x})+\lambda\mathcal{R}(\bm{x}),$
+where we typically adopt the least squares formulation, $f(\bm{x}):=\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}$, as the data fidelity term to enforce consistency with the measurements, $\mathcal{R}(\cdot)$ is the regularizer that enforces the data prior, and $\lambda>0$ is a hyperparameter that balances the fidelity and the regularization strength.
+The idea behind variable splitting is to introduce an auxiliary variable $\bm{v}=\bm{x}$, resulting in an equivalent problem
+$\min_{\bm{x},\bm{v}}f(\bm{x})+\lambda\mathcal{R}(\bm{v}),\quad\mathrm{s.t.}~{}~{}\;\bm{x}=\bm{v},$ (13)
+where the constrained optimization problem above can be further relaxed into an unconstrained problem by the half quadratic splitting technique (HQS) [34, 33]:
+$\min_{\bm{x},\bm{v}}\mathcal{L}(\bm{x},\bm{v}):=f(\bm{x})+\lambda\,\mathcal{R}(\bm{v})+\mu\|\bm{x}-\bm{v}\|_{2}^{2},$ (14)
 for some properly chosen $\mu>0$.
-Then, to solve (7), we can use an iterative method such that for every iteration $k$, we can leverage the extra variable $\bm{v}$ to split the optimization into two sub-problems of the form
-$\bm{x}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{x}}\;\mathcal{L}(\bm{x},\bm{v}_{k-1}),$ (8)
-$\bm{v}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{v}}\;\mathcal{L}(\bm{x}_{k},\bm{v}),$ (9)
-that decouples the data consistency term $\left\|\mathcal{A}(\bm{x})-\bm{y}\right\|_{2}^{2}$ from the regularization term $\mathcal{R}(\bm{v})$.
-Then, instead of having an explicit regularizer $\mathcal{R(\cdot)}$ (e.g., $\ell_{1}$-norm to induce sparsity), we can use implicit regularizers.
-More specifically, we can constrain the output space as the range of the diffusion model via diffusion purification that acts as a form of regularization or prior.
-To enforce this implicit regularizer, our method involves alternating between the two following steps:
-  1. 1.
-    Data Consistency. Generate measurement-consistent reconstructions, which may have various artifacts, by solving the data consistency optimization problem Equation 8.
-  2. 2.
-    Regularization via Diffusion Purification. Refine the obtained reconstruction from the previous step with a pre-trained diffusion model via diffusion purification.
-In the following subsections, we dive into the technical details of each step and present an illustration of our method in Figure 2.
-We also conduct an ablation study in Section 5.3, demonstrating that alternating between these two steps can decrease the distance between the reconstructed and the target image.
-Algorithm 1 Image Restoration with Decoupled Data Consistency and Diffusion Model Prior
-measurements $\bm{y}$, forward operator $\mathcal{A}(\cdot)$, pretrained model $\bm{s}_{\bm{\theta}}$, total number of iterations $K$, decaying time schedule $\{T_{k}\}_{k\in[K]}$ for diffusion purification, iteration number $\tau$, and $\bm{v}_{0}=\bm{0}\in\mathbb{R}^{n}$
+To solve (14), a common strategy is to adopt an iterative method that, at each iteration $k$, uses the auxiliary variable $\bm{v}$ to decompose the optimization into two subproblems of the form:
+$\bm{x}_{k}=\arg\min_{\bm{x}}\;\mathcal{L}(\bm{x},\bm{v}_{k-1}),$ (15)
+$\bm{v}_{k}=\arg\min_{\bm{v}}\;\mathcal{L}(\bm{x}_{k},\bm{v}).$ (16)
+In this way, the data-consistency term $f(\bm{x})$ is decoupled from the regularization term $\mathcal{R}(\bm{v})$.
+Then, instead of having an explicit regularizer $\mathcal{R(\cdot)}$ (e.g., $\ell_{1}$-norm to induce sparsity), the regularization-based step (16) can be replaced with operators that implicitly enforce the data prior.
+In this work, we leverage diffusion purification to enforce the image prior encoded within the diffusion models as we discuss in Section 3.2.
+3.2 Decoupled Data Consistency via Diffusion Purification
+Our DCDP method alternates between the following two phases: (i) a reconstruction phase that enforces data consistency and (ii) a refinement phase that enforces the prior via diffusion purification.
+Below, we discuss these steps in more detail, with an illustration in Figure 1.
+Reconstruction Phase.
+The first step of our algorithm is to solve the sub-problem (15) to enforce data consistency by minimizing the data fidelity term $f(\bm{x})$.
+Given $\bm{v}_{k-1}$ from the previous iteration, optimizing the loss $\mathcal{L}(\bm{x},\bm{v}_{k-1})$ for $\bm{x}$ depends only on the data fidelity term $f(\bm{x})=\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}$ and its associated proximal term $\mu\|\bm{x}-\bm{v}_{k-1}\|_{2}^{2}$, leading to:
+$\bm{x}_{k}=\arg\min_{\bm{x}}\;\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}+\mu\|\bm{x}-\bm{v}_{k-1}\|_{2}^{2}\>.$ (17)
+Although (17) can be solved via iterative methods (e.g., SGD), selecting an appropriate hyperparameter $\mu$ often requires manual tuning.
+To mitigate this issue, we compute an approximate solution $\bm{x}_{k}$ by minimizing only the data consistency term $f(\bm{x})$, initialized at $\bm{v}_{k-1}$ and run for a fixed number of gradient steps.
+Empirically, this approach (i) yields a solution $\bm{x}_{k}$ that stays close to $\bm{v}_{k-1}$ without relying on $\mu$, and (ii) maintains low computational cost while being less sensitive to the hyperparameter.
+Note that enforcing data consistency with (17) is significantly more efficient than DPS-based methods, as it does not require backpropagation over the entire score function.
+In practice, we can use any gradient-based optimizer (e.g., SGD, ADAM) for minimizing (17).
+Throughout the experiments, we fix the number of gradient steps $\tau$ for each round of the data consistency optimization.
+When the measurements are corrupted by noise, one can reduce $\tau$ to prevent overfitting.
+We discuss our choice of optimizers in detail in Appendix B.
+For simplicity, we denote this reconstruction phase by:
+$\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau).$ (18)
+Figure 2: Demonstrating the regularization effects of diffusion purification.
+The estimates from the data fidelity step contain artifacts, making the overall reconstruction inaccurate, whereas regularization via diffusion purification can remove such artifacts.
+Refinement Phase.
+Due to the ill-posed nature of the problem in (1), $\bm{x}_{k}$ obtained in the reconstruction phase could suffer from task-specific artifacts, as shown in Figure 2 (third column).
+To alleviate these artifacts from $\bm{x}_{k}$, we propose to use diffusion purification described in Section 2.3 as an implicit regularizer.
+Specifically, let $t_{k}$ denote a predefined diffusion purification timestep that controls the strength of purification at iteration $k$.
+Following (12), we first define $\bm{x}_{k,0}:=\bm{x}_{k}$ and submerge this sample with noise by running the forward process up to time $t_{k}$:
+$\bm{x}_{k,t_{k}}=\sqrt{\alpha_{t_{k}}}\bm{x}_{k,0}+\sqrt{1-\alpha_{t_{k}}}\bm{\epsilon}.$ (19)
+Then, we run unconditional reverse sampling via either the reverse SDE (3) or the reverse ODE in (4) starting from the initial point $\bm{x}_{k,t_{k}}$ at the time $t_{k}$ to obtain the purified output $\bm{v}_{k}$.
+As demonstrated in Figure 2, diffusion purification effectively removes the artifacts in the reconstructions from the data consistency (fidelity) optimization (17) while preserving the overall image structures at the same time.
+For simplicity, we denote this refinement step by:
+$\bm{v}_{k}=\texttt{DPUR}(\bm{x}_{k},\bm{s}_{\bm{\theta}},t_{k}).$ (20)
+For the choice of $t_{k}$, we propose to use a linear decaying schedule for robust performance, which is capable of addressing the “hallucination” issue brought by the stochasticity of the diffusion purification process.
+In the early iterations of our method, when $t_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though it may be inconsistent with the measurements, will be free from artifacts and can serve as a good starting point for the subsequent iteration of data consistency optimization.
+As $t_{k}$ keeps decreasing when $k$ increases, the algorithm will converge to an equilibrium point which is both consistent with the measurements and free from the artifacts.
+The overall algorithm is presented in Algorithm 1.
+More technical details are deferred to Appendix D, and we empirically study the convergence in Section 4.3.
+Algorithm 1 DCDP using Pixel Space Diffuison Models
+measurements $\bm{y}$, forward operator $\mathcal{A}(\cdot)$, pretrained diffusion model $\bm{s}_{\bm{\theta}}$, total number of reverse sampling steps $K$, decaying time schedule $\{t_{k}\}_{k\in[K]}$ for diffusion purification, the number of data consistency iterations $\tau$, and $\bm{v}_{0}=\bm{0}\in\mathbb{R}^{n}$
 for $k=1,\dots,K$ do
-     $\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau)$triangleright solve via gradient based methods
+     $\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau)$triangleright solve via gradient based methods for $\tau$ steps
      $\bm{v}_{k}=\texttt{DPUR}(\bm{x}_{k},\bm{s}_{\bm{\theta}},t_{k})$triangleright Diffusion Purification
@@
 Return $\bm{v}_{k}$
-3.2 Enforcing Data Consistency
-Recall that the sub-problem in Equation (8) for maintaining data consistency can be written as
-$\bm{x}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{x}}\;\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}+\mu\|\bm{x}-\bm{v}_{k-1}\|_{2}^{2}\>.$
-The difficulty behind directly solving the above problem lies in selecting the proper hyperparameter $\mu$, as it may need to be frequently tuned.
-To tackle this issue, we instead seek an approximate solution $\bm{x}_{k}$ that minimizes solely the data fidelity term $\frac{1}{2}\|\mathcal{A}(\bm{x})-\bm{y}\|_{2}^{2}$ but starting the optimization with $\bm{v}_{k-1}$ as the initial point and run a limited number of $\tau$ gradient steps, resulting in a measurement-consistent reconstruction $\bm{x}_{k}$.
-The rationale behind this is two-fold: (i) by taking $\bm{v}_{k-1}$ as the initial point, we can obtain an output $\bm{x}_{k}$ that is sufficiently close to $\bm{v}_{k-1}$ without having to tune $\mu$, and (ii) taking a limited number of gradient steps $\tau$ reduces the computational burden.
-Overall, for simplicity, we denote this step as
-$\bm{x}_{k}=\texttt{Data-Fidelity}(\bm{v}_{k-1},\bm{y},\mathcal{A},\tau).$ (10)
-Lastly, we want to highlight that our technique for enforcing data consistency is significantly more efficient than existing methods, as our method does not require back propagation over the entire score function (i.e., the neural network) and hence is (very) computationally efficient.
-Furthermore, to enforce data consistency, we can use any momentum-based gradient descent optimizer (e.g., SGD, ADAM) for Equation 10.
-Throughout the experiment we fix the number of gradient steps $\tau$ for each round of the data fidelity optimization.
-When the measurements are corrupted by noise, one can further reduce the number of iteration to prevent overfitting.
-We discuss our choice of optimizers in Appendix A.
-Figure 3: Demonstrating the regularization effects of diffusion purification.
-The estimates from the data fidelity step contain artifacts, making the overall reconstruction inaccurate, whereas regularization via diffusion purification can remove such artifacts.
-3.3 Implicit Regularization by Diffusion Purification
-Due to the underdetermined nature of the problem (1), the initial reconstruction $\bm{x}_{k}$ obtained from Equation 10 will suffer from task specific artifacts, as shown in Figure 3 (the third column).
-The next step is to refine $\bm{x}_{k}$ via enforcing the image prior learned by the diffusion model.
-To do so, we propose to utilize diffusion purification described in Section 2.3 as an implicit regularizer for effectively removing the artifacts in $\bm{x}_{k}$ as shown in Figure 3.
-More specifically, let $T_{k}$ denote the diffusion purification strength at iteration $k$.
-Following the diffusion purification step in Equation (5), we first define $\bm{x}_{k,0}:=\bm{x}_{k}$ and submerge this sample with noise up to time $T_{k}$:
-$\bm{x}_{k,T_{k}}=\sqrt{\bar{\alpha}_{T_{k}}}\bm{x}_{k,0}+\sqrt{1-\bar{\alpha}_{T_{k}}}\bm{\epsilon}.$ (11)
-Then, we run unconditional reverse sampling via the reverse SDE in Equation 3 starting from $\bm{x}_{k,T_{k}}$ to obtain the output $\bm{v}_{k}$.
-Concretely, the discretized version of the reverse SDE can be described as
-$\bm{x}_{k,t-1}=$ $\frac{1}{\sqrt{1-\beta_{t}}}\left(\bm{x}_{k,t}+\beta_{t}\bm{s}_{\bm{\theta}}(\bm{x}_{k,t},t)\right)+\sqrt{\beta_{t}}\bm{\epsilon}_{t},$ (12)
-for $t=T_{k},T_{k}-1,\ldots,1$.
-Here, recall that $\bm{s}_{\bm{\theta}}(\bm{x},t)$ denotes the score function which can be estimated by a pre-trained unconditional diffusion model.
-For simplicity, we denote this overall regularization step by
-$\bm{v}_{k}=\texttt{DPUR}(\bm{x}_{k},\bm{s}_{\bm{\theta}},T_{k}).$ (13)
-For the choice of $T_{k}$, we propose to use a linear decaying schedule for robust performance, which is capable of addressing the “hallucination” issue brought by the stochasticity of the diffusion purification process.
-In the early iterations of our method, when $T_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though may be inconsistent with the measurements, will be free from artifacts and can serve as good starting point for the subsequent iteration of data consistency optimization.
-As $T_{k}$ keeps decreasing, the algorithm will converge to an equilibrium point which is both consistent with the measurements and free from the artifacts.
-More technical details are deferred to Appendix A, and we study the convergence in Section 5.3.
-4 Adaptability to Accelerated Samplers & Latent Diffusion
-One of the main drawbacks of diffusion models is the increased inference time of the iterative sampling process compared to traditional methods such as variational autoencoders (VAEs) [39] and generative adversarial networks (GANs) [40].
-In the literature, many advancements have been made to address this issue, which involve either decreasing the number of iterations via accelerated sampling techniques such as DDIMs [11] or carrying out the sampling process in a lower-dimensional, latent space [24, 25].
-In the following sections, we briefly introduce these techniques and demonstrate how they can be used to decrease the inference time of diffusion purification, thereby accelerating our algorithm.
-4.1 Faster Diffusion Purification
-To improve sampling efficiency, we can adopt the use of accelerated samplers to decrease the sampling time of the reverse process of diffusion purification.
-For example, we can use the accelerated sampler DDIM, which constructs a non-Markovian chain of the reverse process.
-By doing so, we can perform each round of diffusion purification with as few as $20$ unconditional reverse sampling steps (i.e., $20$ number of function evaluations (NFEs)), with minimal performance degradation.
-Here, we would like to remark that existing methods such as DPS do not have this advantage.
-DPS typically requires approximately $1000$ reverse sampling steps, as it requires more iterations to ensure “sufficient” data consistency steps.
-In the following, we discuss the adaptability of our approach using one-step sampling approaches, including consistency models and approximations via Tweedie’s formula.
-Consistency Models.
-Recently, Song et al.
-[19] proposed a class of diffusion models termed consistency models (CMs), which can effectively generate high-quality samples in a single (or at least very few) time steps.
-CMs build on the fact that there exists a deterministic ODE for every corresponding SDE sampling trajectory of Equation 3 in the form of
-$d\bm{x}_{t}\;=\;\left[-\frac{\beta_{t}}{2}{\bm{x}_{t}}-\frac{\beta_{t}}{2}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]dt,$ (14)
-with which it shares the same probability densities $p_{t}(\bm{x}_{t})$ for all $t\in[0,1]$ [9].
-This implies that for any image $\bm{x}_{0}$, there exists a unique ODE trajectory $\{\bm{x}_{t}\}_{t=0}^{1}$.
-By leveraging this property, CMs train a deep network $f_{\bm{\theta}}:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ parameterized by $\bm{\theta}$ to map any point $\bm{x}_{t}$ to its corresponding $\bm{x}_{0}$ in a single step.
-This approach can be seamlessly incorporated into our framework.
-Specifically, with a pre-trained consistency model $f_{\bm{\theta}}$, the diffusion purification step in Equation 13 in Algorithm 1 can be performed in a single step as
-$\bm{v}_{k}=\texttt{DPUR}(\bm{f}_{\bm{\theta}},\bm{x}_{k},T_{k}):=f_{\bm{\theta}}(\bm{x}_{k},T_{k}).$ (15)
-Approximation via Tweedie’s Formula.
-Under a Gaussian assumption, Tweedie’s formula [29, 30] provides a way of computing the posterior mean by using
-$\mathbb{E}\left[\bm{x}_{0}\mid\bm{x}_{t}\right]=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\bm{x}_{t}+(1-\bar{\alpha}_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})),$
+3.3 Fast Diffusion Purification for Refinement
+In this subsection, we demonstrate two approaches to achieve faster diffusion purification: (i) by using accelerated samplers and (ii) by using Tweedie’s formula.
+Accelerated Samplers.
+To improve sampling efficiency, we can adapt accelerated samplers to speed up the reverse process of diffusion purification.
+For example, we can use DDIMs [8], which constructs a non-Markovian chain of the reverse process.
+By doing so, each round of diffusion purification can be performed with as few as $20$ unconditional reverse sampling steps (number of function evaluations, NFEs), with minimal performance degradation.
+We remark that existing methods such as DPS do not offer this advantage, as they typically require over $1000$ reverse sampling steps to achieve sufficient data consistency.
+Approximation via Tweedie’s Formula. By Tweedie’s formula [35, 36], the posterior mean given noisy sample $\bm{x}_{t}$ can be estimated with the score function:
+$\mathbb{E}\left[\bm{x}_{0}\mid\bm{x}_{t}\right]=\frac{1}{\sqrt{\alpha_{t}}}(\bm{x}_{t}+(1-\alpha_{t})\nabla_{\bm{x}_{t}}\log p(\bm{x}_{t})),$ (21)
 where $\bm{x}_{0}$ denotes a sample from the “clean” data distribution.
-Equipped with this formula, instead of running the reverse SDE in Equation 12 for diffusion purification, one can estimate the ground truth in a single step using this formula.
-Specifically, we can approximate $\bm{v}_{k}$ using
-$\bm{v}_{k}$ $=\mathbb{E}\left[\bm{x}_{k,0}\mid\bm{x}_{k,T_{k}}\right]\approx\frac{1}{\sqrt{\bar{\alpha}_{T_{k}}}}(\bm{x}_{k,T_{k}}+(1-\bar{\alpha}_{T_{k}})\bm{s}_{\bm{\theta}}(\bm{x}_{k,T_{k}},T_{k}))\>,$
-where the approximation arises from the score function estimation and $\bm{x}_{k,T_{k}}$ is a sample at iteration $k$ diffused upto switching time $T_{k}$ via the forward process.
-Compared to reverse SDE sampling, employing Tweedie’s formula will lead to a tradeoff between computation time and reconstruction quality.
-As evidenced by our experimental results in Section 5, using Tweedie’s formula for approximating the reverse SDE may result in blurry reconstructions with lower perceptual quality, a phenomenon partly attributed to the ”regression to the mean” effects [41].
-Algorithm 2 LDM Version of Our Method
+Equipped with this formula, one can use the posterior mean as a single-step approximation to the solution of the reverse SDE (or ODE) in diffusion purification:
+$\bm{v}_{k}$ $\approx\mathbb{E}\left[\bm{x}_{k,0}\mid\bm{x}_{k,t_{k}}\right]$ (22)
+$\approx\frac{1}{\sqrt{\alpha_{t_{k}}}}(\bm{x}_{k,t_{k}}+(1-\alpha_{t_{k}})\bm{s}_{\bm{\theta}}(\bm{x}_{k,t_{k}},t_{k}))\>.$ (23)
+Note that applying Tweedie’s formula leads to a trade-off between computation time and reconstruction quality.
+The multi-step reverse SDE produces a sharp sample that approximately follows the ground truth data distribution, whereas Tweedie’s formula returns the mean of all possible clean versions corresponding to the noisy sample $\bm{x}_{t}$, which is perceptually blurrier especially when $t$ is large.
+As evidenced by our experimental results in Section 4, using Tweedie’s formula may result in blurry reconstructions with lower perceptual quality.
+Nevertheless, DCDP with Tweedie’s formula still achieves competitive performance on standard metrics and is extremely fast as it requires only one step.
+3.4 DCDP with Latent Diffusion and Consistency Models
+By decoupling the data consistency steps from the diffusion sampling steps, our method enables the flexibility of utilizing a broad range of diffusion models in a plug-and-play fashion.
+To this end, we illustrate how to adapt DCDP to latent diffusion models and consistency models.
+DCDP with Latent Diffusion Models (LDMs).
+LDMs [28] operate in a lower-dimensional latent space by training a pair of encoder and decoder [37] to improve the efficiency of vanilla diffusion models.
+Specifically, given an encoder $\mathcal{E}:\mathbb{R}^{d}\to\mathbb{R}^{r}$ and a decoder $\mathcal{D}:\mathbb{R}^{r}\to\mathbb{R}^{d}$ where $r\ll d$, LDMs reduce the computational complexity by mapping each $\bm{x}\in\mathbb{R}^{d}$ to $\bm{z}=\mathcal{E}(\bm{x})\in\mathbb{R}^{r}$ and performing the entire diffusion process in the latent space.
+After performing the reverse steps in the latent space, the sampled latent code $\bm{z}_{0}$ is decoded to a synthesized image via $\bm{x}_{0}=\mathcal{D}(\bm{z}_{0})$.
+Unlike previous methods [18, 19], the flexibility of our framework allows us to (almost) seamlessly incorporate LDMs.
+For LDMs, we alternate between enforcing data consistency and diffusion purification, but both in the latent space:
+$\bm{z}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{z}}\;\mathcal{L}_{\mathrm{LDM}}(\bm{z},\widehat{\bm{v}}_{k-1}),$ (24)
+$\widehat{\bm{v}}_{k}$ $=\texttt{DPUR}(\bm{z}_{k},\bm{s}_{\bm{\phi}},t_{k}),$ (25)
+where $\mathcal{L}_{\mathrm{LDM}}$ is a loss function tailored for the latent space (see (26) and (29)), $\bm{s}_{\bm{\phi}}$ is the score function in the latent space, and $\bm{z}=\mathcal{E}(\bm{x})$ is the latent code.
+Once our method converges based on specific stopping criteria, we map the converged latent code back to the pixel space through the decoder $\mathcal{D}$.
+The overall method is presented in Algorithm 2.
+In the following subsections, we discuss two approaches for enforcing data consistency.
+Approach I. Data Consistency in Latent Space
+One approach for enforcing data consistency is to operate directly in the latent space by minimizing the following objective:
+$\bm{z}_{k}=\operatorname*{\text{arg~{}min}}_{\bm{z}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}+\mu\|\bm{z}-\widehat{\bm{v}}_{k-1}\|_{2}^{2}$ (26)
+where $\widehat{\bm{v}}_{k-1}$ denotes the output from the previous diffusion purification step.
+We note that this optimization objective closely resembles the method in [4].
+As discussed in Section 3.2, we approximately solve (26) by performing a limited number of gradient-based updates on the objective $\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}$, initialized from $\widehat{\bm{v}}_{k-1}$ as:
+$\bm{z}_{k}=\texttt{Data-Fidelity-LDM}(\widehat{\bm{v}}_{k-1},\bm{y},\mathcal{A},\mathcal{D},\tau).$ (27)
+Before diffusion purification, we run an extra step:
+$\bm{z}_{k}$ $\leftarrow\mathcal{E}(\mathcal{D}(\bm{z}_{k})),$ (28)
+which we denote by “Re-Encode”.
+The rationale behind Re-Encode is that due to the nonlinearity of the encoder $\mathcal{E}$, there could exist infinitely many solutions $\bm{z}$ for the optimization problem (27).
+However, since the LDMs are trained only on $\bm{z}=\mathcal{E}(\bm{x})$, where $\bm{x}$ follows the image distribution, we need to first map $\bm{z}_{k}$ back to the image space with $\bm{x}_{k}=\mathcal{D}(\bm{z}_{k})$ and then re-encode $\bm{x}_{k}$ as $\bm{z}_{k}=\mathcal{E}(\bm{x}_{k})$.
+Algorithm 2 LDM Version of DCDP
 measurements $\bm{y}$, forward operator $\mathcal{A}(\cdot)$, pretrained LDM $\bm{s}_{\bm{\phi}}$, a pair of pretrained encoder $\mathcal{E}$ and decoder $\mathcal{D}$, total number of iterations $K$, decaying time schedule $\{t_{k}\}_{k\in[K]}$ for diffusion purification, iteration number $\tau$, and $\widehat{\bm{v}}_{0}\sim\mathcal{N}(\bm{0},\bm{I})$ if enforcing data consistency in the latent space or $\mathcal{D}(\widehat{\bm{v}}_{0})=\bm{0}$ if enforcing data consistency in the pixel space.
@@
 Return $\mathcal{D}(\bm{v}_{k})$
-Lastly, notice that this approximation may seem similar to CMs, as they both necessitate only a single NFE.
-However, CMs avoid some of the tradeoffs discussed previously, as CMs were designed specifically for single-step sampling.
-Nevertheless, we illustrate the effectiveness of our framework on both techniques, demonstrating that one can use any model that is more readily available.
-4.2 Adaptation to Latent Diffusion
-Latent diffusion models (LDMs) [24, 25] operate in a lower-dimensional latent space by training a pair of encoder and decoder [42] to address the slow inference time of diffusion models.
-Specifically, given an encoder $\mathcal{E}:\mathbb{R}^{d}\to\mathbb{R}^{r}$ and a decoder $\mathcal{D}:\mathbb{R}^{r}\to\mathbb{R}^{d}$ where $r\ll d$, LDMs reduce the computational complexity by mapping each $\bm{x}\in\mathbb{R}^{d}$ to $\bm{z}=\mathcal{E}(\bm{x})\in\mathbb{R}^{r}$ and performing the entire diffusion process in the latent space.
-Then, upon running the reverse steps within the latent space, one can pass the sampled latent code $\bm{z}_{0}$ through the decoder to obtain a synthesized image $\bm{x}_{0}=\mathcal{D}(\bm{z}_{0})$.
-Unlike previous methods [20, 21], the flexibility of our framework allows us to (almost) seamlessly incorporate LDMs.
-Similar to Section 3.1, we alternate between enforcing data consistency and diffusion purification but both in the latent space:
-$\bm{z}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{z}}\;\mathcal{L}_{\mathrm{LDM}}(\bm{z},\widehat{\bm{v}}_{k-1}),$ (16)
-$\widehat{\bm{v}}_{k}$ $=\texttt{DPUR}(\bm{z}_{k},\bm{s}_{\bm{\phi}},T_{k}),$ (17)
+Approach II. Data Consistency in Pixel Space
+Alternatively, we can directly enforce the data consistency in the pixel space.
+Specifically, given $\widehat{\bm{v}}_{k-1}$ as the outcome of the previous diffusion purification step, we can optimize the following objective:
+$\bm{x}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{x}}\frac{1}{2}||\bm{y}-\mathcal{A}(\bm{x})||_{2}^{2}+\mu\|\bm{x}-\mathcal{D}(\widehat{\bm{v}}_{k-1})\|_{2}^{2}.$ (29)
+Similar to that in the pixel space, (29) can be approximately solved by taking a limited number of gradient-based steps on $\frac{1}{2}||\bm{y}-\mathcal{A}(\bm{x})||_{2}^{2}$ when initialized from $\mathcal{D}(\widehat{\bm{v}}_{k-1})$, with
 […]
-Approach I: Enforcing DC in the Latent Space.
-One way of enforcing data consistency (DC) is strictly in the latent space. Specfically, we can consider the modified problem
-$\bm{z}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{z}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}+\frac{\mu}{2}\|\bm{z}-\widehat{\bm{v}}_{k-1}\|_{2}^{2},$
-where $\widehat{\bm{v}}_{k-1}$ is the outcome of the previous diffusion purification step in the latent space. We remark that this problem is similar to the CSGM problem considered in [4].
-Then, similar to Section 3.2, we can approximately solve the above problem by taking a limited number of gradient based steps on $\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}$ w.r.t.
-$\bm{z}$ starting from $\widehat{\bm{v}}_{k-1}$:
-$\bm{z}_{k}=\texttt{Data-Fidelity-LDM}(\widehat{\bm{v}}_{k-1},\bm{y},\mathcal{A},\mathcal{D},\tau).$ (18)
-Before the diffusion purification step, we run an extra step on the solution obtained from Equation 18
-$\bm{z}_{k}$ $\leftarrow\mathcal{E}(\mathcal{D}(\bm{z}_{k})),$ (19)
-which we label “Re-Encode”.
-The rationale behind this seemingly unnecessary step is that due to the nonlinearity of the encoder $\mathcal{E}$, there exist infinitely many solutions $\bm{z}$ for the optimization problem in Equation 18.
-However, since the LDMs are trained only on $\bm{z}=\mathcal{E}(x)$, we need to first map $\bm{z}_{k}$ back to the image space with $\bm{x}_{k}=\mathcal{D}(\bm{z}_{k})$ and then re-encode $x_{k}$ as $z_{k}=\mathcal{E}(\bm{x}_{k})$.
-Approach II: Enforcing DC in the Pixel Space.
-Alternatively, we can also directly enforce the data consistency in the pixel space.
-Specifically, given $\widehat{\bm{v}}_{k-1}$ as the outcome of the previous diffusion purification step in the latent space, we can run
-$\bm{x}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{x}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\bm{x})\|_{2}^{2}+\frac{\mu}{2}\|\bm{x}-\mathcal{D}(\widehat{\bm{v}}_{k-1})\|_{2}^{2}.$
-Again, we can approximately solve the above problem by taking a limited number of gradient based steps on $\frac{1}{2}||\bm{y}-\mathcal{A}(\bm{x})||_{2}^{2}$ w.r.t.
-$\bm{x}$ starting from $\mathcal{D}(\widehat{\bm{v}}_{k-1})$:
-[…]
-which is the same as Equation 10 with $\bm{v}_{k-1}=\mathcal{D}(\widehat{\bm{v}}_{k-1})$. Followed by this, we run
-$\bm{z}_{k}$ $=\mathcal{E}(\bm{x}_{k}),$
-to use the encoder $\mathcal{E}$ to map $\bm{x}_{k}$ into the latent space $\bm{z}_{k}$ for the subsequent diffusion purification step in Equation 17.
-Compared to Approach I, it is worth noting that Approach II is much more efficient.
-This is because here we do not require back propogation through the deep decoder $\mathcal{D}$ for computing the gradient.
-In practice, both methods are effective for a range of image restoration problems.
-However, we observe that Approach II tends to be more suitable for specific types of problems like motion deblurring and Gaussian deblurring.
-This is because it has a better capability to produce $\bm{x}_{k}$’s with superior quality compared to Approach I.
-See Figures 15 and 14 in Section A.1 for more detail.
+which is equivalent to (18) with $\bm{v}_{k-1}=\mathcal{D}(\widehat{\bm{v}}_{k-1})$.
+Followed by this, we run $\bm{z}_{k}=\mathcal{E}(\bm{x}_{k})$ to map $\bm{x}_{k}$ into the latent space for the subsequent diffusion purification step in (25).
+Compared to Approach I, it is worth noting that Approach-II is much more efficient since it does not require back-propagation through the deep decoder $\mathcal{D}$ when computing the gradient.
+In practice, both methods are effective for a range of inverse problems.
+However, we find that Approach II is better suited for tasks like motion deblurring and Gaussian deblurring, as it consistently yields higher-quality reconstructions than Approach I (See Figures 16 and 15 in Appendix A).
+DCDP with Consistency Models.
+Recently, Song et al.
+[17] proposed a class of diffusion models termed consistency models (CMs), which can effectively generate high-quality samples in a single step.
+CMs can be seamlessly incorporated into our framework.
+Specifically, with a pre-trained consistency model $\bm{f}_{\bm{\theta}}$, which maps a noisy sample $\mathbf{x}_{k}$ at time $t_{k}$ directly to its denoised estimate, the diffusion purification step in (20) can be performed in a single step as
+$\bm{v}_{k}=\texttt{DPUR}(\bm{f}_{\bm{\theta}},\bm{x}_{k},t_{k}):=\bm{f}_{\bm{\theta}}(\bm{x}_{k},t_{k}).$ (31)
+Performing diffusion purification with CMs may seem similar to the approximation via Tweedie’s formula, as they both require only a single NFE.
+However, CMs alleviate the blurriness issue brought by the Tweedie’s formula discussed in Section 3.3.
+This is because unlike the Tweedie’s formula that estimates the posterior mean, CMs are designed specifically for sampling from the data distribution with a single step.
+We illustrate the effectiveness of our framework on both techniques in Section 4.2, demonstrating that one can use any model that is readily available.
+3.5 Comparison with Related Works
+Before demonstrating the effectiveness of our approach in Section 4, we first discuss how it relates to recent closely related works.
+For a more comprehensive overview of other diffusion model-based methods, such as [38], we refer readers to the recent survey by Daras et al.
+[39] and to Section 2 of Zheng et al.
+[40].
+Plug-and-Play Diffusion Priors.
+Our work, together with several concurrent studies [41, 42, 43], falls under the category of plug-and-play diffusion-based inverse problem solvers.
+These methods alternate between a data fidelity step and a prior step during inference.
+While both our approach and [41, 42] use implicit priors of diffusion denoisers, the latter directly performs denoising on $\bm{x}_{k}$ by assuming $\bm{x}_{k}$ contains a certain amount of pseudo noise, which might be a strong assumption in practice.
+Our method explicitly injects noise to $\bm{x}_{k}$ before performing denoising with either iterative sampling or one-step using Tweedie’s formula.
+This explicit noise-then-denoise process—termed diffusion purification in our work—is also implicitly adopted in DAPS [43].
+The main distinction lies in the data fidelity step: DAPS uses Langevin dynamics for posterior sampling, whereas our formulation in (17) can be viewed as a maximum a posterior (MAP) estimation.
+Hence, the two algorithms can be seen as complementary.
+As they were developed concurrently, we leave a thorough comparison of the two approaches for future work.
+Lastly, the study presented by Luther et al.
+[44] shares close ties with DCDP using Tweedie’s formula in the pixel space.
+Their algorithm, Diffusive Denoising of Gradient-based Minimization (DDGM), is designed for solving linear tomographic reconstruction problems, which iteratively applies data consistency steps and image denoising through a pre-trained diffusion model.
+However, they only investigate the use of Tweedie’s formula for regularizing the images, which often results in blurry images (see Section 4.2).
+Our algorithm circumvents this issue via DDIM which performs diffusion purification with multiple diffusion reverse sampling steps.
+Furthermore, our framework can incorporate accelerated samplers, latent diffusion models, and consistency models.
+Consequently, it addresses a broader spectrum beyond tomographic reconstruction and can be readily adapted to a wide range of applications, unlike DDGM.
+Hard Data Consistency Methods.
+Another class of approaches including DiffPIR [33] and ReSample [19] improves upon DPS by alternating between two key steps: (i) at each time step $t$, rather than performing a single-step data consistency update as in DPS, DiffPIR and ReSample solve a multi-step subproblem, which is known as the hard data consistency process, conditioned on the current reconstruction $\bm{x}_{t}$ to obtain an estimate $\bm{x}_{0}$ at the initial timestep, and (ii) map $\bm{x}_{0}$ back to the diffusion trajectory at time $t-1$.
+While these methods may appear conceptually related to the plug-and-play approaches discussed above, hard data consistency methods generate the final reconstruction within a single reverse diffusion process—meaning the data consistency steps of these methods remain tightly coupled with the reverse sampling.
+In comparison, our approach employs $K$ iterations of forward and reverse diffusion processes (diffusion purification), demonstrating that data consistency steps can be fully decoupled from the reverse sampling steps.
+As a result, our method is compatible with one-step CMs, which are not supported by DiffPIR or ReSample.
+Red-Diff.
+RED-Diff [45] is an algorithm that combines regularization by denoising [46] with pre-trained diffusion models within a variational Bayesian framework.
+It involves solving an optimization problem via gradient descent to minimize a data-fitting loss while maximizing the likelihood of the reconstruction under the diffusion-based denoiser prior.
 4 Experiments
@@
 Second, we show our main experimental results in Section 4.2.
-Finally, we provide an ablation study in Section 5.3 in terms of noise robustness, computational efficiency, and convergence.
+Finally, we provide an ablation study in Section 4.3 in terms of noise robustness, computational efficiency, and convergence speed.
 4.1 Experimental Setup
-Types of Image Restoration Problems.
-To demonstrate the effectiveness of our algorithm, we consider the following image restoration problems:
-  • 1)
-    Image inpainting with a $100\times 100$ size box mask.
-  • 2)
-    Image super-resolution with 4×4$\times$4 × down sampling.
-  • 3)
-    Gaussian image deblurring with Gaussian kernel of size $61\times 61$ and standard deviation $\sigma=3.0$.
-  • 4)
-    Motion image deblurring with motion deblurring kernel size $61\times 61$ and intensity value $0.5$.
-These image restoration problems are commonly explored in numerous relevant works [20, 21, 16, 13] and also arise in many practical applications.
-Evaluation Datasets and Metrics.
-To demonstrate the effectiveness of our method, we conducted experiments on the FFHQ [43] and LSUN-Bedroom [44] datasets, which have an image size of $256\times 256\times 3$.
-For CMs, we test our method only on the LSUN-Bedroom dataset as a proof-of-concept.
-Generally, images of bedrooms have more complicated structures than those in “face” image datasets, making them a good indicator of the efficacy of our algorithm.
-To quantitatively evaluate the performance of the algorithms, we use standard metrics: (i) Peak Signal-to-Noise-Ratio (PSNR), (ii) Structural Similarity Index Measure (SSIM) [45] and (iii) Learned Perceptual Image Patch Similarity (LPIPS) [46].
-These metrics are commonly used in existing works that evaluate image restoration techniques [21], and are shown to be an accurate indicator of reconstruction performance.
-Figure 4: Pixel space diffusion models for solving various image restoration problems.
+We evaluate our method based on the following setups. The specific choices of hyperparameters are provided in Appendix B.
+Inverse Problems Used in Evaluation We evaluate our method on both linear and nonlinear inverse problems.
+For linear inverse problems, we consider the following tasks with no additive measurement noise: (i) image inpainting using a $100\times 100$ box mask; (ii) 4×4$\times$4 × image super-resolution via uniform downsampling; (iii) Gaussian deblurring with a $61\times 61$ Gaussian kernel of standard deviation $\sigma=3.0$; and (iv) motion deblurring using a $61\times 61$ motion blur kernel with intensity $0.5$.
+For nonlinear inverse problems, we consider nonlinear deblurring following the experimental setup in [14], with additive Gaussian measurement noise with standard deviation $0.05$.
+Figure 3: Pixel space diffusion models for solving various linear inverse problems.
 These results demonstrate the effectiveness of our algorithm.
 We can consistently recover images with the highest PSNR.
+Evaluation Datasets.
+For linear inverse problems, we conduct experiments on the FFHQ [22] and LSUN-Bedroom [23] datasets, both of which contain 100 images of size $256\times 256\times 3$.
+For nonlinear inverse problems, we conduct experiments on ImageNet [24] with 100 images of size $256\times 256\times 3$.
+For experiments on CMs, we only test on the LSUN-Bedroom dataset and linear inverse problems as a proof-of-concept.
+Evaluation Metrics.
+To quantitatively evaluate the performance, we use standard metrics: (i) Peak Signal-to-Noise Ratio (PSNR), (ii) Structural Similarity Index Measure (SSIM) [47] and (iii) Learned Perceptual Image Patch Similarity (LPIPS) [48].
 Method Super Resolution 4×4$\times$4 × Inpainting (Box)
 […]
-DPS [16] $\underline{0.225}$$\pm 0.05$ $25.37\pm 3.17$ $0.717\pm 0.09$ $0.155\pm 0.04$ $\underline{23.58}\pm 3.49$ $0.795\pm 0.09$
+DPS [14] $0.225\pm 0.05$ $25.37\pm 3.17$ $0.717\pm 0.09$ $0.155\pm 0.04$ $\underline{23.58}\pm 3.49$ $0.795\pm 0.09$
 MCG [11] $0.282\pm 0.11$ $18.94\pm 4.34$ $0.561\pm 0.15$ $0.159\pm 0.04$ $20.32\pm 3.54$ $0.789\pm 0.06$
 ADMM-PnP [49] $0.349\pm 0.04$ $21.56\pm 2.16$ $0.676\pm 0.07$ $0.232\pm 0.02$ $18.41\pm 2.09$ $0.728\pm 0.03$
-Ours Pixel (Version 1) $\bm{0.215}\pm 0.05$ $\underline{26.10}\pm 3.12$ $\underline{0.768}\pm 0.08$ $\underline{0.098}\pm 0.01$ $20.79\pm 2.66$ $\underline{0.880}\pm 0.02$
-Ours Pixel (Version 2) $0.229\pm 0.05$ $\bm{26.92}\pm 3.33$ $\bm{0.800}\pm 0.07$ $\bm{0.094}\pm 0.02$ $\bm{23.66}\pm 3.42$ $\bm{0.906}\pm 0.02$
+DDNM [13] $0.233\pm 0.06$ $26.00\pm 2.56$ $0.756\pm 0.08$ $0.105\pm 0.06$ $23.42\pm 3.45$ $0.882\pm 0.08$
+RED-Diff [45] $0.267\pm 0.05$ $25.78\pm 1.59$ $0.746\pm 0.07$ $0.115\pm 0.08$ $23.22\pm 3.16$ $0.878\pm 0.05$
+DiffPIR [50] $\underline{0.218}\pm 0.06$ $\underline{26.34}\pm 2.34$ $\underline{0.782}\pm 0.06$ $0.108\pm 0.07$ $23.47\pm 1.89$ $\underline{0.884}\pm 0.06$
+DCDP-DDIM (Ours) $\bm{0.215}\pm 0.05$ $26.10\pm 3.12$ $0.768\pm 0.08$ $\underline{0.098}\pm 0.01$ $20.79\pm 2.66$ $0.880\pm 0.02$
+DCDP-Tweedie (Ours) $0.229\pm 0.05$ $\bm{26.92}\pm 3.33$ $\bm{0.800}\pm 0.07$ $\bm{0.094}\pm 0.02$ $\bm{23.66}\pm 3.42$ $\bm{0.906}\pm 0.02$
 Latent-DPS $0.415\pm 0.06$ $22.96\pm 2.62$ $0.573\pm 0.11$ $0.312\pm 0.05$ $18.66\pm 2.53$ $0.656\pm 0.08$
@@
 ReSample [19] $\bm{0.187}\pm 0.04$ $25.92\pm 3.20$ $0.764\pm 0.08$ $\bm{0.195}\pm 0.04$ $21.05\pm 3.11$ $0.762\pm 0.07$
-Ours LDMs (Version 1) $\underline{0.244}\pm 0.05$ $\underline{26.52}\pm 3.21$ $\underline{0.778}\pm 0.08$ $\underline{0.215}\pm 0.04$ $\underline{23.75}\pm 3.32$ $\bm{0.831}\pm 0.05$
-Ours LDMs (Version 2) $0.259\pm 0.05$ $\bm{26.88}\pm 3.28$ $\bm{0.789}\pm 0.08$ $0.239\pm 0.04$ $\bm{24.22}\pm 3.65$ $\underline{0.828}\pm 0.05$
-Ours CMs $0.208\pm 0.05$ $26.17\pm 3.19$ $0.774\pm 0.08$ $0.108\pm 0.02$ $19.66\pm 2.48$ $0.874\pm 0.02$
-Table 1: Quantitative results of super resolution and inpainting on the LSUN-Bedrooms dataset for both pixel-based and LDMs.
+DCDP-LDM-DDIM (Ours) $\underline{0.244}\pm 0.05$ $\underline{26.52}\pm 3.21$ $\underline{0.778}\pm 0.08$ $\underline{0.215}\pm 0.04$ $\underline{23.75}\pm 3.32$ $\bm{0.831}\pm 0.05$
+DCDP-LDM-Tweedie (Ours) $0.259\pm 0.05$ $\bm{26.88}\pm 3.28$ $\bm{0.789}\pm 0.08$ $0.239\pm 0.04$ $\bm{24.22}\pm 3.65$ $\underline{0.828}\pm 0.05$
+DCDP-CM (Ours) $0.208\pm 0.05$ $26.17\pm 3.19$ $0.774\pm 0.08$ $0.108\pm 0.02$ $19.66\pm 2.48$ $0.874\pm 0.02$
+Table 1: Quantitative results of super resolution and inpainting on the LSUN-Bedrooms dataset for both pixel space diffusion models and LDMs.
 The results colored in blue and in orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models.
-The best results are in bold and second best results are underlined for pixel space models and LDMs respectively.
+The best results are in bold and second best results are underlined for pixel space models and LDMs, respectively.
 Comparison Baselines.
-Since we conduct experiments involving both pixel space and latent space, we consider the following baselines:
-  • Diffusion Models in Pixel Space. For the pixel space diffusion models, we compare our algorithm to DPS [16], Manifold Constrained Gradients (MCG) [13] and Plug-and-Play ADMM [47] (PnP-ADMM).
-  • Diffusion Models in Latent Space. For the experiments with LDMs, we compare to Latent-DPS [21], ReSample [21], and Posterior Sampling with Latent Diffusion Models (PSLD) [20].
-To ensure a fair comparison, we use pre-trained pixel-space diffusion models from Chung et al.[16] and Dhariwal et al.[10], as well as pre-trained CMs from Song et al.
-[19], for both the baseline techniques and our methods in the pixel space.
-Similarly, for diffusion models in the latent space, we use models provided by Rombach et al.
-[24], which were previously used in existing methods [21].
+We conduct experiments using both pixel space and latent space diffusion models.
+For the pixel space, we compare our algorithm with DPS [14], Manifold Constrained Gradients (MCG) [11], Plug-and-Play ADMM [49] (PnP-ADMM), DiffPIR [50], RED-Diff [45] and DDNM [13].
+For the latent space, we consider Latent-DPS [19], ReSample [19], and Posterior Sampling with Latent Diffusion Models (PSLD) [18].
+To ensure a fair comparison, both the baselines and DCDP (ours) use the same sets of pre-trained diffusion models, including pixel space diffusion models from [14] and [7], LDMs from [28], and CMs from [17].
+For both DCDP and the baselines considered, we tune the hyperparameters via grid search to obtain the best possible performance.
 Method Gaussian Deblurring Motion Deblurring
@@
 ADMM-PnP [49] $0.437\pm 0.04$ $20.76\pm 1.94$ $0.595\pm 0.09$ $0.524\pm 0.04$ $18.05\pm 2.05$ $0.493\pm 0.10$
-Ours Pixel (Version 1) $\bm{0.196}\pm 0.05$ $\underline{27.13}\pm 3.26$ $\underline{0.804}\pm 0.07$ $\underline{0.065}\pm 0.01$ $\underline{33.56}\pm 3.92$ $\underline{0.947}\pm 0.02$
-Ours Pixel (Version 2) $\underline{0.212}\pm 0.05$ $\bm{27.74}\pm 3.34$ $\bm{0.825}\pm 0.07$ $\bm{0.067}\pm 0.02$ $\bm{34.47}\pm 4.07$ $\bm{0.956}\pm 0.02$
+DDNM [13] $0.217\pm 0.07$ $\underline{27.30}\pm 1.67$ $\underline{0.815}\pm 0.08$
+RED-Diff [45] $0.228\pm 0.12$ $23.98\pm 1.23$ $0.687\pm 0.067$ $0.178\pm 0.06$ $24.15\pm 2.34$ $0.704\pm 0.12$
+DiffPIR [50] $0.214\pm 0.013$ $27.04\pm 2.67$ $0.801\pm 0.12$ $0.114\pm 0.12$ $27.45\pm 1.89$ $0.866\pm 0.65$
+DCDP-DDIM (Ours) $\bm{0.196}\pm 0.05$ $27.13\pm 3.26$ $0.804\pm 0.07$ $\underline{0.065}\pm 0.01$ $\underline{33.56}\pm 3.92$ $\underline{0.947}\pm 0.02$
+DCDP-Tweedie (Ours) $\underline{0.212}\pm 0.05$ $\bm{27.74}\pm 3.34$ $\bm{0.825}\pm 0.07$ $\bm{0.067}\pm 0.02$ $\bm{34.47}\pm 4.07$ $\bm{0.956}\pm 0.02$
 Latent-DPS $0.337\pm 0.05$ $23.75\pm 2.53$ $0.622\pm 0.10$ $0.425\pm 0.06$ $21.90\pm 2.31$ $0.539\pm 0.10$
@@
 ReSample [19] $\underline{0.240}\pm 0.05$ $25.76\pm 3.02$ $0.731\pm 0.09$ $0.188\pm 0.04$ $27.96\pm 3.07$ $0.806\pm 0.07$
-Ours LDMs (Version 1) $0.246\pm 0.05$ $\underline{26.08}\pm 3.02$ $\underline{0.766}\pm 0.07$ $\underline{0.145}\pm 0.03$ $\underline{28.87}\pm 3.77$ $\underline{0.856}\pm 0.06$
-Ours LDMs (Version 2) $\bm{0.217}\pm 0.05$ $\bm{27.11}\pm 3.26$ $\bm{0.807}\pm 0.07$ $\bm{0.121}\pm 0.03$ $\bm{29.40}\pm 3.81$ $\bm{0.875}\pm 0.05$
-Ours CMs $0.202\pm 0.05$ $26.91\pm 3.11$ $0.805\pm 0.07$ $0.061\pm 0.02$ $34.10\pm 3.80$ $0.953\pm 0.02$
-Table 2: Quantitative results of Gaussian and motion deblurring on the LSUN-Bedrooms dataset for both pixel-based and LDMs.
+DCDP-LDM-DDIM (Ours) $0.246\pm 0.05$ $\underline{26.08}\pm 3.02$ $\underline{0.766}\pm 0.07$ $\underline{0.145}\pm 0.03$ $\underline{28.87}\pm 3.77$ $\underline{0.856}\pm 0.06$
+DCDP-LDM-Tweedie (Ours) $\bm{0.217}\pm 0.05$ $\bm{27.11}\pm 3.26$ $\bm{0.807}\pm 0.07$ $\bm{0.121}\pm 0.03$ $\bm{29.40}\pm 3.81$ $\bm{0.875}\pm 0.05$
+DCDP-CMs (Ours) $0.202\pm 0.05$ $26.91\pm 3.11$ $0.805\pm 0.07$ $0.061\pm 0.02$ $34.10\pm 3.80$ $0.953\pm 0.02$
+Table 2: Quantitative results of Gaussian and motion deblurring on the LSUN-Bedrooms dataset for both pixel space diffusion models and LDMs.
 The results colored in blue and orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models.
 The best results are in bold and second best results are underlined for pixel space models and LDMs respectively.
-5.2 Main Experimental Results
-Here, we present both quantitative and qualitative results for our proposed algorithm and the baselines.
-The algorithms are tested with clean measurements generated from $100$ randomly selected images from each of the FFHQ and LSUN-Bedrooms datasets.
-For conciseness, we only display results on the LSUN-Bedroom dataset and defer the results on FFHQ to Appendix B.
-We also defer the choices of hyperparameters in Section A.2.
-5.2.1 Pixel Space Diffusion Models
-In this section, we present the performance of our proposed method when using diffusion models trained in the pixel space.
-Recall that we have two versions of our method for the pixel space: vanilla reverse SDE sampling as well as an approximation method via Tweedie’s formula.
-We label the vanilla reverse sampling method as Version 1 (V1) and the Tweedie’s fomula approximation method as Version 2 (V2).
-We also demonstrate the use case of our algorithm using CMs as discussed in Section 4.1.
-Figure 5: Demonstrating the differences between Versions 1 and 2 on super resolution (4×4$\times$4 ×).
-The two columns show the reconstructions with PSNRs and their corresponding “canny edges”, respectively.
-The qualitative results are reported in Figure 4 and the quantitative comparison of pixel space diffusion model algorithms is shown in Table 1 and Table 2.
-In these tables, it is evident that our methods noted as V1 and V2 consistently outperform the existing baselines.
-However, even though the quantitative results imply that V2 is able to achieve higher performance on standard metrics such as PSNR and SSIM, the reconstructions are often blurrier and contain less perceptual detail compared to V1.
-As shown in Figure 5, the reconstruction of V1 has more high frequency information, visualized with “canny edges”, despite having lower PSNR.
-This phenomenon is commonly referred to as the “perception-distortion tradeoff” [48, 49].
-Surprisingly, we observe that MCG seems to have a lot of variability in the results on the LSUN-Bedroom dataset, but performs well on FFHQ (shown in Appendix B).
-This might suggest that datasets such as faces are generally easier than datasets with more complicated structures such as bedrooms.
-For the results with CMs, we present the quantitative results in Table 1 and  Table 2 highlighted in orange, with visual results in Figure 6.
-This highlights another use case of our algorithm, as it demonstrates that our algorithm can flexibly utilize CMs compared to existing methods.
+4.2 Experimental Results
+In this section we present both quantitative and qualitative results for our proposed algorithm and the baselines.
+For conciseness, we defer the quantitative results for FFHQ to Appendix D.
+As discussed in Section 3.3, the diffusion purification can be performed by either (i) employing accelerated samplers such as DDIM or (ii) applying a single-step approximation using Tweedie’s formula.
+We denote these variants as DCDP-DDIM and DCDP-Tweedie, respectively—and as DCDP-LDM-DDIM and DCDP-LDM-Tweedie when implemented with LDMs—and evaluate both of them in our experiments.
+Figure 4: Demonstrating the differences between the DCDP methods on SR (4×4$\times$4 ×). The two columns show the reconstructions with PSNRs and their corresponding canny edges.
 Figure 5: Qualitative results for various image restoration problems by using CMs. Our proposed framework is capable of utilizing CMs for image restoration.
+Comparison of Pixel Space Diffusion Models.
+We first test DCDP and the baselines based on pixel space diffusion models.
+The qualitative results are reported in Figure 3, and the quantitative results are presented in Tables 1, 2 and 3.
+In these tables, it is evident that our methods consistently outperform or achieve performance on par with the baselines.
+We note that Red-Diff [45] performs extremely well on nonlinear deblurring but is less competitive on other tasks.
+When comparing DCDP-DDIM and DCDP-Tweedie, we observe that DCDP-Tweedie generally achieves higher PSNR and SSIM, which mainly measure the data fidelity.
+However, its reconstructions are often blurrier and contain less perceptual detail.
+As shown in Figure 4, DCDP-DDIM qualitatively preserves richer high-frequency content—evident in the Canny-edge visualizations—and leads to a better LPIPS score, which more aligns with the perceptual quality.
+This phenomenon is commonly referred to as the “perception-distortion trade-off” [51].
+For the results with CMs, we present the quantitative results in Tables 1 and 2 highlighted in orange, and qualitative results in Figure 5, which demonstrate that DCDP with CMs leads to decent reconstruction quality for the inverse problems considered.
+This result highlights the flexibility of our method to accommodate a broad range of diffusion-based models.
+In contrast, none of the baselines considered are able to utilize CMs, as they require a sufficiently large number of reverse sampling steps.
+Lastly, we note that the CMs used in this work are adopted from  [17], which is the first ever prototype that introduces the concept of CMs.
+The performance of DCDP with CMs could potentially be further improved by incorporating more advanced consistency models proposed in recent works such as sCMs [52], Latent Consistency Models [53] and InstaFlow [54], which we leave for future investigation.
 Figure 6: Latent space diffusion models for solving various image restoration problems.
@@
 We can consistently recover images with the highest PSNR.
-5.2.2 Latent Space Diffusion Models
-For the results with LDMs, we display quantitative results in Table 1 and Table 2 highlighted in blue, with qualitative results in Figure 7.
-Recall that as outlined in Section 4.2, there are two approaches for ensuring data consistency.
-For the best performance, we use pixel space optimization (Approach 2) for motion deblurring and Gaussian deblurring and latent space optimization (Approach 1) for box inpainting and super-resolution.
-As observed in the tables, we observe that our algorithm outperforms the existing baselines on many of the inverse problems.
-Interestingly, we observe that PSLD seems to reconstruct images that are often noisy and blurry, where the learning rate plays a huge role on the reconstruction quality.
-While ReSample may have higher LPIPS scores on some inverse problem tasks, our algorithm is able to achieve better PSNR and SSIM scores.
+Comparison of Latent Space Diffusion Models.
+For the results with LDMs, we present quantitative results in Tables 1, 2 and 3 highlighted in blue, and qualitative results in Figure 6.
+Recall that, as outlined in Section 3.4, there are two approaches for enforcing data consistency.
+We use pixel space data consistency (Approach II) for motion deblurring and Gaussian deblurring, and latent space optimization (Approach I) for box inpainting, super-resolution, and nonlinear deblurring, as doing so empirically leads to the best reconstruction quality.
+As shown in the tables, our algorithm consistently matches or outperforms all existing baselines.
+Interestingly, we observe that PSLD seems to reconstruct images that are often noisy and blurry, where the learning rate plays a huge role in the reconstruction quality.
+While ReSample could achieve higher LPIPS scores on some inverse problem tasks, our algorithm is able to achieve better PSNR and SSIM scores.
+Method Nonlinear Deblurring
+[…]
+DPS [14] $0.355\pm 0.09$ $22.60\pm 3.36$ $0.588\pm 0.16$
+RED-Diff [45] $\textbf{0.211}\pm 0.05$ $\textbf{29.51}\pm 0.76$ $\textbf{0.828}\pm 0.08$
+DCDP-DDIM (Ours) $0.323\pm 0.06$ $24.08\pm 2.42$ $0.634\pm 0.112$
+DCDP-Tweedie (Ours) $\underline{0.320}\pm 0.08$ $\underline{25.36}\pm 2.77$ $\underline{0.712}\pm 0.10$
+Latent-DPS $0.483\pm 0.08$ $21.77\pm 3.02$ $0.513\pm 0.16$
+ReSample [19] $\bm{0.279}\pm 0.07$ $\textbf{26.36}\pm 2.90$ $0.717\pm 0.11$
+DCDP-LDM-DDIM (Ours) $\underline{0.286}\pm 0.08$ $25.97\pm 3.23$ $\underline{0.735}\pm 0.12$
+DCDP-LDM-Tweedie (Ours) $0.302\pm 0.09$ $\underline{26.34}\pm 3.34$ $\textbf{0.747}\pm 0.12$
+Table 3: Quantitative results of nonlinear deblurring on the ImageNet dataset for both pixel space diffusion models and LDMs.
+The results colored in blue are for LDMs, while the others are for pixel space diffusion models.
+The best results are in bold and second best results are underlined for pixel space models and LDMs respectively.
 4.3 Ablation Studies
-Finally, we conduct ablation studies demonstrating computational efficiency, convergence, and robustness to measurement noise.
+Finally, we conduct ablation studies demonstrating computational efficiency, convergence, and robustness of the proposed DCDP method.
+Figure 7: Comparison of inference times on Motion Deblurring.
+The experiment setting is the same as that introduced in Section 4.1 and all methods are run using the exact same compute.
+Here V1 and V2 stands for Approach I and Approach II introduced in Section 3.3 respectively.
+The inference time for our algorithm is the fastest, while enjoying similar (or better) performance.
+The runtime is normalized by dividing by the largest value, which corresponds to DPS with a total runtime of 387 seconds.
 Computational Efficiency.
-First, we compare the inference times (the time it takes to reconstruct a single image) of our method with the baselines and display the results in Table 3.
-Clearly, for pixel based diffusion models, our methods are much more efficient than DPS [16] and MCG [13], as we do not require backpropogation through the score function for data consistency.
-Furthermore, using the single step Tweedie’s formula approximation for diffusion purification is much faster than using reverse SDE which involves multiple time steps.
-Second, for latent space methods, our methods are also faster than existing methods PSLD [20] and ReSample [21] as observed in Table 3.
-For our own methods, we observe that optimizing at the pixel level (Approach 2) for data consistency proves to be considerably more efficient than optimizing in the latent space (Approach 1).
-As previously mentioned, this is due to the need for backpropagating through the complex deep decoder $\mathcal{D}(\cdot)$.
-This is also one of the main reasons our methods are more computationally efficient than methods such as ReSample.
-Algorithm Time (sec)
-DPS [16] 387
-MCG [13] 385
-Our Pixel (Reverse SDE, Version 1) 28.7
-Our Pixel (Tweedie’s Formula, Version 2) 2.14
-PSLD [20] 188
-ReSample [21] 118
-Our Latent (Latent DC + Reverse SDE) 109
-Our Latent (Pixel DC + Reverse SDE) 12.2
-Our Latent (Latent DC + Tweedie’s Formula) 100
-Our Latent (Pixel DC + Tweedie’s Formula) 2.59
-Table 3: Comparison of inference times. The results colored in blue are for LDMs, whereas the others are for pixel space diffusion models.
+First, we compare the inference speed (the time it takes to reconstruct a single image) of our method with the baselines and display the results in Figure 7.
+Clearly, for pixel-based diffusion models, our methods are much more efficient than DPS [14] and MCG [11], as we do not require backpropagation through the score function for data consistency.
+Furthermore, using the single-step Tweedie’s formula approximation for diffusion purification (DCDP-Tweedie and DCDP-LDM-Tweedie) is much faster than using the reverse ODE (DDIM) which involves multiple reverse diffusion steps (DCDP-DDIM and DCDP-LDM-DDIM).
+Second, for LDMs, our methods are faster than both ReSample and PSLD.
+Moreover, enforcing data consistency in the pixel space (Approach II) is considerably more efficient than enforcing data consistency in the latent space (Approach I), as the former does not require backpropagation through the deep decoder $\mathcal{D}(\cdot)$.
 Figure 8: Convergence of our methods in terms of LPIPS.
-The orange, green, red and blue curves correspond to the reconstruction process with data fidelity optimization only, our proposed method with a constant (large) purification strength, constant (small) purification strength and a linear decaying purificaiton strength, respectively.
+The orange, green, red and blue curves correspond to the reconstruction process with data fidelity optimization only, DCDP with a constant (large) purification strength, constant (small) purification strength and a linear decaying purificaiton strength, respectively.
 The x-axis denotes the total number of data fidelity gradient steps.
-Convergence Study.
-In this section, we investigate the convergence of our algorithm in terms of LPIPS.
-This study also serves to demonstrate the efficiency of a linear decaying time schedule for $T_{k}$ for diffusion purification across various image restoration problems.
+See Figure 20 in Appendix D for convergence behavior measured using PSNR and SSIM.
+Empirical Convergence Study.
+We further investigate the convergence of our algorithm in terms of LPIPS.
+This study also serves to demonstrate the efficiency of a linear decaying time schedule for $t_{k}$ in diffusion purification.
 As shown in Figure 8, at the beginning of each diffusion purification and data fidelity optimization cycle, the LPIPS increases after performing diffusion purification due to stochasticity introduced by the forward-backward diffusion process.
-However, the subsequent iterations can successfully decrease LPIPS and result in reconstructions with higher qualities compared to pure data fidelity optimization, illustrating that the resulting purified image is able to serve as a good initialization point.
-In the same figure, also notice that the linear decaying schedule greatly enhances the performance of our algorithm compared to using a constant time schedule for $T_{k}$.
+However, the subsequent iterations successfully decrease LPIPS and result in reconstructions with higher quality compared to pure data fidelity optimization, implying that the purified image is able to serve as a good initialization point.
+In the same figure, note that the linear decaying schedule greatly improves the performance of our algorithm compared to using a constant time schedule for $t_{k}$.
 This is expected, as a large constant purification strength will result in serious hallucination while a small constant purification strength is not able to fully alleviate the artifacts in the data fidelity based estimations.
-In Figure 9, we illustrate an example reconstruction process for super-resolution with noisy measurements, with more examples in the Appendix.
-It is evident that the quality of the reconstructed image gradually increases with more iterations.
+We note that similar convergence behavior holds when measuring the reconstruction quality with PSNR and SSIM, as shown in Figure 20 in Appendix D.
+Lastly, in Figure 9, we illustrate an example reconstruction process for super-resolution with noisy measurements, with more qualitative examples in the Appendix.
+It is evident that the quality of the reconstructed image gradually increases with each iteration, implying that the diffusion purification is an effective operator for enforcing the image priors embedded in the pre-trained diffusion models.
 Figure 9: Reconstruction process for super-resolution (4x) in the pixel space.
@@
 Super Resolution (4×4$\times$4 ×) Motion Deblurring
-Metric DPS [16] Version 1 Version 2 DPS [16] Version 1 Version 2
+Metric DPS DCDP-DDIM DCDP-Tweedie DPS DCDP-DDIM DCDP-Tweedie
 […]
@@
 Noise Robustness.
-Thus far, we considered experiments on image restoration tasks in which the measurements were noiseless.
-However, in many practical settings of interest, the measurements $\bm{y}$ can be corrupted with additive noise.
-In this section, we perform an ablation study on the performance of our algorithm in which the measurements have additive Gaussian noise $\mathcal{N}(0,\sigma^{2}_{\bm{y}})$, with varying values of $\sigma_{\bm{y}}$ for super resolution and motion deblurring.
-We compare our algorithm to DPS, as their theory states that the approximation of their gradient likelihood term actually becomes tighter as the measurement noise $\sigma_{\bm{y}}$ increases [16].
-We conduct experiments on $100$ randomly selected images from the LSUN-Bedroom dataset for super-resolution, motion deblurring on standard deviations $0.05,0.075$, and $0.1$.
-For the hyperparameters, we tune the parameters according to $\sigma_{\bm{y}}=0.01$ and fix them throughout the experiments.
+In many practical settings of interest, the measurements $\bm{y}$ can be corrupted with additive noise.
+Here, we perform an ablation study on the performance of our algorithm in which the measurements have additive Gaussian noise $\mathcal{N}(0,\sigma^{2}_{\bm{y}}\bm{I})$ with varying values of $\sigma_{\bm{y}}$ for super-resolution and motion deblurring.
+We compare our algorithm to DPS, as their theory states that the approximation of the likelihood gradient becomes tighter as the measurement noise $\sigma_{\bm{y}}$ increases [14].
+We conduct experiments on $100$ randomly selected images from the LSUN-Bedroom dataset for super-resolution, motion deblurring with Gaussian measurement noise of standard deviations $0.05,0.075$, and $0.1$.
+For the hyperparameters, we tune the parameters according to $\sigma_{\bm{y}}=0.1$ and fix them throughout the experiments.
 The rest of the setup is the same as previously described.
 We present the results in Table 4.
-Here we observe that as the standard deviation increases, the performance of all algorithms slightly degrades, as expected.
+Note that as the standard deviation increases, the performance of all algorithms degrades.
 However, our algorithm consistently outperforms DPS in metrics such as PSNR and SSIM, whereas DPS often exhibits higher perceptual quality measured by LPIPS.
@@
 Our algorithm is less susceptible to such artifacts, resulting in “smoother” images with higher signal-to-noise ratios.
-Details of the hyperparameters and visual results can be found in Section A.3.
-6 Relationship to Prior Arts
-Here, we discuss the relationship of our work with existing methods for image restoration, including both methods derived in the pixel space and algorithms tailored for the latent space.
-Pixel Space Diffusion Models.
-Other than DPS [16], there are other diffusion model-based inverse problem solvers available for addressing image restoration tasks.
-Song et al.
-[17] introduced $\Pi$GDM, an algorithm that approximates the likelihood gradient using a pseudo-inverse of the linear sensing matrix used for generating the measurements.
-When the operator $\mathcal{A}(\cdot)$ is linear, it can be represented by a matrix $\bm{A}\in\mathbb{R}^{m\times n}$, for which one can calculate its pseudo-inverse.
-They leverage this property by first constructing the conditional distribution $p(\bm{y}|\bm{x}_{t})$ in terms of the sensing matrix $\bm{A}$ and derive the likelihood gradient using this conditional distribution.
-The work by Wang et al.
-[15] has a similar flavor; since the image restoration problem is ill-posed, the reconstructions without priors generally are a mixture of a sample in both the range and the nullspace of the matrix $\bm{A}$.
-They use the diffusion model as a prior to essentially “fill in” this nullspace, which is constructed by taking the pseudo-inverse of $\bm{A}$ and used during the reverse sampling process.
-The work presented in [50] introduced the Direct Diffusion Bridge method building upon the DPS approach.
-Here, the data fidelity is enforced during the reverse sampling procedure by taking a deeper gradient step which is designed for improved data consistency.
-Another popular method for addressing image restoration problems is the Denoising Diffusion Restoration Models (DDRMs) proposed by Kawar et al.
-[51].
-DDRMs construct a probabilistic graph conditioned on the measurements, allowing the use of a pre-trained unconditional diffusion model.
-Subsequently, they perform the reverse sampling process in the “spectral space”, which is constructed by taking the SVD of the sensing matrix $\bm{A}$.
-These works generally share two common themes: they require the inverse problem to be linear, and they incorporate the data consistency steps in the reverse diffusion process.
-Although we do not test on nonlinear inverse problems, our algorithm does not face these limitations and is flexible in this regard.
-Lastly, the study presented by Luther et al.
-[52] shares close ties with Version 2 our algorithm in the pixel space.
-They introduced Diffusive Denoising of Gradient-based Minimization (DDGM), an image-space method for solving linear tomographic reconstruction problems, which iteratively applies data consistency steps and image denoising through a pre-trained diffusion model.
-However, as we have shown in Section 5.2, directly regularizing images with Tweedie’s formula will result in blurry images.
-In contrast, the proposed Version 1 of our algorithm alleviates this issue by performing diffusion purification with multiple diffusion reverse sampling steps.
-Furthermore, we propose a unified framework that can incorporate accelerated samplers, latent diffusion models, and consistency models, thereby significantly reducing the inference time.
-Hence, our work addresses a broader spectrum, not only limited to tomographic reconstruction, and can be readily adapted to applications requiring fast inference.
-Latent Space Diffusion Models.
-Throughout our experiments, we also presented results using LDMs, using PSLD [20] and ReSample [21] as baselines.
-ReSample [21] is an algorithm designed to solve inverse problems in the latent space that involves solving a data fidelity optimization problem during the reverse sampling process, a process they term “hard data consistency”.
-The data fidelity objective is similar to the approaches we take, where they alternate between minimizing the objective in either the latent space or the pixel space.
-Upon solving this objective, they devise a resampling scheme to project the data-consistent sample back onto the noisy manifold to continue the reverse sampling process.
-Nonetheless, by decoupling these two processes, our algorithm does not require a resampling stage.
-Additionally, ReSample cannot use accelerated samplers because of the coupling, as it would hinder the data fidelity stage by reducing the number of sampling steps, unlike our framework.
-Posterior Sampling with Latent Diffusion (PSLD) [20] extends the idea of DPS to the latent space by addressing the challenge that the decoder and encoder constitute a many-to-one mapping.
-This implies that there may be multiple latent codes that fit the measurements, and they devise a gradient scheme to ensure the “correct” latent code is estimated during the reverse sampling process.
-However, their technique involves two gradient computations—one through the LDM decoder and one through the pre-trained LDM score function.
-Thus, PSLD suffers from increased inference time.
-Furthermore, it is observed in our experiment that the performance of PSLD is sensitive to the choice of hyperparameters.
-Plug-and-Play (PnP) Methods.
-The Plug-and-Play Priors (PnP) framework combines physical and learned models for effective computational imaging.
-PnP alternates between minimizing a data-fidelity term for enforcing data consistency and applying a learned regularizer, typically an image denoiser [53, 36, 37].
-The PnP framework is commonly formulated using the Alternating direction method of multipliers (ADMM) [36] and fast iterative shrinkage/thresholding algorithms (FISTA) [54].
-Our proposed method shares similarity with the PnP in the sense that both methods utilize image denoisers as the implicit prior.
-However, while PnP performs denoising directly on $\bm{x}_{k}$ by assuming that $\bm{x}_{k}$ contains certain pseudo noise, our proposed method uses diffusion models to explicitly add noise to $\bm{x}_{k}$ and then denoises it by either iterative sampling or Tweedie’s formula [52] estimation.
-7 Discussion & Conclusion
-Throughout this paper, we have discussed both the efficiency and flexibility of our algorithm, demonstrating that we can compete with many state-of-the-art techniques while offering numerous advantages.
-However, despite these strengths, our algorithm has a few steps that warrant further investigation.
-Recall that the input to the diffusion purification stage is an initial reconstruction estimated from minimizing the data fidelity objective.
-Notice that if this step does not return an estimate that is meaningful, then the diffusion purification stage cannot produce a high-quality reconstruction.
-Although this was not an issue for the image restoration problem that we considered, it may be the case for more complex problems, as simply minimizing the data fidelity objective in Equation (10) can lead to poor local minima.
-For these tasks, it would be of great interest to explore replacing Equation (10) with more advanced solvers.
-Furthermore, while our algorithm generally outperforms others in metrics such as PSNR and SSIM, it often only shows marginal improvements in LPIPS, which measures perceptual quality.
-This outcome is expected, as coupled methods like DPS prioritize the prior step more meticulously, potentially leading to hallucinations.
-On the other hand, by decoupling data fidelity from the prior, our algorithm places greater emphasis on data fidelity.
-Therefore, exploring methods to carefully balance between the prior and data consistency with decoupling is an important avenue for future research.
-8 Data Availability Statement
+Details of the hyperparameters and visual results can be found in Appendix C.
+Impact of Number of DDIM Steps.
+Next, we investigate how the number of DDIM steps for diffusion purification affects the reconstruction quality.
+We focus on solving the nonlinear deblurring with pixel space diffusion models (DCDP-DDIM), varying number of DDIM steps among $\{10,20,30\}$, while fixing all other hyperparameters (see Appendix D for details).
+As shown in Figure 10(a), increasing the number of DDIM steps degrades PSNR and SSIM—metrics that primarily capture data fidelity—while improving the LPIPS score, which better reflects perceptual quality.
+This reveals a trade-off between fidelity and perceptual quality controlled by the number of DDIM steps.
+Based on this observation, we fix the number of DDIM steps to 20 in all DCDP-DDIM and DCDP-LDM-DDIM experiments to strike a balance between the two.
+Figure 10: Impact of number of DDIM steps and total number of iterations $K$.
+Figure (a) illustrates the performance of DCDP-DDIM with different number of DDIM steps for diffusion purification.
+Figure (b) and (c) illustrate the impact of the total number of iterations $K$ on the performance of DCDP-DDIM and DCDP-Tweedie respectively.
+Impact of Total Number of Iterations $K$.
+Lastly, we investigate how the number of total iterations $K$ affects reconstruction quality.
+Again, we focus on solving the nonlinear deblurring using pixel space diffusion models (both DCDP-DDIM and DCDP-Tweedie), varying $K$ over $\{5,10,20,30,40\}$ while keeping all other hyperparameters fixed (see Appendix D for details).
+In each iteration, we use a fixed number of steps $\tau=100$ for data fidelity optimization, resulting in a total of $\{500,1000,2000,3000,4000\}$ data consistency steps.
+As shown in Figure 10(b) and (c), increasing $K$ initially improves performance across standard metrics, but the gains plateau or even degrade beyond a certain point.
+In practice, we find that choosing $K$ between 10 and 20 strikes a good balance between reconstruction quality and inference speed.
+5 Conclusion
+In this study, we introduce a novel diffusion-based method for solving general inverse problems, which decouples the data consistency steps from the reverse diffusion process.
+Our technique alternates between a reconstruction phase that enforces data consistency and a refinement stage that uses diffusion purification to enforce the diffusion image prior.
+We demonstrated the effectiveness of our algorithm across a wide range of inverse problems, including deblurring, inpainting, and super-resolution.
+Notably, by decoupling the two processes, we showed that our algorithm can incorporate a broader range of diffusion methods, such as accelerated samplers, latent diffusion models, and consistency models.
+Acknowledgements
+QQ, XL, SMK acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, MICDE Catalyst Grant, and a gift grant from KLA.
+SL and SR acknowledge support from NSF CCF-2212065.
+Data Availability Statement
 The code and instructions for reproducing the experiment results will be made available in the following link: https://github.com/Morefre/Decoupled-Data-Consistency-with-Diffusion-Purification-for-Image-Restoration.
-Acknowledgement
-QQ, XL, SMK acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, MICDE Catalyst Grant, and a gift grant from KLA.
-IA and SR acknowlege support from NSF CCF-2212065.
-Appendix A Implementation Details
-Notation Definition
-$K\in\mathbb{N}$ Number of iterations for overall algorithm
-$\tau\in\mathbb{N}$ Total number of iterations for data consistency optimization
-$\alpha\in\mathbb{R}$ Learning rate for data consistency optimization
-$T_{k}\in\mathbb{R}$ Diffusion purification strength at the $k^{\text{th}}$ iteration
-$\bm{v}_{k}\in\mathbb{R}^{n}$ Prior consistent sample at $k^{\text{th}}$ iteration
-$\bm{x}_{k}\in\mathbb{R}^{n}$ Data consistent sample at $k^{\text{th}}$ iteration
-Table 5: Summary of the notation used throughout this work.
-A.1 Discussion on Hyperparameters
-Here, we discuss the hyperparameters associated with our algorithm and some typical choices we made throughout our experiments.
+Appendix A Discussion on Hyperparameters
+Here, we discuss the hyperparameters of our algorithm and some typical choices we make throughout our experiments.
   • Total Number of Iterations.
@@
 Empirically, we observe a $K$ ranging from 10 to 20 is sufficient for generating high quality reconstructions.
-We suggest users start with $K=10$ and then gradually decrease it if a faster computation speed is required.
+We suggest users start with $K=10$ and then gradually decrease it if a faster computation speed is required or increase it if higher reconstruction quality is required.
   • Optimizer for Data Fidelity Optimization.
@@
 The optimal learning rate ($lr$) for data fidelity optimization varies from task to task.
-In general, the LR should neither be too large, which will not result in images with meaningful structures, nor too small, in which case the convergence speed of the overall algorithm will be too slow.
+In general, $lr$ should neither be too large, which will not result in images with meaningful structures, nor too small, in which case the convergence speed of the overall algorithm will be too slow.
 Throughout the experiment we fix the number of gradient steps for each round of the data fidelity optimization.
@@
   • Decaying Time Schedule for Diffusion Purification.
-A decaying time schedule plays the key role in the success of our proposed algorithm.
+The decaying time schedule plays the key role in the success of our proposed algorithm.
 As described in  Section 3.2, given a data fidelity based reconstruction $\bm{x}_{k}$, the goal of diffusion purification is to generate $\bm{v}_{k}$ that shares the overall structural and semantic information of $\bm{x}_{k}$, such that it is consistent with the measurements while at the same time being free from the artifacts visible in $\bm{x}_{k}$.
-Unfortunately, the $\bm{v}_{k}$ resulting from diffusion purification is sensitive to the choice of $T$, i.e., the amount of noise added to $\bm{x}_{k}$.
-When $T$ is small, the resulting $\bm{v}_{k}$ will exhibit similar structure to $\bm{x}_{k}$ but with artifacts, as there is not enough noise to submerge these artifacts through diffusion purification.
-On the other hand, when $T$ is too large, though $\bm{v}_{k}$ will be free from artifacts, its semantic information and overall structure will significantly deviate from $\bm{x}_{k}$.
+Unfortunately, the $\bm{v}_{k}$ resulting from diffusion purification is sensitive to the choice of $t_{k}$, i.e., the amount of noise added to $\bm{x}_{k}$.
+When $t_{k}$ is small, the resulting $\bm{v}_{k}$ will exhibit similar structure to $\bm{x}_{k}$ but with artifacts, as there is not enough noise to submerge these artifacts through diffusion purification.
+On the other hand, when $t_{k}$ is too large, though $\bm{v}_{k}$ will be free from artifacts, its semantic information and overall structure will significantly deviate from $\bm{x}_{k}$.
 As a result, the resulting reconstructions will not be measurement-consistent.
     To this end, previous works that utilize diffusion purification [29, 12] depend on fine-tuning an optimal $t^{\star}$.
-However, we find that the fine-tuning strategy can be infeasible since the optimal $T^{\star}$ varies from task to task and even from image to image.
+However, this fine-tuning strategy can be infeasible since the optimal $t^{\star}$ varies from task to task and even from image to image.
 To alleviate the need to identify the optimal $t^{\star}$, we implement a linear decaying schedule $\{t_{k}\}_{k\in[K]}$ for each iteration $k$.
-In the early iterations when $T_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though inconsistent with the measurements, will be free from artifacts and can serve as good starting point for the subsequent iteration of data fidelity optimization.
+In the early iterations when $t_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though inconsistent with the measurements, will be free from artifacts and can serve as a good starting point for the subsequent iteration of data fidelity optimization.
 As $t_{k}$ keeps decreasing, the algorithm will converge at an equilibrium point which is both consistent with the measurements and free from the artifacts.
-When the measurements are noiseless, we observe that for a diffusion model trained with 1000 total timesteps, a linear decaying time schedule starting from $T_{1}=400$ and ending at $T_{K}=0$ results in reconstructions with high quality for Super-Resolution, Gaussian Deblurring and Motion Deblurring.
-For Box Inpainting, we choose a higher initial timestep $T_{1}=700$ and $T_{1}=500$ for pixel space diffusion models and LDMs respectively to compensate the severe artifacts in the large box region.
-When the measurements are corrupted by noise, one may instead end at $T_{K}=T$, where $T>0$ to compensate the negative effect of the measurement noise.
-We demonstrate the reconstruction process for various tasks in Figures 10, 11, 12, 13 and 16.
-Notice that the diffusion purification removes the artifacts from the initial reconstructions obtained via data fidelity optimization especially in the early iterations and results in high quality reconstructions consistent with the measurements as timestep decreases.
+When the measurements are noiseless, we observe that for a diffusion model trained with 1000 total timesteps, a linear decaying time schedule starting from $t_{1}=400$ and ending at $t_{K}=0$ results in reconstructions with high quality for Super-Resolution, Gaussian Deblurring, and Motion Deblurring.
+For Box Inpainting, we choose a higher initial timestep $t_{1}=700$ and $t_{1}=500$ for pixel space diffusion models and LDMs respectively to compensate for the severe artifacts in the large box region.
+When the measurements are corrupted by noise, one may instead end at $t_{K}=\tilde{t}$, where $\tilde{t}>0$, to compensate for the negative effect of the measurement noise.
+We demonstrate the reconstruction process for various tasks in Figures 11, 12, 13, 14 and 19.
+Notice that the diffusion purification removes the artifacts from the initial reconstructions obtained via data fidelity optimization, especially in the early iterations, and results in high-quality reconstructions consistent with the measurements as the timestep decreases.
     Figure 11: Reconstruction process for super-resolution with pixel space diffusion models.
@@
   • DDIM Steps for Diffusion Purification.
-With the help of accelerated diffusion samplers such as DDIMs [11], we can perform diffusion purification with as low as 20 unconditional sampling steps.
+With the help of accelerated diffusion samplers such as DDIMs [8], we can perform diffusion purification with as few as 20 unconditional sampling steps.
 As mentioned before, the total number of iterations $K$ ranging from 10-20 is sufficient for generating high quality reconstructions.
@@
   • Pixel Space Optimization and Latent Space Optimization.
-As discussed in Section 4.2, in the case of LDMs, one can use either Pixel Space Optimization Equation (20) or Latent Space Optimization (18) to enforce the data consistency.
-Depending on the specific task, one approach may generate $\bm{x}_{k}$’s with higher quality than the other.
-In practice, as shown in Figures 14 and 15, we observe that for motion deblurring and Gaussian deblurring, the initial reconstructions generated with Pixel Space Optimization have significantly higher quality compared to the ones generated with Latent Space Optimization.
+As discussed in Section 3.4, in the case of LDMs, one can use either Pixel Space Optimization (30) or Latent Space Optimization (27) to enforce the data consistency.
+Depending on the specific task, one approach may generate $\bm{x}_{k}$ with higher quality than the other.
+In practice, as shown in Figures 15 and 16, for motion deblurring and Gaussian deblurring, the initial reconstructions generated with Pixel Space Optimization have significantly higher quality compared to the ones generated with Latent Space Optimization.
 Furthermore, Pixel Space Optimization is much faster than the Latent Space Optimization since the gradient computation of the former approach does not require backpropagation through the deep decoder $\mathcal{D}(\cdot)$.
@@
 Figure 15: Comparison between Pixel Space Optimization and Latent Space Optimization for Motion Deblurring.
-Initial Reconstructions obtained via Pixel Space Pptimization have much better quality compared to the ones obtained from Latent Space Optimization which in turn result in final reconstructions with higher quality.
+Initial reconstructions obtained via Pixel Space Optimization have much better quality compared to the ones obtained from Latent Space Optimization which in turn result in final reconstructions with higher quality.
 Figure 16: Comparison between Pixel Space Optimization and Latent Space Optimization for Gaussian Deblurring.
 Initial reconstructions obtained via Pixel Space Optimization have much better quality compared to the ones obtained from Latent Space Optimization which in turn result in final reconstructions with higher quality.
-A.2 Choice of Hyperparameters
-In this section, we specify the hyperparameters used for generating the results in Section 5.
-For the baselines, we either followed the instructions given in each respective work or used the parameters provided in their code.
-A.2.1 Experiments with Pixel Space Diffusion Models
-For the hyperparameters used for the pixel space experiments, we tabulate all of our values for each dataset in Table 6.
-Recall that for $T_{k}$, we choose a linearly decaying time schedule.
-We start at a value $T_{0}$, which is specified in the table.
-For box inpainting, we used a larger value of $T_{0}$ to compensate for the large artifacts that exist within the box region.
-The diffusion purification is performed with 20 deterministic DDIM steps.
-We optimize the data fidelity objective with momentum-based gradient descent with the momentum hyperparameter set to $0.9$ for both FFHQ and LSUN-Bedrooms datasets.
-To achieve the best performance, we use Pixel Space Optimization for Gaussian deblurring and motion deblurring while using Latent Space Optimization for box inpainting and super-resolution.
-Again, we fix the total number of gradient steps as 1000 throughout the $K$ iterations, which means we take $\frac{1000}{K}$ gradient steps in each round of data fidelity optimization.
+Appendix B Choice of Hyperparameters
+In this section, we specify the hyperparameters of DCDP for generating the results in Section 4. For the baselines, we tune the hyperparameters via grid search to obtain the best possible performance.
+Experiments with Pixel Space Diffusion Models
+For the pixel space experiments, we tabulate the hyperparameters for each dataset in Table 6 (linear inverse problems) and in Table 8 (nonlinear deblurring).
+Recall that for $t_{k}$, we use a linearly decaying time schedule, which starts at a initial value $t_{1}$ and ends at $t_{K}$.
+For linear inverse problems, the DCDP-DDIM and DCDP-Tweedie share the same set of hyperparameters.
+The diffusion purification in DCDP-DDIM is performed with 20 deterministic DDIM steps.
+We optimize the data fidelity objective with momentum-based gradient descent with the momentum hyperparameter set to $0.9$ for all tasks.
+For linear inverse problems, we fix the total number of gradient steps as 1000 throughout the $K$ iterations, which means $\tau=\frac{1000}{K}$ in each round of data fidelity optimization.
+For nonlinear deblurring, we instead use a total number of 2000 gradients steps, i.e., $\tau=\frac{2000}{K}$ in each round of data fidelity optimization.
 Dataset Super Resolution (4×4$\times$4 ×) Gaussian Deblurring Motion Deblurring Inpainting (Box)
-$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
-LSUN-Bedrooms $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $20$ $400$ $10^{3}$ $20$ $700$
-FFHQ $10^{3}$ 10 $400$ $10^{5}$ $10$ $400$ $10^{5}$ $20$ $400$ $10^{3}$ $20$ $700$
-Table 6: Table of hyperparameters used for experiments in the pixel space.
-$\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule.
+$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
+LSUN-Bedrooms $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{5}$ $20$ $50$ $400$ 00 $10^{3}$ $20$ $50$ $700$ 00
+FFHQ $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{5}$ $20$ $50$ $400$ 00 $10^{3}$ $20$ $50$ $700$ 00
+Table 6: Hyperparameters of DCDP with pixel space diffusion models for solving linear inverse problems. For linear inverse problems, DCDP-Tweedie and DCDP-DDIM share the same set of hyperparameters.
 Dataset Super Resolution (4×4$\times$4 ×) Gaussian Deblurring Motion Deblurring Inpainting (Box)
-$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
-LSUN-Bedrooms $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{3}$ $20$ $500$
-FFHQ $10^{3}$ $10$ $400$ $10^{5}$ $10$ $400$ $10^{5}$ $4$ $400$ $10^{3}$ $20$ $500$
-Table 7: Table of hyperparameters used for experiments in the latent space.
-$\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule.
+$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
+LSUN-Bedrooms $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{3}$ $20$ $50$ $500$ 00
+FFHQ $10^{3}$ $10$ $100$ $400$ 00 $10^{5}$ $4$ $250$ $400$ 00 $10^{5}$ $10$ $100$ $400$ 00 $10^{3}$ $20$ $50$ $500$ 00
+Table 7: Hyperparameters of DCDP with latent space diffusion models for solving linear inverse problems. For linear inverse problems, DCDP-Tweedie and DCDP-DDIM share the same set of hyperparameters.
+Dataset DCDP-DDIM DCDP-Tweedie DCDP-LDM-DDIM DCDP-LDM-Tweedie
+$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
+ImageNet $10^{3}$ $20$ $100$ $400$ $100$ $10^{3}$ $20$ $100$ $400$ $50$ $5\times 10^{3}$ $20$ $100$ $300$ $50$ $5\times 10^{3}$ $20$ $100$ $300$ $50$
+Table 8: Hyperparameters of DCDP for solving nonlinear deblurring.
 Dataset Super Resolution (4×4$\times$4 ×) Gaussian Deblurring Motion Deblurring Inpainting (Box)
-$\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$ $\alpha$ $K$ $T_{0}$
-LSUN-Bedrooms $10^{3}$ $20$ $1$ $10^{5}$ $20$ $1$ $10^{5}$ $20$ $1$ $10^{3}$ $20$ $5$
-Table 8: Table of hyperparameters used for experiments with CMs.
-$\alpha$ refers to the learning rate used in the data consistency optimization, $K$ refers to the total number of iterations, and $T_{0}$ refers to the starting point for the linearly decaying time schedule.
-Different from the typical diffusion models used in previous experiments which use a noise scale in the range of [0,1000], CMs are trained in the range of [0.002,80].
+$lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$ $lr$ $K$ $\tau$ $t_{1}$ $t_{K}$
+LSUN-Bedrooms $10^{3}$ $20$ $50$ $1$ 00 $10^{5}$ $20$ $50$ $1$ 00 $10^{5}$ $20$ $50$ $1$ 00 $10^{3}$ $20$ $50$ $5$ 00
+Table 9: Hyperparameters of DCDP with CMs.
+Different from the diffusion models in previous experiments which specify the noise level via discrete forward steps (ranging from 0 to 1,000), CMs are trained on a continous range of Gaussian noise with standard deviation spanning [0.002,80].
+Therefore, $t_{1}$ and $t_{K}$ in this table stands for the standard deviation of the noise rather than the timestep.
 Method Super Resolution 4×4$\times$4 × Inpainting (Box)
@@
 ADMM-PnP [49] $0.307\pm 0.03$ $21.54\pm 3.12$ $0.766\pm 0.03$ $0.251\pm 0.02$ $18.26\pm 1.86$ $0.724\pm 0.02$
-Ours Pixel (Version 1) $\bm{0.154}\pm 0.03$ $\underline{29.43}\pm 2.01$ $\underline{0.844}\pm 0.03$ $\underline{0.052}\pm 0.01$ $\underline{25.54}\pm 3.01$ $\underline{0.917}\pm 0.01$
-Ours Pixel (Version 2) $0.185\pm 0.03$ $\bm{30.86}\pm 2.11$ $\bm{0.878}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\bm{28.02}\pm 2.74$ $\bm{0.935}\pm 0.01$
+DDNM [13] $0.197\pm 0.04$ $28.03\pm 1.46$ $0.795\pm 0.08$ $0.235\pm 0.06$ $24.47\pm 2.38$ $0.837\pm 0.08$
+RED-Diff [45] $0.212\pm 0.06$ $27.18\pm 1.23$ $0.781\pm 0.07$ $0.165\pm 0.08$ $24.22\pm 3.16$ $0.824\pm 0.05$
+DiffPIR [50] $0.260\pm 0.06$ $26.64\pm 1.78$ $0.767\pm 0.06$ $0.108\pm 0.07$ $23.89\pm 1.89$ $0.814\pm 0.06$
+DCDP-DDIM $\bm{0.154}\pm 0.03$ $\underline{29.43}\pm 2.01$ $\underline{0.844}\pm 0.03$ $\underline{0.052}\pm 0.01$ $\underline{25.54}\pm 3.01$ $\underline{0.917}\pm 0.01$
+DCDP-Tweedie $0.185\pm 0.03$ $\bm{30.86}\pm 2.11$ $\bm{0.878}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\bm{28.02}\pm 2.74$ $\bm{0.935}\pm 0.01$
 PSLD [18] $0.309\pm 0.06$ $26.22\pm 2.07$ $0.758\pm 0.06$ $0.248\pm 0.06$ $25.42\pm 2.15$ $0.788\pm 0.05$
-Ours LDMs (Version 1) $\bm{0.183}\pm 0.04$ $\underline{29.16}\pm 1.65$ $\underline{0.826}\pm 0.03$ $\bm{0.166}\pm 0.03$ $\underline{26.50}\pm 2.20$ $\underline{0.860}\pm 0.02$
-Ours LDMs (Version 2) $\underline{0.215}\pm 0.04$ $\bm{30.61}\pm 1.97$ $\bm{0.866}\pm 0.03$ $\underline{0.204}\pm 0.04$ $\bm{27.47}\pm 2.05$ $\bm{0.868}\pm 0.03$
+DCDP-LDM-DDIM $\bm{0.183}\pm 0.04$ $\underline{29.16}\pm 1.65$ $\underline{0.826}\pm 0.03$ $\bm{0.166}\pm 0.03$ $\underline{26.50}\pm 2.20$ $\underline{0.860}\pm 0.02$
+DCDP-LDM-Tweedie $\underline{0.215}\pm 0.04$ $\bm{30.61}\pm 1.97$ $\bm{0.866}\pm 0.03$ $\underline{0.204}\pm 0.04$ $\bm{27.47}\pm 2.05$ $\bm{0.868}\pm 0.03$
 Table 10: Quantitative results of super resolution and inpainting on the FFHQ dataset for both pixel-based and LDMs.
@@
 MCG [11] $0.164\pm 0.03$ $25.03\pm 3.62$ $0.790\pm 0.06$ $0.343\pm 0.04$ $24.88\pm 2.01$ $0.708\pm 0.05$
+DDNM [13] $0.217\pm 0.04$ $28.23\pm 1.69$ $0.805\pm 0.08$
+RED-Diff [45] $0.242\pm 0.06$ $25.25\pm 2.17$ $0.791\pm 0.05$ $0.268\pm 0.08$ $24.32\pm 2.67$ $0.754\pm 0.05$
+DiffPIR [50] $0.236\pm 0.06$ $27.34\pm 2.15$ $0.778\pm 0.04$ $0.255\pm 0.07$ $26.69\pm 1.49$ $0.856\pm 0.06$
 ADMM-PnP [49] $0.421\pm 0.03$ $20.86\pm 2.16$ $0.664\pm 0.05$ $0.522\pm 0.02$ $17.91\pm 1.99$ $0.547\pm 0.06$
-Ours Pixel (Version 1) $\bm{0.140}\pm 0.02$ $\underline{30.41}\pm 1.93$ $\underline{0.864}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\underline{37.56}\pm 1.77$ $\underline{0.963}\pm 0.01$
-Ours Pixel (Version 2) $0.173\pm 0.03$ $\bm{31.67}\pm 2.12$ $\bm{0.891}\pm 0.03$ $\underline{0.060}\pm 0.01$ $\bm{38.37}\pm 1.88$ $\bm{0.970}\pm 0.01$
+DCDP-DDIM $\bm{0.140}\pm 0.02$ $\underline{30.41}\pm 1.93$ $\underline{0.864}\pm 0.03$ $\bm{0.051}\pm 0.01$ $\underline{37.56}\pm 1.77$ $\underline{0.963}\pm 0.01$
+DCDP-Tweedie $0.173\pm 0.03$ $\bm{31.67}\pm 2.12$ $\bm{0.891}\pm 0.03$ $\underline{0.060}\pm 0.01$ $\bm{38.37}\pm 1.88$ $\bm{0.970}\pm 0.01$
 PSLD [18] $0.270\pm 0.05$ $27.46\pm 1.84$ $0.790\pm 0.05$ $0.379\pm 0.05$ $22.35\pm 1.85$ $0.641\pm 0.07$
-Ours LDMs (Version 1) $\underline{0.198}\pm 0.03$ $\underline{28.91}\pm 1.98$ $\underline{0.825}\pm 0.04$ $\underline{0.152}\pm 0.03$ $\underline{30.11}\pm 1.93$ $\underline{0.851}\pm 0.03$
-Ours LDMs (Version 2) $\bm{0.171}\pm 0.02$ $\bm{30.45}\pm 1.86$ $\bm{0.877}\pm 0.03$ $\bm{0.108}\pm 0.02$ $\bm{31.48}\pm 1.94$ $\bm{0.894}\pm 0.02$
+DCDP-LDM-DDIM $\underline{0.198}\pm 0.03$ $\underline{28.91}\pm 1.98$ $\underline{0.825}\pm 0.04$ $\underline{0.152}\pm 0.03$ $\underline{30.11}\pm 1.93$ $\underline{0.851}\pm 0.03$
+DCDP-LDM-Tweedie $\bm{0.171}\pm 0.02$ $\bm{30.45}\pm 1.86$ $\bm{0.877}\pm 0.03$ $\bm{0.108}\pm 0.02$ $\bm{31.48}\pm 1.94$ $\bm{0.894}\pm 0.02$
 Table 11: Quantitative results of Gaussian and motion deblurring on the FFHQ dataset for both pixel-based and LDMs.
@@
 The Best Results are in bold and second best results are underlined for pixel space models and LDMs respectively.
-A.2.2 Experiments with LDMs
-Similarly, we specify the hyperparameters used for experiments using LDMs in Table 7.
-For box inpainting, similar to the pixel space experiments, we used a larger value of $T_{0}$ to compensate for the artifacts seen in the box region.
-The diffusion purification is performed with 20 deterministic DDIM steps.
-For the data fidelity objectives, the same procedure done in the pixel space was applied for the LDM case.
-A.2.3 Experiments with CMs
+Experiments with LDMs
+We specify the hyperparameters used for experiments using LDMs in Table 7 (linear inverse problems) and in Table 8 (nonlinear deblurring).
+Again, DCDP-LDM-DDIM and DCDP-LDM-Tweedie share the same set of hyperparameters for linear inverse problems.
+The diffusion purification in DCDP-LDM-DDIM is performed with 20 deterministic DDIM steps and for the data fidelity objectives, the same procedure done in the pixel space is applied on the LDMs.
+To achieve the best performance, we use Pixel Space Optimization for Gaussian deblurring and motion deblurring while using Latent Space Optimization for box inpainting, super-resolution and nonlinear deblurring.
+Again, we fix the total number of gradient steps for linear inverse problems as 1000 while using 2000 steps for nonlinear deblurring.
+Experiments with CMs
 Lastly, we specify the hyperparameters used for experiments using CMs in Table 9.
-Notice that different from the typical diffusion model used previously which use a noise scale in the range of [0,1000], CMs are traind in the range of [0.002,80].
-Again, the diffusion purification is performed with 20 deterministic DDIM steps.
-A.3 Dealing with Measurement Noise
-When the measurements are noiseless, a linear decaying schedule starts from $T_{0}$ and ends at $T_{K}=0$ is capable of resulting in high quality reconstructions and we can use a relatively high learning rate for the data fidelity optimization.
+Note that different from the typical diffusion models which specify the noise level via discrete forward steps (commonly ranging from 0 to 1,000), CMs are trained on a continuous range of Gaussian noise with standard deviation spanning [0.002,80].
+Appendix C Dealing with Measurement Noise
+When the measurements are noiseless, a linear decaying schedule starting from $t_{1}$ and ending at $t_{K}=0$ is capable of resulting in high-quality reconstructions, and we can use a relatively high learning rate for the data fidelity optimization.
 On the contrary, in the presence of measurement noise, the data fidelity optimization in (18) will inevitably introduce additional artifacts even if $\bm{v}_{k-1}$ is the ground truth image.
-To achieve robust recovery, one may instead: (i) use a relatively small learning rate for data fidelity optimization to prevent overfitting to the noise, and (ii) end the time schedule at $t_{K}=T$, where $T>0$, to compensate for the artifacts resulting from the measurement noise.
-As shown in  Figure 16, compared to the clean measurements case, due to the additive noise, the data fidelity optimization always introduce artifacts in $\bm{x}_{k}$.
+To achieve robust recovery, one may instead: (i) use a relatively small learning rate for data fidelity optimization to prevent overfitting to the noise, and (ii) end the time schedule at $t_{K}=\tilde{t}$, where $\tilde{t}>0$, to compensate for the artifacts resulting from the measurement noise.
+As shown in Figure 19, compared to the clean measurements case, due to the additive noise, the data fidelity optimization always introduces artifacts in $\bm{x}_{k}$.
 This issue can be alleviated by utilizing a non-zero diffusion purification strength at the final iteration.
-For the experiments in Section 5.3, all the hyperparameters are kept the same as described in Section A.2.1, but with the following changes:
+For the experiments in Section 4.3, all the hyperparameters are kept the same as described in Appendix B, but with the following changes:
   • The learning rate for motion deblurring changes from $1e^{5}$ to $1e^{4}$.
-  • The time schedules for super-resolution and motion deblurring start from $t_{0}=400$ and decay linearly to $250$ and $300$ respectively.
+  • The time schedules for super-resolution and motion deblurring start from $t_{1}=400$ and decay linearly to $250$ and $300$ respectively.
 With the changes made above, our algorithm achieves reconstruction performance comparable to that of DPS with measurement noise.
@@
 Better performance can be achieved by decreasing $t_{K}$ in the case of $\sigma=0.05$ or $\sigma=0.075$.
-Appendix B Additional Quantitative Results
-In this section, we provide additional quantitative results for solving inverse problems on the FFHQ dataset.
-The results are presented in Tables 9 and 10.
-Throughout all of our experiments, we observe that our proposed algorithms outperform the baselines for various image restoration problems.
-(a) Reference and Measurements for Super-Resolution (4x). The noisy measurements are corrupted by random Gaussian noise with $\sigma=0.1.$
-(b) Reconstruction Process with Clean Measurements.
+Figure 17: Reference and Measurements for Super-Resolution (4x). The noisy measurements are corrupted by random Gaussian noise with $\sigma=0.1$.
+Figure 18: Reconstruction Process with Clean Measurements.
 The reconstruction process takes $K=10$ iterations in total with a time schedule starting at $t_{1}=400$ and linearly decaying to $t_{10}=0$.
@@
 Due to the space limit, only reconstructions at timesteps 1, 3, 5, 7 and 10 are shown.
-(c) Reconstruction Process with Noisy Measurements.
+Figure 19: Reconstruction Process with Noisy Measurements.
 To compensate the artifacts introduced by the measurement noise, the reconstruction process takes $K=10$ iterations in total with a time schedule starting at $t_{1}=400$ and linearly decaying to $t_{10}=200$.
 The learning rate for data fidelity optimization is $lr=100$.
-Figure 16: Reconstruction Process for Super-Resolution (4x) with Pixel Space Diffusion Models.
+Appendix D Additional Experiment Results and Implementation Details
+Quantitative Results on FFHQ.
+Quantitative results for solving inverse problems on the FFHQ dataset are presented in Tables 10 and 11.
+Note that DCDP outperforms the baselines for various inverse problems.
+Convergence Behavior.
+In Section 4.3 (Figure 8), we show that with a linear decaying time schedule for diffusion purification, DCDP iteratively improves the LPIPS scores and progressively converges towards a high-quality reconstruction.
+Similar convergence behavior holds when measuring the reconstruction quality using PSNR and SSIM, as shown in Figure 20.
+In this experiment, DCDP is applied to the nonlinear deblurring task, with the standard metrics averaged over 10 randomly selected images from the ImageNet dataset.
+The hyperparameters are the same as those reported in Appendix B (Table 8).
+Note that DCDP substantially improves upon the pure data fidelity optimization, implying diffusion purification is an efficient operator to enforce the image priors embedded in the pre-trained diffusion models.
+Impact of Number of DDIM Steps.
+Here we provide experimental details for the ablation study on the impact of DDIM steps in Section 4.3.
+The experiments are conducted using a pixel space diffusion model based on DCDP (DCDP-DDIM) for solving the nonlinear deblurring problem.
+Except for the number of DDIM steps, all other hyperparameters are the same as those reported in Appendix B (Table 8).
+The standard metrics are averaged over 10 randomly selected images from the ImageNet dataset.
+Impact of Number of Total Number of Iterations of $K$.
+Here we provide experimental details for the ablation study on the impact of $K$ in Section 4.3.
+The experiments are conducted using pixel space diffusion models (both DCDP-DDIM and DCDP-Tweedie).
+Except for $K$, all other hyperparameters are the same as those reported in Appendix B (Table 8).
+The standard metrics are averaged over 10 randomly selected images from the ImageNet dataset.
+Figure 20: Convergence of our methods in terms of PSNR, SSIM and LPIPS for solving nonlinear deblurring.
+DCDP greatly improves upon the pure data fidelity optimization by leveraging the image priors embedded in the diffusion models.